### v1

# Landing with the Score: Riemannian Optimization through Denoising

Abstract
Under the data manifold hypothesis, high-dimensional data are concentrated near a low-dimensional manifold. We study the problem of Riemannian optimization over such manifolds when they are given only implicitly through the data distribution, and the standard manifold operations required by classical algorithms are unavailable. This formulation captures a broad class of data-driven design problems that are central to modern generative AI. Our key idea is to introduce a link function that connects the data distribution to the geometric operations needed for optimization. We show that this function enables the recovery of essential manifold operations, such as retraction and Riemannian gradient computation. Moreover, we establish a direct connection between our construction and the score function in diffusion models of the data distribution. This connection allows us to leverage well-studied parameterizations, efficient training procedures, and even pretrained score networks from the diffusion model literature to perform optimization. Building on this foundation, we propose two efficient inference-time algorithms—Denoising Landing Flow (DLF) and Denoising Riemannian Gradient Descent (DRGD)—and provide theoretical guarantees for both feasibility (approximate manifold adherence) and optimality (small Riemannian gradient norm). Finally, we demonstrate the effectiveness of our approach on finite-horizon reference tracking tasks in data-driven control, highlighting its potential for practical generative and design applications.
1 Introduction
Riemannian optimization Boumal (2023); Absil et al. (2008); Hu et al. (2020) considers minimizing an objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ over an explicitly known embedded submanifold $\mathcal{M}\subseteq\mathbb{R}^{d}$,
$\min_{x\in\mathcal{M}}f(x).$ (1)
Problem (1) is ubiquitous in fields of machine learning and control and encompasses problems such as independent component analysis Nishimori (1999), low-rank matrix completion Vandereycken (2013), training of orthogonally normalized neural networks Bansal et al. (2018), the control of rigid bodies Duong et al. (2024), as well as sensor network localization Patwari & Hero (2004) and many others. Compared to general constrained optimization, Riemannian optimization promises the advantage of exploiting the natural geometry of the problem, producing feasible iterates and increased numerical robustness Boumal (2023).
In contrast to the above classical setup, in this work we focus on the setting where the manifold $\mathcal{M}$ is given implicitly through a measure $\mu=\mu_{\mathrm{data}}$ supported on $\mathcal{M}$. This perspective is especially relevant in view of the manifold hypothesis (Loaiza-Ganem et al., 2024), which posits that many real-world data sets lie (approximately) on a manifold with dimension much smaller than that of the ambient space Fefferman et al. (2016). Importantly, such data manifolds are not just low-dimensional geometric structures—they also capture rich semantic meaning. For instance, the image manifold corresponds to photo-realistic images Pope et al. , the system behavior manifold to dynamically feasible input-output trajectories Willems & Polderman (1997), while the manifold of airfoils represents aerodynamically viable shapes Zheng et al. (2025). The optimization problem (1) in this setting thus encompasses a broad class of modern tasks, including airfoil design (Chen et al., 2025), reinforcement learning (Lee & Choi, 2025), and Bayesian inverse problems (Chung et al., 2022b). Driven by the growing demand in generative AI and data-driven design, this calls for a paradigm shift: moving from optimization over explicitly known manifolds, where classical Riemannian optimization applies, to optimization over data manifolds that are accessible only implicitly through samples. In this data-driven regime, methods from classical Riemannian optimization cannot be applied directly, since they rely on explicit manifold operations such as tangent-space projection, retraction, or exponential maps (Boumal et al., 2019; Boumal, 2023).
While there has been extensive research on manifold learning Meilă & Zhang (2024); Lin & Zha (2008); Cayton et al. (2005); Belkin & Niyogi (2005) in the past, none of these works addressed (1) from an optimization point of view and focused instead on learning the manifold geometry. Moreover, the non-parametric nature of these methods limits their applicability to very low-dimensional manifolds and cannot exploit the successful inductive bias and the exceptional representation power of modern neural networks.
In this work, we propose a data-driven approach to recover the fundamental operations needed for optimization on manifolds. Starting from the data distribution $\mu_{\mathrm{data}}$, we smooth it with a Gaussian kernel to obtain
$p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu_{\mathrm{data}},$ (2)
and define the associated link function
$\ell_{\sigma}(x)=\frac{1}{2}\lVert x\rVert^{2}-\sigma^{2}\log p_{\sigma}(x).$ (3)
We show that, as the smoothing parameter $\sigma$ decreases, the gradient $\nabla\ell_{\sigma}$ recovers the projection back to the manifold, while the Hessian $\nabla^{2}\ell_{\sigma}$ recovers the projection onto its tangent space. These results reveal that core ingredients of Riemannian optimization—such as retraction and gradient computation—can be implemented directly from the derivative information of $\ell_{\sigma}$, even when the manifold itself is only given implicitly through data.
Given the above novel theoretical findings, a key practical challenge is how to access (even approximately) the gradient $\nabla\ell_{\sigma}$ and Hessian $\nabla^{2}\ell_{\sigma}$ when only samples from $\mu_{\mathrm{data}}$ are available. A crucial observation is that $p_{\sigma}$ coincides with the marginal distribution of the Variance-Exploding SDE (VE-SDE), with analogous arguments for VP-SDE and DDPM formulations (Ho et al., 2020; Song et al., 2020). In this setting, the gradient of $\log p_{\sigma}$ – commonly referred to as the score function – plays a central role. In diffusion models, this score is parameterized by a neural network and learned directly from samples of $\mu_{\mathrm{data}}$, typically via denoising score matching (Vincent, 2011; Song et al., 2020). It is therefore natural to adopt the same methodology here: a neural network can be trained to represent $\nabla\ell_{\sigma}$, while the Hessian $\nabla^{2}\ell_{\sigma}$ can be recovered by computing its Jacobian. This approach offers two key advantages: (i) strong inductive biases can be incorporated through the neural network parameterization, and (ii) efficient, well-established training techniques from the diffusion model literature can be directly leveraged. Taken together, the theoretical link between $\ell_{\sigma}$ and manifold geometry, and the practical machinery for learning the score, form the foundation of the paradigm shift: from classical Riemannian optimization with explicit manifold knowledge to a data-driven framework where geometry is recovered from samples—thus enabling principled manifold optimization in generative and design-driven applications.
Building on these insights, we propose two algorithms for optimization over data manifolds – denoising landing flow (DLF) and denoising Riemannian gradient descent (DRGD) – to the best of our knowledge, the first such framework in the literature. The prerequisite is access to an approximate function $s_{\theta}^{\sigma}\simeq\nabla\ell_{\sigma}$, which can be learned from data using standard diffusion model parameterization and training given samples from $\mu_{\mathrm{data}}$.
  • Our first method directly employs $s_{\theta}^{\sigma}$ and its Jacobian (with $\sigma$ fixed as a small constant) to substitute the corresponding manifold operations in Riemannian gradient descent: each iteration computes a Riemannian gradient step and then applies retraction to (approximately) return to the manifold.
  • Since learning the score is inevitably imperfect when $\sigma$ is small, we also design a complementary landing-type algorithm that avoids staying close to the feasible set at every step. Instead, it augments the target function $f$ with the regularization term $\ell_{\sigma}$ and performs gradient descent on this regularized objective.
For both algorithms, we establish non-asymptotic convergence guarantees to approximate stationary points, measured by small Riemannian gradient norm, as $\sigma\to 0$. Importantly, our methods require only inference of the neural network and gradients with respect to its inputs—not with respect to the network parameters. Thus, if a pretrained score network is already available for a given task, no additional training is required to enable Riemannian optimization on the corresponding data manifold. Viewed from this perspective, our approach can also be interpreted as an inference-time algorithm, aligning with a growing trend in modern machine learning research.
1.1 Our contributions
We summarize our contributions and give the outline of the paper as follows.
  • Data-driven recovery of manifold operations. By introducing the link function $\ell_{\sigma}$, we develop strategies to recover core manifold operations directly from data. In Section 3, by leveraging the score function from diffusion models to train and parameterize $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$, we bridge the gap between the requirements of classical Riemannian optimization (which assumes explicit manifold knowledge) and the emerging need to optimize over data manifolds that are only implicitly available.
  • First algorithms for optimization over data manifolds. We propose two algorithms – denoising landing flow (DLF) via (8) in Section 4 and denoising Riemannian gradient descent (DRGD) via (12) in Section 5to the best of our knowledge, the first in the literature – that exploit operations enabled by a pretrained score function. Our methods require only inexpensive inference-time queries and back-propagation with respect to the input of the neural network, making them computationally efficient and readily applicable when pretrained scores are available.
  • Rigorous non-asymptotic guarantees. We provide the first non-asymptotic convergence analysis for optimization over data manifolds. Our main results (Theorem 3 and Theorem 5) ensure both approximate feasibility (output is close to the manifold) and approximate optimality (small Riemannian gradient norm). A key technical contribution is a novel non-asymptotic analysis of how $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$ approximate the true manifold operations when $\sigma$ is small but non-zero (Theorem 1).
Finally in Section 6 we validate our findings by numerically simulating the proposed flow on well-known manifolds from Riemannian optimization literature and data-driven optimal control for reference trajectory tracking. Our finding are that we can generate feasible points on the manifold with objective values far lower than the ones available in the training set – see Figure 1 – and shows the effectiveness of exploiting the strong inductive biases of modern deep learning for the classical field of constrained optimization.
Figure 1: Optimized trajectory (orange) on the system trajectory manifold for the unicycle car model that is desired to track the reference trajectory (black, thin). The closest tracking trajectory in the set of available manifold samples is given in (green, dotted). See Section 6 for more details.
1.2 Related Work
Riemannian Optimization.
Riemannian optimization (RO) was originally developed under the assumption that the constraint manifold is explicitly known, either through closed-form descriptions such as matrix manifolds or via nonlinear equality constraints (Sato, 2021; Boumal, 2023). A prototypical algorithm in this setting is Riemannian gradient descent, which updates by taking a gradient step along the tangent space and then retracting back to the manifold (Boumal et al., 2019). Such algorithms guarantee that the iterates remain on the manifold at every iteration and are thus referred to as feasible methods. In contrast, infeasible methods, where the iterates are not constrained to stay on $\mathcal{M}$, have also been studied—for example, augmented Lagrangian approaches (Xie & Wright, 2021). However, these typically involve solving complicated optimization subproblems in each inner loop, making them computationally expensive in practice. More recently, a new class of landing-type algorithms has emerged: instead of enforcing feasibility at every step, they regularize the objective function with the distance to the manifold and then perform gradient flow or descent on this regularized objective. Such methods have demonstrated strong empirical performance, offering a promising alternative to classical approaches (Ablin & Peyré, 2022; Schechtman et al., 2023).
Manifold Learning.
High-dimensional data in modern machine learning often exhibit an intrinsic lower-dimensional structure. Such structure is of central importance: it enables more efficient representation and compression of data, facilitates interpretability by revealing meaningful semantic organization, and provides a foundation for designing algorithms that exploit geometry rather than ambient dimensionality. The task of uncovering this structure is commonly referred to as manifold estimation or manifold learning, with a large body of work devoted to this goal. Classical approaches include Isomap (Tenenbaum et al., 2000), Laplacian Eigenmaps (Belkin & Niyogi, 2003), and Locally Linear Embedding (LLE) (Roweis & Saul, 2000), Diffusion Map (Coifman & Lafon, 2006), among many subsequent developments, e.g. (Zhou et al., 2020). We defer the discussion on diffusion-model-related manifold learning literature to the next paragraph.
Diffusion models.
Diffusion models have achieved remarkable success in generative modeling, where the goal is to generate new samples consistent with an underlying data distribution $\mu_{\mathrm{data}}$. A central ingredient of these methods is the learning of the score function—the gradient of the log-density of the diffused distribution $p_{\sigma}$ Tang & Zhao (2025); Song et al. (2020). In practice, the score is parameterized by a neural network whose architectural design has been extensively studied, and its training is carried out using well-established techniques such as denoising score matching (Song et al., 2020). This combination of principled theory and mature practice has made diffusion models one of the most effective tools for data-driven generative modeling.
When the data distribution resides on a manifold, a recent observation in the diffusion model community is that the score is asymptotically orthogonal to the manifold surface Stanczuk et al. (2024). This observation has been exploited for the estimation of the manifold dimension. See (Kamkari et al., 2024) for the same task. Furthermore, Ventura et al. (2024) has shown that in the case of linear manifolds (i.e. affine subspaces), the Jacobian of the score scaled by the diffusion temperature asymptotically approximates the projection of the manifold onto the normal space and used it to study the geometric phases of diffusion models.
Recent work on the statistical complexity of diffusion models under the manifold hypothesis provides indirect evidence that these models capture geometric information about the underlying data manifold (Oko et al., 2023; Tang & Yang, 2024). In particular, the sample complexity required to learn the data distribution $\mu_{\mathrm{data}}$ depends only on the intrinsic dimension of the manifold, rather than on the ambient dimension.
Two Formulations: Optimization and Posterior Sampling.
To avoid possible confusion, we stress the distinction between our optimization formulation and the posterior sampling literature. In short, our optimization formulation in eq. 1 enforces the manifold constraint directly. This ensures (approximate) feasibility at the final step and guarantees that the optimization process remains semantically meaningful, which is not guaranteed by the sampling formulation, as discussed below.
Across the literatures Classifier(-Free) Guidance in Diffusion Models (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) and the Plug-and-Play Framework in Bayesian Inverse Problems (Venkatakrishnan et al., 2013; Laumont et al., 2022; Pesme et al., 2025; Graikos et al., 2022; Chung et al., 2022a), a unifying perspective is the task of sampling from a posterior distribution of the form
$p_{\mathrm{post}}\propto p_{\mathrm{pre}}\exp\left(-\tfrac{r}{\alpha}\right),$
where $r$ denotes a cost function to be minimized—corresponding to our objective ff—and $\alpha>0$ is a temperature-like parameter (Domingo-Enrich et al., 2024). In Classifier(-Free) Guidance, $r$ encodes a classifier signal (e.g., specified by a prompt) and in Bayesian inverse problems, $r$ is the negative log-likelihood of observations. Meanwhile, $p_{\mathrm{pre}}$ serves as a prior distribution, typically derived from a large-scale pretrained generative model, which is expected to capture the semantic structure of the data manifold. Sampling from $p_{\mathrm{post}}$ thus aims to balance semantic plausibility with low cost.
If $p_{\mathrm{pre}}$ were exactly supported on the data manifold $\mathcal{M}$ and $\alpha\to 0$, this formulation would reduce to the constrained optimization problem eq. 1. In practice, however, the situation is very different: the distribution induced by a pretrained diffusion model is not concentrated on a low-dimensional manifold but rather has support of non-zero Lebesgue measure in the ambient space—indeed, in many cases, essentially the full space (due to the noisy generation process of $p_{\mathrm{pre}}$). As a result, when $\alpha$ is set too small, the posterior $p_{\mathrm{post}}$ becomes dominated by the exponential tilt $\exp(-r/\alpha)$, pushing samples into regions far from the true data manifold $\mathcal{M}$ and thereby losing semantic meaning. Consequently, these sampling-based frameworks must carefully tune $\alpha$ to trade off semantic fidelity (staying close to $\mathcal{M}$) against optimization quality (achieving low $r$).
2 Preliminaries
Here we introduce some preliminary notation and concepts to state our results. We refer the reader to the Appendices A and C for more details.
2.1 Manifolds and distance functions
Let $\mathcal{M}\subseteq\mathbb{R}^{d}$ be a $k$-dimensional embedded compact $C^{2}$-submanifold without boundary. For any point $p\in\mathcal{M}$ we denote by $\operatorname{T}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ and $\operatorname{N}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ the tangent and normal spaces of $\mathcal{M}$ at $p$, respectively, and their orthogonal projections by $\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}$ and $\operatorname{P}_{\operatorname{N}_{p}\mathcal{M}}$. The squared distance function is defined by $\operatorname{d}(x)=\inf_{p\in\mathcal{M}}\frac{1}{2}\lVert x-p\rVert^{2}$. For a $C^{2}$-submanifold, then there exists a radius $\tau_{\mathcal{M}}>0$ such that every point in $x\in\mathcal{T}(\tau)=\bigcup_{p\in\mathcal{M}}B_{\tau}(p)$ has a unique projection $\pi(x)\in\mathcal{M}$ such that $\frac{1}{2}\lVert x-\pi(x)\rVert^{2}=\operatorname{d}(x)$ and $x-\pi(x)\in\operatorname{N}_{\pi(x)}\mathcal{M}$. Sets of the form $\mathcal{T}(\tau)$ will be called tubular neighborhoods of $\mathcal{M}$ with closure denoted by $\overline{\mathcal{T}}(\tau)$. One can prove that $\operatorname{d}$ and $\pi$ are differentiable on $\mathcal{T}(\tau_{\mathcal{M}})$ with $\operatorname{d}^{\prime}(x)=x-\pi(x)$ and $\pi^{\prime}(x)$ given explicitly in Appendix A. The maximal principal curvature of a manifold will be denoted by $\kappa_{\mathcal{M}}$. It always holds that $\tau_{\mathcal{M}}\leq 1/\kappa_{\mathcal{M}}$. Finally, for a function $f:\mathcal{M}\to\mathbb{R}$ we denote the Riemannian gradient by $\operatorname{grad}_{\mathcal{M}}f(p)$, which in the case of $f:\mathbb{R}^{d}\to\mathbb{R}$ is given by $\operatorname{grad}_{\mathcal{M}}f(p)=\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}\nabla f(p)$. We write
2.2 The Stein score function and score-based diffusion models
For a Borel probability measure we denote its Gaussian blurring by $p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu$ for $\sigma>0$ with $p_{0}=\mu$. The score function $\nabla\log p_{\sigma}$ of $p_{\sigma}$ and its Jacobian allow for the following interpretation (see Jaffer & Gupta (1972), also Appendix B):
$x+\sigma^{2}\nabla\log p_{\sigma}(x)=\nabla\ell_{\sigma}(x)=\mathbb{E}\nu_{x,\sigma}\,,\text{\ \ }I+\sigma^{2}\nabla^{2}\log p_{\sigma}(x)=\nabla^{2}\ell_{\sigma}(x)=\frac{1}{\sigma^{2}}\operatorname{Cov}(\nu_{x,\sigma})\,$, (4)
where $\ell_{\sigma}$ is the link function (3) and $\nu_{x,\sigma}$ is the posterior distribution observing $x$ under the noise model $p_{\sigma}$ and prior $\mu$. The representation for $\mathbb{E}\nu_{x,\sigma}$ has also been known under the name of Tweedie’s formula Robbins (1992); Efron (2011). The score function has gained recent attention due to its use in score-based diffusion models in the field of generative modelling. Specifically in the so-called variance exploding (VE) diffusion scheme one seeks to learn $\nabla\log p_{\sigma}$ for different noise scales $\sigma$ via a neural network $\texttt{s}(\cdot,\sigma)$ by minimizing the conditional score matching loss $L_{\operatorname{CSM}}(\texttt{s}(\cdot,\sigma))$, which attains its unique minimum in $\texttt{s}(\cdot,\sigma)=\nabla\log p_{\sigma}$. Then $\texttt{s}(\cdot,\sigma)$ is used for sampling from $\mu$ by following a particular reverse-time SDE or ODE flow in the noise scale $\sigma$ (see Appendix C for more details).
3 Score as retraction and tangent space projection
In this section we give another interpretation to the score function $\nabla\log p_{\sigma}(x)$ and its Jacobian $\nabla^{2}\log p_{\sigma}(x)$. Namely, it has been already observed in Stanczuk et al. (2024) that the score function is for $\sigma\to 0$ asymptotically orthogonal to the tangent space of the data manifold. Our first contribution is showing that when the support of the distribution $\mu$ is a manifold $\mathcal{M}$ and $\mu$ is absolutely continuous w.r.t. its volume measure, then both quantities (4) approximate the projection operator $\pi(x)$ and its Jacobian $\pi^{\prime}(x)$ uniformly on tubular neighborhoods of $\mathcal{M}$. Formally we establish the following
Theorem 1 (Main).
Let $\mathcal{M}\subseteq\mathbb{R}^{d}$ be a compact, embedded $C^{3}$-submanifold and $\mu\in\mathcal{P}(\mathbb{R}^{d})$ a Borel probability measure with $\operatorname{supp}\mu=\mathcal{M}$ and $\mu\ll\operatorname{Vol}_{\mathcal{M}}$ such that $\frac{\operatorname{d}\mu}{\operatorname{d}\operatorname{Vol}_{\mathcal{M}}}\in C^{3}(\mathcal{M})$. Then for any $\tau\in(0,\tau_{\mathcal{M}})$ there exist some constants $K=K(\tau,\mathcal{M},\mu)>0$ and $\overline{\sigma}=\overline{\sigma}(\tau,\mathcal{M},\mu)>0$ depending on $\tau$, $\mathcal{M}$ and $\mu$ such that
$\lVert\mathbb{E}\nu_{x,\sigma}-\pi(x)\rVert\leq K\sigma\lvert\log(\sigma)\rvert^{3}\text{\ \ and\ \ }\left\lVert\frac{1}{\sigma^{2}}\operatorname{Cov}(\nu_{x,\sigma})-\pi^{\prime}(x)\right\rVert\leq K\sigma\lvert\log(\sigma)\rvert^{3}$ (5)
for all $\sigma\in(0,\overline{\sigma})$ and $x\in\mathcal{T}(\tau)$.
The proof is deferred to Appendix D and is based on a careful non-asymptotic estimate of the Laplace integral method. As a consequence and together with fact that $\pi^{\prime}(x)$ coincides with $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ for $x\in\mathcal{M}$ (see Appendix A) we obtain the following result.
Corollary 2.
Let $\mathcal{M}$ and $\mu$ be as in Theorem 1 and suppose that $x\in\mathcal{M}$. Then
$\lim_{\sigma\to 0}I+\sigma^{2}\nabla^{2}\log p_{\sigma}(x)=P_{\operatorname{T}_{x}\mathcal{M}}\,.$
In view of Theorem 1 let us abbreviate (4) into
$\operatorname{d}_{\sigma}(x)=-\sigma^{2}\log p_{\sigma}(x)\,,\quad\pi_{\sigma}(x)=x+\sigma^{2}\nabla\log p_{\sigma}(x)\,,\quad P_{\sigma}(x)=I+\sigma^{2}\nabla^{2}\log p_{\sigma}(x)\,$, (6)
with the limiting cases $\operatorname{d}_{0}=\operatorname{d}$, $\pi_{0}=\pi$ and $P_{0}(x)=\pi^{\prime}(x)$. We stress that the expressions in (6) are defined for all $x\in\mathbb{R}^{d}$, whereas $\operatorname{d}$, $\pi$ and $\pi^{\prime}$ in are only sensible in a tubular neighborhood $\mathcal{T}$ of $\mathcal{M}$.
Thus, in theory, a well-trained diffusion model score $\texttt{s}(\cdot,\sigma)$ and its Jacobian $\texttt{s}^{\prime}(\cdot,\sigma)$ allow us to approximate the closest-point projection $\pi(x)$ and the tangent space projection $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ as $\sigma\to 0$ arbitrarily well via its Tweedie score $x+\sigma^{2}\texttt{s}(x,\sigma)$ and its Jacobian, respectively.
4 Denoising Riemannian gradient flow with landing
In this section we show how to use the score function from Section 3 for Riemannian optimization of (1) for some smooth $f\in C^{1}(\mathbb{R}^{d})$. Assuming that we have access to a sufficiently accurate estimate of the Tweedie score in form of a vector function $s\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$ such that $s(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and
$\lVert s(x)-\pi_{\sigma}(x)\rVert\leq\epsilon\text{\ and\ }\lVert s^{\prime}(x)-P_{\sigma}(x)\rVert\leq\epsilon\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$ (7)
for some $\tau\in(0,\tau_{\mathcal{M}})$, we propose for $\eta\geq 0$ the denoising landing flow (DLF)
$\dot{x}$ $=-s^{\prime}(x)\nabla f(s(x))+\eta(s(x)-x)\,.$ (8)
In the exact case $s(x)=\pi_{\sigma}(x)$ and $s^{\prime}(x)=P_{\sigma}(x)$ (i.e. $\epsilon=0$ in (7)), flow (8) is the gradient flow
$\dot{x}$ $=-\nabla F_{\sigma}^{\eta}(x)=-P_{\sigma}(x)\nabla f(\pi_{\sigma}(x))+\eta(\pi_{\sigma}(x)-x)\text{\ \ with\ \ }F_{\sigma}^{\eta}(x)=f(\pi_{\sigma}(x))+\eta\operatorname{d}_{\sigma}(x)$ (9)
and the dynamics in (9) consists of two parts: An approximate projection $P_{\sigma}(x)\nabla f(\pi_{\sigma}(x))$ of the gradient $\nabla f(\pi_{\sigma}(x))$ and an approximate landing term $\eta(\pi_{\sigma}(x)-x)$ corresponding to the penalty function $\eta\operatorname{d}_{\sigma}(x)$. In the further case of $\sigma=0$ and $x(0)\in\mathcal{M}$ the flow (9) reduces to the ordinary Riemannian gradient flow, which has been extensively studied Helmke & Moore (2012); Ambrosio et al. (2005). Interestingly, when $\sigma=0$, but only $x(0)\in\mathcal{T}(\tau)$, then (9) reduces to
[…]
with $H_{x}^{-1}$ a linear operator on $\operatorname{T}_{\pi(x)}\mathcal{M}$ given in Appendix A. In particular the two terms in (10) belong to $\operatorname{T}_{\pi(x)}\mathcal{M}$ and $\operatorname{N}_{\pi(x)}\mathcal{M}$, respectively, and are orthogonal to each other, implying that the distance between $x$ and $\mathcal{M}$ is non-increasing, which allows for perfect landing on the manifold via similar arguments as in Ablin & Peyré (2022); Schechtman et al. (2023), see Theorem 20 in Appendix E.3. For $\sigma>0$ or $\epsilon>0$, the two summands in (9) and (8) in general not perpendicular to each other and the landing is not exact. However, using Theorem 1 we show the following.
Theorem 3.
Consider the flow (8) and $\tau\in(0,\tau_{\mathcal{M}})$. Set $C=\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f)$. Suppose that for some $\epsilon>0$ and $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ with $\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\min\{\tau,\frac{2\tau}{1+C/\eta}\}$ the function $s$ satisfies (7). Then for any $x(0)\in\mathcal{T}(\tau)$ the solution $x(t)$ to (8) exists for all $t\geq 0$ and is contained in $\mathcal{T}(\tau)$. Moreover, every accumulation point $x_{*}$ of this flow satisfies
$\operatorname{dist}_{\mathcal{M}}(x_{*})\leq\tau_{0}:=\frac{1}{2}\left(\frac{C}{\eta}+1\right)(\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3})\,$,
and for the projection $p_{*}=\pi(x_{*})$ it holds that
$\lVert\operatorname{grad}_{\mathcal{M}}f(p_{*})\rVert\leq\left(2(L+C+2\eta)+\frac{(1+C/\eta)/\tau_{\mathcal{M}}}{1-\tau/\tau_{\mathcal{M}}}C\right)(\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3})\,.$ (11)
Thus, Theorem 3 shows that one can still use the flow (8) for a fixed $\sigma>0$ to converge to approximate critical points of the objective $f$, at which the approximation error and norm of the Riemannian gradient are both $\widetilde{O}(\sigma)$ plus the score error $\epsilon$.
Remark 4.
We can evaluate the right hand side of the flow (8) in a single forward-backward pass of the network $s$. Namely, given an input $x$, we compute and store $p=s(x)$ by a forward pass of $s$, while keeping the computational graph of $s(x)$. Then we evaluate $v=\nabla f(p)$ and build the computational graph of $y=\left\langle s(x),v\right\rangle$, while detaching $v$. Finally we backpropagate on $x$ in $y$ to obtain $s^{\prime}(x)v=s^{\prime}(x)\nabla f(s(x))$.
5 Denoising Riemannian gradient descent
In a practical implementation one has to consider a discretized version of the flow (8). A natural alternative is to study the following approximate version of the Riemannian gradient descent Absil et al. (2008); Boumal (2023)
$x_{k+1}=s(x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k}))\,$, (12)
which we term the denoising Riemannian gradient descent (DRGD). Here $s$ acts as an approximate retraction and $s$ as an approximate projection onto the tangent space. We obtain the following convergence result for this algorithm.
Theorem 5.
Let $\tau\in(0,\tau_{\mathcal{M}}/2)$ and set $C=\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f)$, $D=\lVert f|_{\mathcal{T}(\tau)}\rVert_{\infty}$ and
$L_{0}=8C\left(2(\frac{3}{\tau_{\mathcal{M}}}+\tau M)+\frac{1}{\tau_{\mathcal{M}}}\right)+2L\,.$
Suppose that for some $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ and $\epsilon>0$ with $\epsilon^{\prime}:=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$ the function $s$ satisfies (7) and that the step-size $\gamma_{k}$ is constrained by $\gamma_{k}\in[\gamma_{\min},\gamma_{\max}]$ with
$0<\gamma_{\min}<\gamma_{\max}<\min\left\{\frac{2}{L_{0}},\frac{\tau}{C(4+\tau)}\right\}\,.$
Then for any $x_{0}\in\mathcal{T}(\tau/2)$ the iterates $x_{k}$ of (12) satisfy $\{x_{k}\}_{k=1}^{\infty}\subseteq\mathcal{T}(\epsilon^{\prime})\subseteq\mathcal{T}(\tau/2)$ and for the projection $p_{k}=\pi(x_{k})$ the following average-of-gradient-norm condition holds:
$\frac{1}{N}\sum_{k=0}^{N}\lVert\operatorname{grad}_{\mathcal{M}}f(p_{k})\rVert^{2}\leq\frac{4D/N+(8C^{2}\epsilon^{\prime}/\tau_{\mathcal{M}}^{2}+2(2C+L_{0}\gamma_{\max}(C+L)))\epsilon^{\prime}}{\gamma_{\min}(1-\frac{L_{0}}{2}\gamma_{\max})}\,.$
In particular there exists at least one accumulation point $x_{*}\in\overline{\mathcal{T}}(\epsilon)$ of $\{x_{k}\}_{k=0}^{\infty}$ such that its projection $p_{*}=\pi(x_{*})$ satisfies
$\lVert\operatorname{grad}_{\mathcal{M}}f(p_{*})\rVert^{2}\leq\frac{(8C^{2}\epsilon^{\prime}/\tau_{\mathcal{M}}^{2}+2(2C+L_{0}\gamma_{k}(C+L)))\epsilon^{\prime}}{\gamma_{\min}(1-\frac{L_{0}}{2}\gamma_{\max})}\,.$
Remark 6.
Note that both Theorem 3 and Theorem 5 require the rather strong $L^{\infty}$-approximation assumption (7) on the Tweedie score $s$ and its Jacobian. The analysis under a weaker $L^{2}$-bound is out of score for this paper and left for future work.
6 Numerical Experiments and Applications
In this section we provide some numerical results for our proposed algorithms, namely the denoising landing flow (8) (more precisely, the discretized version (48)) and the denoising Riemannian gradient descent (12).
6.1 Optimization on orthogonal group $O(n)$
In this section we evaluate flow (8) on a synthetic example. In order to compare the error of our method to classical Riemannian optimization techniques, we consider distributions supported on manifolds that are the focus of study in Absil et al. (2008); Boumal (2023). Specifically we consider the orthogonal group manifold $\mathcal{M}=\operatorname{O}(n)\subseteq\mathbb{R}^{n\times n}$ with $\mu=\operatorname{Vol}_{\mathcal{M}}$ being the uniform volume measure and Brockett’s cost function Helmke & Moore (2012) defined by
$\min_{X\in\operatorname{O}(n)}f(X):=\operatorname{tr}(AXQX^{\top})\,$,
where $A,Q\in\mathbb{S}^{n\times n}$ are given. We consider the cases $n\in\{10,20\}$ and assume that we are given a set $\mathcal{D}_{\operatorname{train}}\subseteq\mathcal{M}$ of $N_{\operatorname{data}}=20000$ data points from $\mu$ and train the score function $s$ with denoising score matching (see Appendix C for diffusion models and Appendix G.1.1 for implementation details). In Figure 3 (left) (in Appendix G.1.2 for space reasons) we compare the evolution of the objective value of our approximation of (8) to the exact landing flow (10) for different noise levels $\sigma>0$ and dimension $n$. We observe that we can obtain objective values with cost lower than the best possible point in the training set and that the accuracy improves as $\sigma\to 0$.
6.2 Reference tracking via data-driven control
Problem definition: In this example we consider applying our method to the control of discrete-time dynamical systems on a finite horizon. Specifically we assume that we are given a discrete-time state-space system
$x_{k+1}=\bm{f}(x_{k},u_{k})\,,\quad y_{k}=\bm{g}(x_{k},u_{k})\,,\quad k=0,\ldots,N_{h}-1\,$, (13)
on a finite time horizon $N_{h}$ with state $x_{k}\in\mathbb{R}^{n_{x}}$, input $u_{k}\in\mathbb{R}^{n_{u}}$, output $y_{k}\in\mathbb{R}^{n_{y}}$ and a fixed initial state $x_{0}=0\in\mathbb{R}^{n_{x}}$. The task is to find inputs $\bm{u}=(u_{0},\ldots,u_{N_{h}-1})$ such that the corresponding outputs $\bm{y}=(y_{0},\ldots,y_{N_{h}})$ closely track a prespecified reference trajectory $\bm{r}=(r_{0},\ldots,r_{N_{h}})$ by solving the optimal control problem
$\min_{(\bm{u},\bm{y})\in\mathcal{M}_{\operatorname{IO}}}f(\bm{u},\bm{y})$ (14)
with tracking objective
$f(\bm{u},\bm{y})=\sum_{k=0}^{N_{\operatorname{h}}-1}u_{k}^{\top}Ru_{k}+(y_{k}-r_{k})^{\top}Q(y_{k}-r_{k})+(y_{N_{h}}-r_{N_{h}})^{\top}Q(y_{N_{h}}-r_{N_{h}})$ (15)
for positive-definite weight matrices $R\in\mathbb{S}^{n_{u}\times n_{u}}$ and $Q\in\mathbb{S}^{n_{y}\times n_{y}}$ and feasible input-output set
[…]
Under certain smoothness assumptions on the dynamics $\bm{f}$ and $\bm{g}$ the set $\mathcal{M}_{\operatorname{IO}}$ is a (non-compact) embedded smooth submanifold of $(\mathbb{R}^{n_{u}})^{N_{h}}\times(\mathbb{R}^{n_{y}})^{N_{h}+1}$. The problem (14) is ubiquitous in receding horizon control applications such as model predictive control (MPC) and used for e.g. autonomous driving Vu et al. (2021), motion planning Cohen et al. (2020), optimizing HVAC system energy efficiency Serale et al. (2018) and inventory control Kostić (2009). In many of these applications the dynamics (13) governing the system are not known explicitly. Instead, in data-driven control Dörfler (2023a; b); Markovsky et al. (2023) one assumes that (13) is given implicitly by a finite number of measured input-output trajectories
$\mathcal{D}_{\operatorname{train}}=\{(\bm{u}_{i},\bm{y}_{i})\mid i=1,\ldots,N_{\operatorname{data}}\}\subseteq\mathcal{M}_{\operatorname{IO}}\,$,
where the input $\bm{u}$ is persistently exciting Willems et al. (2005), e.g. given by (white) noise Ljung (1999). In particular the so-called system behavior manifold $\mathcal{M}_{\operatorname{IO}}$ Willems & Polderman (1997) is given by samples from a distribution $\mu$ on its in- and outputs and fits precisely into our framework of data-driven Riemannian optimization (1). We test our proposed denoising Riemannian gradient descent on two classical systems from the control domain: The discretized double pendulum system and the unicycle car model LaValle (2006) (see Appendix G.2.1 detailed information on the systems and the particular choice of $\bm{r}$, $R$ and $Q$ in (15)), each on a horizon of $N_{h}=100$. To apply our proposed methods, we train a diffusion model (see Appendix G.2.2 for implementation details) on the measured trajectories $\mathcal{D}_{\operatorname{train}}$ and solve (14) via the denoising Riemannian gradient descent to obtain a solution $(\bm{u}^{*},\bm{y}^{*})$. As initial values we take the trajectories from the training set that minimize the objective cost $f$, i.e. $(\bm{u}_{0},\bm{y}_{0})=\operatorname*{arg\,min}_{(\bm{u},\bm{y})\in\mathcal{D}_{\operatorname{train}}}f(\bm{u},\bm{y})$. Note that, as seen in Section 5, in general the final iterate will not exactly lie on the input-output manifold, i.e. $(\bm{u}^{*},\bm{y}^{*})\notin\mathcal{M}_{\operatorname{IO}}$. To account for this deviation we back-test our generated input trajectory $\bm{u}^{*}$ by implementing it on the true underlying system (13) to obtain the real output $\bm{y}^{\operatorname{true}}$.
Results and discussion: In Figure 4 (in Appendix G.2.3) we depict the evolution of the objective value w.r.t. iteration count and in Figure 2 we depict the final optimizing trajectories $\bm{y}^{*}$ and $\bm{y}^{\operatorname{true}}$. We can observe that the error $\lVert\bm{y}^{*}-\bm{y}^{\operatorname{true}}\rVert$ is small, which shows that $(\bm{u}^{*},\bm{y}^{*})$ is close to the true system behavior $\mathcal{M}_{\operatorname{IO}}$. Moreover, we can see that the trajectory $\bm{y}^{\operatorname{true}}$ tracks the corresponding reference $\bm{r}$ much better than the train set minimum $\bm{y}_{0}$, which shows a generalization capability of our diffusion model. We from Figure 4 (left) that the current objective can depart from the true objective significantly. This is due to the iterates deviating from $\mathcal{M}_{\operatorname{IO}}$. In this example the algorithm (DRGD) recovers and we have found it to be robust w.r.t. moderate deviations from the manifold. Note that we have set a iteration budget of $N_{\operatorname{iter}}=3000$ and $N_{\operatorname{iter}}=2500$, respectively, while the objective is still decreasing. Accelerating the convergence of DRGD is left for future work.
Figure 2: Denoising Riemannian gradient descent: Angle of the first pendulum (left) and unicycle car position (right) with the optimized output trajectory $\bm{y}^{*}$ (blue, dashed), the true system trajectory $\bm{y}^{\operatorname{true}}$ (orange), the initial trajectory $\bm{y}_{0}$ (green, dotted) and the reference trajectory $\bm{r}$ (red)
7 Conclusion and Future Work
In this paper, we show that the denoising score and its Jacobian allow to perform manifold operations such as the closest-point and tangent space projection without the explicit knowledge of the manifold. We then propose a landing flow for the corresponding manifold-constrained Riemannian optimization problem and show that its limit points approximate critical points of the original problem. Moreover, we investigate the approximate version of the Riemannian gradient descent and provide a bound on the average-gradient-norm of its iterates, which converges to zero as the manifold operations become more exact. We apply both algorithms on known manifolds and to finite-horizon reference tracking in the domain of data-driven control. Future work will consist of deriving error bounds for this flow when the denoising score is trained with a non-zero $L^{2}$-error as well as the study of more sophisticated classical Riemannian optimization algorithms such as Newton and trust region methods when using the approximate manifold operations with the trained score to accelerate convergence.
Appendix A More on manifolds and distance functions
In addition to the notation introducted in Section 2.1, we note that $\tau_{\mathcal{M}}$ is also the largest $\tau\geq 0$ such that the map $\{(p,v)\in\operatorname{N}\mathcal{M}\mid\lVert v\rVert<\tau\}\to\mathbb{R}^{d}:(p,v)\mapsto p+v$ is a diffeomorphism. By a tubular neighborhood of radius $\tau\in(0,\tau_{\mathcal{M}}]$ we mean a set of the form $\mathcal{T}(\tau)=\{p+v\mid p\in\mathcal{M}\,,v\in\operatorname{N}_{p}\mathcal{M}\,,\lVert v\rVert<\tau\}$. Moreover, for $x\in\mathbb{R}^{d}$ let $\operatorname{dist}_{\mathcal{M}}(x)=\inf_{p\in\mathcal{M}}\lVert x-p\rVert$ denote the distance function so that $\operatorname{d}(x)=\frac{1}{2}\operatorname{dist}_{\mathcal{M}}(x)^{2}$. The second fundamental form of $\mathcal{M}$ at a point $p\in\mathcal{M}$ will be denoted by $\mathrm{I\!I}_{p}$ and is a symmetric bilinear map $\mathrm{I\!I}_{p}:\operatorname{T}_{p}\mathcal{M}\times\operatorname{T}_{p}\mathcal{M}\to\operatorname{N}_{p}\mathcal{M}$ intrinsic to the manifold $\mathcal{M}$. Fixing some $u\in\operatorname{N}_{p}\mathcal{M}$ we also define the directed second fundamental form $\mathrm{I\!I}_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\times\operatorname{T}_{p}\mathcal{M}\to\mathbb{R}:(v,w)\mapsto\left\langle\mathrm{I\!I}_{p}(v,w),u\right\rangle_{\operatorname{N}_{p}\mathcal{M}}$. The Weingarten map $S_{p}^{u}$ at a point $p\in\mathcal{M}$ in the direction $u\in\operatorname{N}_{p}\mathcal{M}$ is defined as the unique self-adjoint linear operator $S_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\to\operatorname{T}_{p}\mathcal{M}$ such that $\left\langle w,S_{p}^{u}(v)\right\rangle_{\operatorname{T}_{p}\mathcal{M}}=\mathrm{I\!I}_{p}^{u}(v,w)$ for all $v,w\in\operatorname{T}_{p}\mathcal{M}$. A useful operator that has been studied in Abatzoglou (1978); Breiding & Vannieuwenhoven (2021); Alvarez et al. (2004) is
$H_{x}=I_{T_{\pi(x)}\mathcal{M}}+S_{\pi(x)}^{\pi(x)-x}:\operatorname{T}_{\pi(x)}\mathcal{M}\to\operatorname{T}_{\pi(x)}\mathcal{M}\,.$ (16)
In Breiding & Vannieuwenhoven (2021) it has been shown (see Lemma 8) that $H_{x}$ is invertible on $\mathcal{T}$. Now we have the following useful identities that hold for $x\in\mathcal{T}$
$\pi^{\prime}(x)$ $=\operatorname{I}_{\operatorname{T}_{\pi(x)}\mathcal{M}}H_{x}^{-1}\operatorname{P}_{\operatorname{T}_{\pi(x)}\mathcal{M}}$
$\nabla\operatorname{d}(x)$ $=x-\pi(x)$
$\nabla^{2}\operatorname{d}(x)$ $=I-\pi^{\prime}(x)=I-\operatorname{I}_{\operatorname{T}_{\pi(x)}\mathcal{M}}H_{x}^{-1}\operatorname{P}_{\operatorname{T}_{\pi(x)}\mathcal{M}}\,.$
[…]
Lemma 7 (Lemma A.1 in Divol (2022)).
[…]
  1. (i)
    For all $r\leq\min\{\tau_{\mathcal{M}},M\}/4$ it holds that $B_{r}(p)\cap\mathcal{M}\subseteq\psi_{p}(B_{r}^{\operatorname{T}_{p}\mathcal{M}}(0))\subseteq B_{8r/7}(p)\cap\mathcal{M}$. For $z\in B_{\min\{\tau_{\mathcal{M}},M\}/4}^{\operatorname{T}_{p}\mathcal{M}}(0)$ it holds that $\lVert z\rVert\leq\lVert\psi_{p}(z)-p\rVert\leq 8\lVert z\rVert/7$.
  2. (ii)
    There exists a map $W_{p}:B_{\min\{\tau_{\mathcal{M}},M\}/4}^{\operatorname{T}_{p}\mathcal{M}}(0)\to\operatorname{N}_{p}\mathcal{M}$ with $W_{p}^{\prime}(0)=0$ and such that $\psi_{p}(z)=p+z+W_{p}(z)$ and $\lVert W_{p}(z)\rVert\leq M\lVert z\rVert^{2}$ for all $z\in B_{\min\{\tau_{\mathcal{M}},M\}/4}^{\operatorname{T}_{p}\mathcal{M}}(0)$.
  3. (iii)
    For $G_{\psi_{p}}:B_{\min\{\tau_{\mathcal{M}},M\}/4}^{\operatorname{T}_{p}\mathcal{M}}(0)\to\mathbb{R}:z\mapsto\sqrt{\det\psi_{p}^{\prime}(z)^{\top}\psi_{p}^{\prime}(z)}$ it holds that $G_{\psi_{p}}(0)=1$ and $\nabla G_{\psi_{p}}(0)=0$.
Note that for the graph chart $\psi_{p}$ we always have
$\psi_{p}^{\prime}(0)=\operatorname{I}_{\operatorname{T}_{p}\mathcal{M}}\,,\quad\psi_{p}^{\prime\prime}(0)[\cdot,\cdot]=W_{p}^{\prime\prime}(0)[\cdot,\cdot]=\mathrm{I\!I}_{p}(\cdot,\cdot)\,$, (17)
and hence $\lVert\psi_{p}^{\prime}(0)\rVert\leq 1$ and $\lVert\psi_{p}^{\prime\prime}(0)\rVert=\lVert\mathrm{I\!I}_{p}\rVert\leq 1/\tau_{\mathcal{M}}$.
A.1 Properties of (16)
We study the invertibility and boundedness of the operator (16). For this purpose, let us recall first the definition of the (normalized) curvature radius of $\mathcal{M}$ at $p$ in the direction of $u\in\operatorname{N}_{p}\mathcal{M}$:
$\frac{1}{\rho(p,u)}=\max_{\begin{subarray}{c}v\in\operatorname{T}_{p}\mathcal{M}\\ \mathrm{I\!I}_{p}^{u}(v,v)\geq 0\end{subarray}}\frac{\mathrm{I\!I}_{p}^{u/\lVert u\rVert}(v,v)}{\lVert v\rVert^{2}}=\max(\operatorname{eig}(S_{p}^{u/\lVert u\rVert})\cup\{0\})\,.$
If $S_{p}^{u}$ has only non-positive eigenvalues, then $\mathcal{M}$ is curved away from the unit vector $u$ and thus the curvature radius is infinite. Moreover, we define the (normalized) curvature of $\mathcal{M}$ to be
$\kappa_{p}^{\mathcal{M}}(u)=\max\left\lvert\operatorname{eig}(S_{p}^{u/\lVert u\rVert})\right\rvert\text{\ for\ }u\in\operatorname{N}_{p}\mathcal{M}\,$,
and the maximal curvature by
$\kappa_{\mathcal{M}}=\max_{(p,u)\in\operatorname{N}\mathcal{M}}\kappa_{p}^{\mathcal{M}}(u)\,.$ (18)
We have the following
Lemma 8.
The operator (16) is invertible on $\mathcal{T}(\tau_{\mathcal{M}})$. Moreover, (16) satisfies [^1] In (Breiding & Vannieuwenhoven, 2021, Theorem 4.3) the quantity $\lVert H_{x}^{-1}\rVert$ has been shown to equal the condition number of a certain critical point problem associated with $\mathcal{M}$.
$\lVert P_{0}(x)\rVert=\lVert H_{x}^{-1}\rVert=\left(1-\frac{\lVert\pi(x)-x\rVert}{\rho(\pi(x),x-\pi(x))}\right)^{-1}\leq\frac{1}{1-\lVert x-\pi(x)\rVert\kappa_{\mathcal{M}}}\leq\frac{1}{1-\lVert x-\pi(x)\rVert/\tau_{\mathcal{M}}}\,$,
and if $\psi$ is a local parametrization of $\mathcal{M}$ with $\psi(0)=p$, then
$P_{0}(x)=\psi^{\prime}(0)\left(\psi^{\prime}(0)^{\top}\psi^{\prime}(0)+\sum_{i=1}^{d}(p-x)_{i}\nabla^{2}\psi_{i}(0)\right)^{-1}\psi^{\prime}(0)^{\top}\,.$
In particular, for any $\tau\in(0,\tau_{\mathcal{M}})$, it holds
$\sup_{x\in\mathcal{T}(\tau)}\lVert H_{x}^{-1}\rVert<\infty\,.$
Proof.
In (Breiding & Vannieuwenhoven, 2021, Lemma A.2) the following condition has been established: Let $\mathcal{S}=\{(a,p)\in\mathbb{R}^{n}\times\mathcal{M}\mid a-p\in\operatorname{N}_{p}\mathcal{M}\}$. Then $\mathcal{S}$ is diffeomorphic to the normal bundle $\operatorname{N}\mathcal{M}$ via the diffeomorphism $\Phi:\operatorname{N}\mathcal{M}\to\mathcal{S}:(v,p)\mapsto(p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v),p)$. Consider the operator $\Pi:\mathcal{S}\to\mathbb{R}^{n}:(a,p)\mapsto a$ and the domain where its differential is invertible $\mathcal{W}=\{(a,p)\in\mathcal{S}\mid\Pi^{\prime}(a,p):\operatorname{T}_{(a,p)}\mathcal{S}\to\mathbb{R}^{n}\text{\ invertible}\}$. Then $H_{x}$ is invertible iff $(x,\pi(x))\in\mathcal{W}$. But $\Pi\circ\Phi:\operatorname{N}\mathcal{M}\mapsto\mathbb{R}^{n}:(v,p)\mapsto p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v)$ being a diffeomorphism (and thus having an inveritble differential) is precisely the condition in the definition of the tubular neighborhood $\mathcal{T}$. The formulas for $P_{0}(x)$ in local coordinates as well as $\lVert P_{0}(x)\rVert$ are given in (Abatzoglou, 1978, Theorem 4.1, Corollary 4.1). To see that $P_{0}(x)$ is bounded on $\mathcal{T}(\tau)$ for any $\tau\in(0,\tau_{\mathcal{M}})$ it sufficies to note that in the tubular neighborhood $\mathcal{T}=\mathcal{T}(\tau_{\mathcal{M}})$ we always have $\lVert\pi(x)-x\rVert<\rho(\pi(x),x-\pi(x))$ and that $\overline{\mathcal{T}(\tau)}$ is a compact subset thereof. The second inequality follows from $1/\rho(p,u)\leq\kappa_{\mathcal{M}}$. ∎
Now let us derive bounds for the quantity $\lVert P_{0}(\pi(x))-P_{0}(x)\rVert$ when $x\in\mathcal{T}(\tau_{\mathcal{M}})$.
Lemma 9.
If $x\in\mathcal{T}(\tau_{\mathcal{M}})$, then
$\lVert P_{0}(\pi(x))-P_{0}(x)\rVert$ $\leq\kappa_{\pi(x)}^{\mathcal{M}}(x-\pi(x))\left(1-\frac{\lVert\pi(x)-x\rVert}{\rho(\pi(x),x-\pi(x))}\right)^{-1}\lVert x-\pi(x)\rVert\,$,
$\leq\frac{\lVert x-\pi(x)\rVert\kappa_{\mathcal{M}}}{1-\lVert x-\pi(x)\rVert\kappa_{\mathcal{M}}}\leq\frac{\lVert x-\pi(x)\rVert/\tau_{\mathcal{M}}}{1-\lVert x-\pi(x)\rVert/\tau_{\mathcal{M}}}\,.$
Proof.
We clearly have for $x\in\mathcal{T}$
$P_{0}(\pi(x))-P_{0}(x)=\operatorname{I}_{\operatorname{T}_{\pi(x)}\mathcal{M}}(I-H_{x}^{-1})\operatorname{P}_{\operatorname{T}_{\pi(x)}\mathcal{M}}\,.$
Moreover, $(I-H_{x}^{-1}):\operatorname{T}_{\pi(x)}\mathcal{M}\to\operatorname{T}_{\pi(x)}\mathcal{M}$ is symmetric with eigenvalues
$\operatorname{eig}(I-H_{x}^{-1})=\left\{\frac{\zeta}{1+\zeta}\mid\zeta\in\operatorname{eig}(S_{\pi(x)}^{\pi(x)-x})\right\}=\left\{\frac{-\lVert\pi(x)-x\rVert\zeta}{1-\lVert\pi(x)-x\rVert\zeta}\mid\zeta\in\operatorname{eig}(S_{\pi(x)}^{u})\right\}\,$,
where $u=\frac{x-\pi(x)}{\lVert x-\pi(x)\rVert}$. Thus
$\lVert P_{0}(\pi(x))-P_{0}(x)\rVert$ $\leq\max_{\zeta\in\operatorname{eig}(S_{\pi(x)}^{u})}\left\lvert\frac{\zeta}{1-\lVert x-\pi(x)\rVert\zeta}\right\rvert\lVert x-\pi(x)\rVert$
$\leq\lVert S_{\pi(x)}^{u}\rVert\left(1-\frac{\lVert\pi(x)-x\rVert}{\rho(\pi(x),x-\pi(x))}\right)^{-1}\lVert x-\pi(x)\rVert\,$,
which shows the first inequality. The second and third inequalities follow from $1/\rho(p,u)\leq\kappa_{\mathcal{M}}\leq 1/\tau_{\mathcal{M}}$ . ∎
[…]
Lemma 10.
[…]
$\sup_{x\in\mathcal{T}(\tau)}\lVert P_{0}^{\prime}(x)\rVert\leq(\frac{1}{1-\tau/\tau_{\mathcal{M}}})^{2}\left((\frac{3}{\tau_{\mathcal{M}}}+\tau M)(\frac{1}{1-\tau/\tau_{\mathcal{M}}})^{2}+\frac{2}{\tau_{\mathcal{M}}}\right)$
Proof.
Let $x\in\mathcal{T}(\tau)$ where $\tau\in(0,\tau_{\mathcal{M}})$. First let us relate the quantity $\pi(x)$ and a fixed chart $\psi:\mathcal{V}\to\mathcal{U}$ with $\pi(x)\in\mathcal{U}$ for all $x\in\mathcal{W}$ for some (small enough) open set $\mathcal{W}\subseteq\mathbb{R}^{d}$. The map $h_{\psi}(z)=\frac{1}{2}\lVert x-\psi(z)\rVert^{2}$ attains its minimum in some $z_{\psi}(x)\in\mathcal{V}$ and if $F^{\psi}(z,x)=h_{\psi}^{\prime}(z)=\psi^{\prime}(z)(\psi(z)-x)$, then $F^{\psi}(z_{\psi}(x),x)=0$. Moreover, we have $\pi(x)=\psi(z_{\psi}(x))$ and hence for $w\in\mathbb{R}^{d}$
$P_{0}^{\prime}(x)[w,w]=\pi^{\prime\prime}(x)[w,w]=\psi^{\prime\prime}(z_{\psi}(x))[z_{\psi}^{\prime}(x)[w],z_{\psi}^{\prime}(x)[w]]+\psi^{\prime}(z_{\psi}(x))[z_{\psi}^{\prime\prime}(x)[w,w]]\,$,
and hence
$\lVert P_{0}^{\prime}(x)\rVert\leq\lVert\psi^{\prime\prime}(z_{\psi}(x))\rVert\lVert z_{\psi}^{\prime}(x)\rVert^{2}+\lVert\psi^{\prime}(z_{\psi}(x))\rVert\lVert z_{\psi}^{\prime\prime}(x)\rVert$
By the implicit function theorem we can bound the derivatives of $z_{\psi}$ in terms of derivatives of $F^{\psi}$. Indeed, we have (here again $w\in\mathbb{R}^{d}$ and $v\in\mathbb{R}^{k}$ are place-holder vectors to express the differentials)
[…]
with
$F_{z}^{\psi}(z,x)[v,v]$ $=\left\langle\psi^{\prime}(z)[v],\psi^{\prime}(z)[v]\right\rangle+\left\langle\psi(z)-x,\psi^{\prime\prime}(z)[v,v]\right\rangle\,$,
$F_{x}^{\psi}(z,x)[w,v]$ $=\left\langle\psi^{\prime}(z)[v],w\right\rangle\,$,
$F_{zz}^{\psi}(z,x)[v,v,v]$ $=3\left\langle\psi^{\prime\prime}(z)[v,v],\psi^{\prime}(z)[v]\right\rangle+\left\langle\psi(z)-x,\psi^{\prime\prime\prime}(z)[v,v,v]\right\rangle\,$,
$F_{xz}^{\psi}(z,x)[w,v,v]$ $=\left\langle\psi^{\prime\prime}(z)[v,v],w\right\rangle\,$,
$F_{xx}^{\psi}(z,x)[w,v,v]$ $=0\,.$
Now we set $\psi=\psi_{p}$ for the graph chart defined on $\mathcal{V}=B_{\min\{\tau,M\}/4}^{\operatorname{T}_{p}\mathcal{M}}(0)$, where $p=\pi(x)$. In this case $z=z_{\psi}(x)=0$ and $\lVert F_{z}^{\psi}(0,x)^{-1}\rVert=\lVert P_{0}(x)\rVert\leq(1-\tau/\tau_{\mathcal{M}})^{-1}$ by Lemma 8. Solving for $z_{\psi}^{\prime}(x)$ and $z_{\psi}^{\prime\prime}(x)$ in (19), using (17) with $\lVert\mathrm{I\!I}_{p}\rVert\leq 1/\tau_{\mathcal{M}}$ and taking norms yields then
$\lVert z_{\psi}^{\prime}(x)\rVert$ $\leq\frac{1}{1-\tau/\tau_{\mathcal{M}}}\,$,
$\lVert z_{\psi}^{\prime\prime}(x)\rVert$ $\leq\frac{1}{1-\tau/\tau_{\mathcal{M}}}\left((\frac{3}{\tau_{\mathcal{M}}}+\tau M)\lVert z_{\psi}^{\prime}(x)\rVert^{3}+\frac{1}{\tau_{\mathcal{M}}}\lVert z_{\psi}^{\prime}(x)\rVert\right)\,.$
Plugging this back into the upper bound of $\lVert P_{0}^{\prime}(x)\rVert$ and using once again (17)finishes the proof ∎
The next lemma establishes a lower bound on the distance between a point $x$ in a tubular neighborhood and any other point that is sufficiently away from $p=\pi(x)$ uniformly in $x$.
Lemma 11.
Let $\tau\in(0,\tau_{\mathcal{M}})$ and let $\eta>0$. Then
$\inf_{x\in\mathcal{T}(\tau)}\frac{1}{2}\operatorname{dist}_{\mathcal{M}\setminus B_{\eta}(\pi(x))}(x)^{2}-\frac{1}{2}\operatorname{dist}_{\mathcal{M}}(x)^{2}\geq\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}\eta^{2}>0\,.$
Proof.
Let $\tau^{\prime}=\frac{\tau+\tau_{\mathcal{M}}}{2}\in(0,\tau_{\mathcal{M}})$. Take $x\in\mathcal{T}(\tau)$ and set $p=\pi(x)$ and $y=p+\tau^{\prime}\widehat{n}$ with $\widehat{n}=\frac{x-p}{\lVert x-p\rVert}$. Then $y\in\mathcal{T}(\tau_{\mathcal{M}})$ with $\operatorname{dist}_{\mathcal{M}}(y)=\tau^{\prime}$ and thus $\overline{B}_{\tau^{\prime}}(y)\cap\mathcal{M}=\{p\}$. Then, since $\mathcal{M}\setminus B_{\eta}(p)\subseteq\mathbb{R}^{d}\setminus(\overline{B}_{\tau^{\prime}}(y)\cup B_{\eta}(p))$, it is sufficient to show
$\inf_{q\in\mathbb{R}^{d}\setminus(\overline{B}_{\tau^{\prime}}(y)\cup B_{\eta}(p))}\lVert x-q\rVert^{2}-\lVert x-p\rVert^{2}\geq\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}\eta^{2}\,.$
To see this, note that $q\in\mathbb{R}^{d}\setminus(\overline{B}_{\tau^{\prime}}(y)\cup B_{\eta}(p))$ implies $\lVert y-q\rVert\geq\tau^{\prime}=\lVert y-p\rVert$ and $\lVert q-p\rVert\geq\eta$, as well as $2\left\langle p-q,y-q\right\rangle\geq\lVert p-q\rVert^{2}$. Then, abbreviating $t=\frac{\lVert x-p\rVert}{\lVert y-p\rVert}\in(0,1)$, we have x=ty+(1t)p=p+t(yp)x=ty+(1-t)p=p+t(y-p) and
$\lVert x-q\rVert^{2}-\lVert x-p\rVert^{2}$ $=\lVert ty+(1-t)p-q\rVert^{2}-t^{2}\lVert y-p\rVert^{2}$
$\geq\lVert t(y-q)+(1-t)(p-q)\rVert^{2}-t^{2}\lVert y-q\rVert^{2}$
$=(1-t)^{2}\lVert p-q\rVert^{2}+2t(1-t)\left\langle p-q,y-q\right\rangle$
$\geq(1-t)\lVert p-q\rVert^{2}$
$\geq\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}\eta^{2}\,.$
A.2 Densities on manifolds
Given a manifold $\mathcal{M}$ as in Section A and any chart $\psi:\mathcal{V}\to\mathcal{U}\subseteq\mathcal{M}$, the volume measure $\operatorname{Vol}_{\mathcal{M}}$ is uniquely defined by
$\operatorname{Vol}_{\mathcal{M}}(E)=\int_{\psi^{-1}(E)}G_{\psi}(z)\,\text{d}z\text{\ \ for\ \ }E\subseteq\mathcal{U}\text{\ Borel measurable}.$
Here $G_{\psi}(z)=\sqrt{\det\psi^{\prime}(z)^{\top}\psi^{\prime}(z)}$ and $\operatorname{Vol}_{\mathcal{M}}$ is independent of the chart $\psi$. If $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is absolutely continuous w.r.t. $\operatorname{Vol}_{\mathcal{M}}$, then for the density $\mu(y):=\frac{\operatorname{d}\mu}{\operatorname{d}\operatorname{Vol}_{\mathcal{M}}}(y)$ on $\mathcal{M}$ we have the local representations
$\mu(\psi(z))=\frac{\operatorname{d}\lambda}{\operatorname{d}(\psi^{-1}\#\operatorname{Vol}_{\mathcal{M}})}(z)\,,\quad\lambda(z)=G_{\psi}(z)\mu(\psi(z))$
[…]
$\lambda(0)=\mu(p)\,,\quad\nabla\lambda(0)=\operatorname{grad}_{\mathcal{M}}\mu(p)\in\operatorname{T}_{p}\mathcal{M}\cong\mathbb{R}^{k}\,.$ (20)
Appendix B More on the Stein score function
In this section we derive the representations (4). For $\sigma\geq 0$ let $\mathcal{N}(0,\sigma^{2}I_{d})$ denote the Gaussian distribution with mean zero and variance $\sigma^{2}$ and for $\sigma>0$ the Gaussian (heat) kernel in the ambient space $\mathbb{R}^{d}$ by
$\varphi_{\sigma}(x)=\frac{1}{Z_{\sigma}}e^{-\lVert x\rVert^{2}/2\sigma^{2}}\,,\quad Z_{\sigma}=\frac{1}{(2\pi\sigma^{2})^{d/2}}\,.$
Then $p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu$ has for $\sigma>0$ a fully supported $C^{\infty}$-density, denoted by $p_{\sigma}$ as well, given by
$p_{\sigma}(x)=\int_{\mathcal{M}}\varphi_{\sigma}(x-y)\,\text{d}\mu(y)\,,\quad x\in\mathbb{R}^{d}\,.$
Let $\nu_{x,\sigma}$ be the posterior of observing $x$ under $p_{\sigma}$ with prior $\mu$, i.e.
$\nu_{x,\sigma}(E)=\frac{1}{p_{\sigma}(x)}\int_{E}\varphi_{\sigma}(x-y)\,\text{d}\mu(y)\,,\quad E\subseteq\mathbb{R}^{d}\text{\ Borel}\,.$
We have the following representation
Lemma 12.
For each $\sigma>0$ representations (4) hold.
Proof.
Clearly
$\nabla p_{\sigma}(x)$ $=-\frac{1}{\sigma^{2}}\int_{\mathcal{M}}(x-y)\varphi_{\sigma}(x-y)\,\text{d}\mu(y)\,,\quad x\in\mathbb{R}^{d}\,$,
$\nabla^{2}p_{\sigma}(x)$ $=-\frac{1}{\sigma^{2}}\left(p_{\sigma}(x)I-\frac{1}{\sigma^{2}}\int_{\mathcal{M}}(x-y)(x-y)^{\top}\varphi_{\sigma}(x-y)\,\text{d}\mu(y)\right)\,,\quad x\in\mathbb{R}^{d}\,.$
and hence
$\nabla\log p_{\sigma}(x)=\frac{\nabla p_{\sigma}(x)}{p_{\sigma}(x)}=-\frac{1}{\sigma^{2}}(x-\mathbb{E}\nu_{x,\sigma})\,$,
which yields the first representation in (4). Further
$\frac{\nabla^{2}p_{\sigma}(x)}{p_{\sigma}(x)}=-\frac{1}{\sigma^{2}}\left(I-\frac{1}{\sigma^{2}}\int_{\mathcal{M}}(x-y)(x-y)^{\top}\,\text{d}\nu_{x,\sigma}(y)\right)$
and hence
$\nabla^{2}\log p_{\sigma}(x)$ $=\frac{\nabla^{2}p_{\sigma}(x)}{p_{\sigma}(x)}-\frac{\nabla p_{\sigma}(x)\nabla p_{\sigma}(x)^{\top}}{p_{\sigma}(x)^{2}}$
$=-\frac{1}{\sigma^{2}}\left(I-\frac{1}{\sigma^{2}}\int_{\mathcal{M}}(x-y)(x-y)^{\top}\,\text{d}\nu_{x,\sigma}(y)\right)$
$\qquad-\frac{1}{\sigma^{4}}(x-\mathbb{E}\nu_{x,\sigma})(x-\mathbb{E}\nu_{x,\sigma})^{\top}$
$=-\frac{1}{\sigma^{2}}I+\frac{1}{\sigma^{4}}\left(\int_{\mathcal{M}}yy^{\top}\,\text{d}\nu_{x,\sigma}(y)-(\mathbb{E}\nu_{x,\sigma})(\mathbb{E}\nu_{x,\sigma})^{\top}\right)$
$=-\frac{1}{\sigma^{2}}I+\frac{1}{\sigma^{4}}\operatorname{Cov}(\nu_{x,\sigma})\,$,
which yields the second representation in (4). ∎
Appendix C Variance-exploding diffusion models
In the variance-exploding scheme of score-based diffusion models Song et al. (2020); Tang & Zhao (2025), one considers the following stochastic differential equation on the finite interval $[0,T]$:
$\operatorname{d}X_{t}=\sqrt{2t}\operatorname{d}W_{t}\,,\quad t\in[0,T]\,,\quad X_{0}\sim\mu\,.$ (21)
The solution $X_{t}$ of this SDE is distributed according to $X_{t}\sim p_{\sigma(t)}$, where $\sigma^{2}(t)=t^{2}$. To sample from $\mu$ one exploits the fact that the reverse SDE
$\operatorname{d}\overline{X}_{t}=2(T-t)\nabla\log p_{\sigma(T-t)}(\overline{X}_{t})\operatorname{d}t+\sqrt{2(T-t)}\operatorname{d}W_{t}\,,\quad t\in[0,T]\,,\quad\overline{X}_{0}\sim p_{\sigma(T)}\,$, (22)
satisfies $\overline{X}_{t}\sim p_{\sigma(T-t)}$ and in particular $\overline{X}_{0}\sim p_{0}=\mu$. Here the initial distribution is approximated by $p_{\sigma(T)}\approx\mathcal{N}(0,\sigma(T)^{2}I)$ and unknown score $\nabla\log p_{\sigma(T-t)}$ is learned by minimizing the conditional score matching loss defined as
$L_{\operatorname{CSM}}(s)=\mathbb{E}_{t\sim\operatorname{Unif}[0,T]}\mathbb{E}_{x_{0}\sim\mu}\mathbb{E}_{x\sim p_{\sigma(t)}(\cdot\mid x_{0})}\sigma(t)^{2}\lVert s_{\sigma(t)}(x)-\nabla\log p_{\sigma(t)}(x\mid x_{0})\rVert^{2}\,$, (23)
and which admits the unique minimizer $s_{\sigma}(x)=\nabla\log p_{\sigma}(x)$. The loss $L_{\operatorname{CSM}}$ can be evaluated because $\nabla\log p_{\sigma}(x\mid x_{0})=-\frac{1}{\sigma^{2}}(x-x_{0})$ is known explicitly.
Appendix D Proof of Theorem 1
D.1 Nonasymptotic Laplace method
In this section we derive some non-asymptotic error estimates for the Laplace method Hashorva et al. (2015); Hwang (1980), which concerns itself with the asymptotic of integrals of the form
$\int_{\mathcal{V}}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z\quad(\sigma\to 0)$
for some functions $f,h:\mathcal{V}\to\mathbb{R}$ from an open set $\mathcal{V}\subseteq\mathbb{R}^{k}$, where $h$ is non-negative and attains a unique minimum in, say, $z=0\in\mathcal{V}$. While there have been results providing such an error estimation to the first order expansion Inglot & Majerski (2014); Majerski (2015); Lapinski (2019), we need an error estimate up to the second order expansion, as provided in Theorem 14. For the sake of completeness we also include the statement and proof for the first order expansion in Theorem 13. Before we state the results, we need to introduce some notation and conventions. For any $f\in C^{0}(\mathcal{V})$ and $h\in C^{2}(\mathcal{V})$, respectively we write
$\zeta_{f}(z)$ $=f(z)-f(0)\overset{z\to 0}{=}o(1)$
$\chi_{h}(z)$ $=h(z)-h(0)-\nabla h(0)^{\top}z-\frac{1}{2}z^{\top}\nabla^{2}h(0)z\overset{z\to 0}{=}o(\lVert z\rVert^{2})$
for its first and second order remainder terms and abbreviate
$\zeta_{f}^{\sigma}(w)=\zeta_{f}(\sigma w)\text{\ and\ }\chi_{h}^{\sigma}(w)=\frac{1}{\sigma^{2}}\chi_{h}(\sigma w)\text{\ for\ }\sigma>0\,.$
Let us note the following: If additionally $f\in C^{1}(\mathcal{V})$ and $g\in C^{3}(\mathcal{V})$, then $\zeta_{f}(z)=O(\lVert z\rVert)$ and $\chi_{g}(z)=O(\lVert z\rVert^{3})$ for $z\to 0$, i.e. there exist some $\eta>0$ and $C>0$ such that $B_{\eta}(0)\subseteq\mathcal{V}$ and
$\lvert\zeta_{f}(z)\rvert\leq C\lVert z\rVert\text{\ and\ }\lvert\chi_{h}(z)\rvert\leq C\lVert z\rVert^{3}\text{\ for all\ }z\in B_{\eta}(0)\,.$ (24)
In particular for $\sigma>0$ and $\beta:[0,\infty)\to[0,\infty)$ it holds
$\sup_{\lVert w\rVert\leq\min\{\eta/\sigma,\beta(\sigma)\}}\lvert\zeta_{f}^{\sigma}(w)\rvert\leq C\sigma\beta(\sigma)\text{\ and\ }\sup_{\lVert w\rVert\leq\min\{\eta/\sigma,\beta(\sigma)\}}\lvert\chi_{g}^{\sigma}(w)\rvert\leq C\sigma\beta(\sigma)^{3}\,.$ (25)
For a non-negative function $h$ with a global minimum at $z=0$ we define
$\gamma_{h}(\delta)=\inf_{\begin{subarray}{c}z\in\mathcal{V}\\ \lVert z\rVert\geq\delta\end{subarray}}h(z)-h(0)\,.$
We say that $h$ satisfies a local quadratic growth condition in $B_{\eta}(0)\subseteq\mathcal{V}$ if there exists a $c>0$ with
$h(z)-h(0)\geq c\lVert z\rVert^{2}\text{\ for all\ }z\in B_{\eta}(0)\,$, (26)
and that $h$ admits minimum separation outside of $B_{\eta}(0)$ in $\mathcal{V}$ if there exists a $\Delta>0$ with
$h(z)-h(0)\geq\Delta>0\text{\ for all\ }z\in\mathcal{V}\setminus B_{\eta}(0)\,.$ (27)
Clearly, if $h$ satisfies (24), then $h$ also satisfies (26) (for a potentially smaller $\eta$). On the other hand, the global assumption (27) cannot be inferred from properties of $h$ around $z=0$ alone. Together, (26) and (27) imply that $\gamma_{h}$ can be bounded below by
$\gamma_{h}(\delta)\geq\min\{c\delta^{2},\Delta\}\,$, (28)
which is crucial to obtain a quantitative bound on the convergence of the Laplace method. While we state the following results in this full generality, in our application we have $B_{\eta}(0)=\mathcal{V}$ and thus (27) is not needed, with (24) and (26) holding globally on $\mathcal{V}$ and implying that
$\gamma_{h}(\delta)\geq c\delta^{2}\,.$ (29)
Theorem 13 (First Order Laplace method).
Let $h\in C^{3}(\mathcal{V})$ with $\nabla h(0)=0$ and $f\in C^{1}(\mathcal{V})$ for some open $B_{r}(0)\subseteq\mathcal{V}\subseteq\mathbb{R}^{k}$. Let $\eta\in(0,r]$ and $C,c,\Delta>0$ be such that (24), (26) and (27) hold. Then for all $\sigma\in(0,\overline{\sigma})$ with $\overline{\sigma}=\min\{\eta^{2},(\log(2)/(2C))^{2}\}$ with it holds
$\left\lvert e^{\frac{1}{\sigma^{2}}h(0)}\frac{\sqrt{\det\Sigma}}{Z_{\sigma}}\int_{\mathcal{V}}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z-f(0)\right\rvert\leq E_{1}(\sigma;f,h,r)$ (30)
with $\Sigma=\nabla^{2}h(0)$ and $E_{1}(\sigma;f,h,r)=O(\sigma\lvert\log\sigma\rvert^{3})$ for $\sigma\to 0$ given in (31).
Proof.
Without loss of generality we can assume $h(0)=0$. For brevity we also write $Z_{\sigma}^{\Sigma}=\frac{Z_{\sigma}}{\sqrt{\det\Sigma}}$ for the normalizing constant. Let $\alpha:[0,\infty)\to[0,\infty)$ be any function and denote $B(\sigma)=B_{\alpha(\sigma)}(0)$. We can split
$\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\mathcal{V}}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z=\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\mathcal{V}\setminus B(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z+\frac{1}{Z_{\sigma}^{\Sigma}}\int_{B(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z\,.$
We will craft $\alpha:[0,\infty)\to[0,\infty)$ in such a way that the first integral vanishes and the second converges to $f(0)$ for $\sigma\to 0$. For the first integral we have
$\left\lvert\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\mathcal{V}\setminus B(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z\right\rvert\leq\frac{1}{Z_{\sigma}^{\Sigma}}e^{-\frac{1}{\sigma^{2}}\gamma_{h}(\alpha(\sigma))}\lVert f\rVert_{L^{1}(\mathcal{V})}\,.$
For the second integral let us write $f(z)=f(0)+\zeta_{f}(z)$ and split
$\frac{1}{Z_{\sigma}^{\Sigma}}\int_{B(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z$ $=\frac{f(0)}{Z_{\sigma}^{\Sigma}}\int_{B(\sigma)}e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z+\frac{1}{Z_{\sigma}^{\Sigma}}\int_{B(\sigma)}\zeta_{f}(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z$
$=:I(\sigma)+J(\sigma)\,.$
First let us estimate the difference between $I(\sigma)$ and $f(0)$. We have, using the substitution $z=\sigma w$ and abbreviation $B_{\mathcal{V}}(\sigma)=(\mathcal{V}\cap B(\sigma))/\sigma$,
$\lvert I(\sigma)-f(0)\rvert$
$\qquad\leq f(0)\left\lvert\frac{1}{Z_{1}^{\Sigma}}\int_{B_{\mathcal{V}}(\sigma)}e^{-\frac{1}{\sigma^{2}}h(\sigma w)}\,\text{d}w-1\right\rvert$
[…]
Estimating $I_{1}(\sigma)$ we obtain
$\lvert I_{1}(\sigma)\rvert\leq Z_{1}^{\Sigma}\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert e^{-\chi_{h}^{\sigma}(w)}-1\right\rvert\,$.
To estimate $I_{2}(\sigma)$, let us note that $\Sigma^{1/2}(\mathbb{R}^{k}\setminus B_{R}(0))\subseteq\mathbb{R}^{k}\setminus B_{\lambda_{\min}(\Sigma)R}(0)$. Then, via the substitution $u=\Sigma^{1/2}w$, we obtain
$\lvert I_{2}(\sigma)\rvert\leq\left\lvert\int_{\mathbb{R}^{k}\setminus B_{\mathcal{V}}(\sigma)}e^{-\frac{1}{2}w^{\top}\Sigma w}\,\text{d}w\right\rvert\leq\left\lvert\int_{\mathbb{R}^{k}\setminus B_{\min\{r,\alpha(\sigma)\}/\sigma}(0)}e^{-\frac{1}{2}w^{\top}\Sigma w}\,\text{d}w\right\rvert$
$\leq Z_{1}^{\Sigma}\frac{1}{(2\pi)^{k/2}}\left\lvert\int_{\Sigma^{1/2}(\mathbb{R}^{k}\setminus B_{\min\{r,\alpha(\sigma)\}/\sigma}(0))}e^{-\frac{1}{2}\lVert u\rVert^{2}}\,\text{d}u\right\rvert\leq Z_{1}^{\Sigma}\operatorname{G}_{0}\left(\lambda_{\min}(\Sigma)\frac{\min\{r,\alpha(\sigma)\}}{\sigma}\right)\,$,
with the Gaussian tail
$\operatorname{G}_{0}(R)=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}\setminus B_{R}(0)}e^{-\frac{1}{2}\lVert u\rVert^{2}}\,\text{d}u\,.$
Next we estimate $J(\sigma)$ via
$\left\lvert J(\sigma)\right\rvert$ $\leq\frac{1}{Z_{1}^{\Sigma}}\int_{B_{\mathcal{V}}(\sigma)}\zeta_{f}(\sigma w)e^{-\frac{1}{2}w^{\top}\Sigma w}e^{-\chi_{h}^{\sigma}(w)}\,\text{d}w$
$\leq\left(\sup_{w\in B_{\mathcal{V}}(\sigma)}\lvert\zeta_{f}(\sigma w)\rvert\right)\left(\sup_{w\in B_{\mathcal{V}}(\sigma)}\lvert e^{-\chi_{h}^{\sigma}(w)}-1\rvert+1\right)\,.$
In all, we obtain the following error estimate
$\left\lvert e^{\frac{1}{\sigma^{2}}h(0)}\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\mathcal{V}}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z-f(0)\right\rvert$
$\leq\frac{1}{Z_{\sigma}^{\Sigma}}e^{-\frac{1}{\sigma^{2}}\gamma(\alpha(\sigma))}\lVert f\rVert_{L^{1}(\mathcal{V})}+\frac{f(0)}{Z_{1}^{\Sigma}}(\lvert I_{1}(\sigma)\rvert+\lvert I_{2}(\sigma)\rvert)+\left\lvert J(\sigma)\right\rvert$
$\leq\frac{1}{Z_{\sigma}^{\Sigma}}e^{-\frac{1}{\sigma^{2}}\gamma(\alpha(\sigma))}\lVert f\rVert_{L^{1}(\mathcal{V})}$
$\qquad+f(0)\left(\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert e^{-\chi_{h}^{\sigma}(w)}-1\right\rvert+\operatorname{G}_{0}\left(\lambda_{\min}(\Sigma)\frac{\min\{r,\alpha(\sigma)\}}{\sigma}\right)\right)$
$\qquad+\left(\sup_{w\in B_{\mathcal{V}}(\sigma)}\lvert\zeta_{f}(\sigma w)\rvert\right)\left(\sup_{w\in B_{\mathcal{V}}(\sigma)}\lvert e^{-\chi_{h}^{\sigma}(w)}-1\rvert+1\right)\,.$
Thus we observe that we need to select $\alpha$ in such a way that for $\sigma\to 0$
  1. (i)
    $(Z_{\sigma}^{\Sigma})^{-1}\exp(-\gamma_{h}(\alpha(\sigma))/\sigma^{2})\to 0$
  2. (ii)
    $\operatorname{G}_{0}\left(\lambda_{\min}(\Sigma)\frac{\min\{r,\alpha(\sigma)\}}{\sigma}\right)\to 0$
  3. (iii)
    $\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert\chi_{h}^{\sigma}(w)\right\rvert\to 0$
  4. (iv)
    $\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert\zeta_{f}(\sigma w)\right\rvert\to 0$
Let us make the Ansatz $\alpha(\sigma)=\sigma\beta(\sigma)$ for some $\beta:(0,\infty)\to(0,\infty)$ with $\beta(\sigma)\to\infty$ for $\sigma\to 0$. Taking $\beta(\sigma)=\lvert\log(\sigma)\rvert$, noting that $\lvert e^{-x}-1\rvert\leq 2\lvert x\rvert$ for $\lvert x\rvert\leq\log 2$ as well as using (28), (25) and the fact that $\sigma\lvert\log(\sigma)\rvert,\sigma\lvert\log(\sigma)\rvert^{3}\leq\sigma^{1/2}$ when $\sigma\in(0,1)$, yields the following estimates
$\exp(-\frac{1}{\sigma^{2}}\gamma(\alpha(\sigma)))\leq\exp(-\min\{c\log(\sigma)^{2},\Delta\sigma^{-2}\})\,$,
[…]
$\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert e^{-\chi_{h}^{\sigma}(w)}-1\right\rvert\leq 2\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert\chi_{h}^{\sigma}(w)\right\rvert\leq 2C\sigma\beta(\sigma)^{3}=2C\sigma\lvert\log(\sigma)\rvert^{3}\,$,
$\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert\zeta_{f}(\sigma w)\right\rvert\leq C\sigma\beta(\sigma)=C\sigma\lvert\log(\sigma)\rvert\,$,
when $0<\sigma\leq\overline{\sigma}:=\min\{\eta^{2},(\log 2/(2C))^{2}\}$. Plugging all these estimates back yields (30) with
$\begin{aligned} E(\sigma)&=\frac{1}{Z_{\sigma}^{\Sigma}}\exp(-\min\{c\log(\sigma)^{2},\Delta\sigma^{-2}\})\lVert f\rVert_{L^{1}(\mathcal{V})}\\ &\qquad+f(0)\left(2C\sigma\lvert\log(\sigma)\rvert^{3}+\operatorname{G}_{0}(\lambda_{\min}(\Sigma)\min\{r\sigma^{-1},\lvert\log(\sigma)\rvert\})\right)\\ &\qquad+C\sigma\lvert\log(\sigma)\rvert(1+2C\sigma\lvert\log(\sigma)\rvert^{3})\\ &\overset{\sigma\to 0}{=}O(\sigma\log(\sigma)^{3})=\widetilde{O}(\sigma)\,.\end{aligned}$ (31)
The next result provides an nonasymptotic estimate of the Laplace method for the second order expansion. We will need a version that allows for a non-zero gradient of $f$, as long as it is contained in the subspace $\{(q,-q)\mid q\in\mathbb{R}^{n}\}\subseteq\mathbb{R}^{n}\times\mathbb{R}^{n}$.
Theorem 14 (Second Order).
Let $f,h\in C^{3}(\mathcal{V}\times\mathcal{V})$ be such that $f(0)=0$ and $\nabla h(0)=0$ for some open $B_{r}(0)\subseteq\mathcal{V}\subseteq\mathbb{R}^{n}$ for some $r>0$. Further, suppose that $h$ is additively separable as $h(z)=\overline{h}(z_{1})+\overline{h}(z_{2})$ and $\nabla f(0)=\left(\begin{smallmatrix}q\\ -q\end{smallmatrix}\right)$ for some $q\in\mathbb{R}^{n}$. Additionally, let $\eta\in(0,r]$ and $C,c,\Delta>0$ be such that (24), (26) and (27) hold, where this time (24) applies also for $h=f$. Then for $\sigma\in(0,\overline{\sigma})$ with $\overline{\sigma}=\min\{\eta^{2},(\log(2)/(2C))^{2}\}$ it holds
$\left\lvert e^{\frac{1}{\sigma^{2}}h(0)}\frac{\sqrt{\det\Sigma}}{Z_{\sigma}}\int_{\mathcal{V}\times\mathcal{V}}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z-\sigma^{2}A(f,h)\right\rvert\leq E_{2}(\sigma;f,h,r)\,$, (32)
where $\Sigma=\nabla^{2}h(0)$ with $E_{2}(\sigma;f,h,r)=O(\sigma^{3}\lvert\log(\sigma)\rvert^{3})$ for $\sigma\to 0$. Here the second order coefficient $A_{2}(f,h)$ is given by
$\begin{aligned} A_{2}(f,h)&=\frac{1}{2}\operatorname{tr}(\Sigma^{-1}\nabla^{2}f(0))\,.\end{aligned}$ (33)
and $E_{2}(\sigma;f,h,r)$ is given in (34).
Proof.
The proof is similar to the proof of Theorem 13 with the modification for the product space $\mathbb{R}^{k}=\mathbb{R}^{n}\times\mathbb{R}^{n}$. Again we can assume $h(0)=0$. Let $\alpha:[0,\infty)\to[0,\infty)$ be any function and denote $\widehat{B}(\sigma)=B_{\alpha(\sigma)}(0)\times B_{\alpha(\sigma)}(0)\subseteq\mathbb{R}^{n}\times\mathbb{R}^{n}$ and $\widehat{\mathcal{V}}=\mathcal{V}\times\mathcal{V}$. Write
$\frac{1}{\sigma^{2}}\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\widehat{\mathcal{V}}}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z$
$\quad=\frac{1}{\sigma^{2}}\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\widehat{\mathcal{V}}\setminus\widehat{B}(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z+\frac{1}{\sigma^{2}}\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\widehat{\mathcal{V}}\cap\widehat{B}(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z\,.$
We will carefully craft $\alpha$ such that $\lim_{\sigma\to 0}\alpha(\sigma)=0$ and the first and second integral converge to 0 and $\frac{1}{2}\operatorname{tr}(\Sigma^{-1}\nabla^{2}f(0))$, respectively, with explicit error bounds. For the first one we have
$\left\lvert\frac{1}{\sigma^{2}}\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\widehat{\mathcal{V}}\setminus\widehat{B}(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z\right\rvert\leq\frac{e^{-\frac{1}{\sigma^{2}}\gamma_{h}(\alpha(\sigma))}}{Z_{\sigma}^{\Sigma}\sigma^{2}}\lVert f\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}\,.$
Now consider the second integral. By a substitution $z=\sigma w$ and abbreviating $\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)=(\widehat{\mathcal{V}}\cap\widehat{B}(\sigma))/\sigma=\widetilde{B}_{\mathcal{V}}(\sigma)\times\widetilde{B}_{\mathcal{V}}(\sigma)$ with $\widetilde{B}_{\mathcal{V}}(\sigma)=(\mathcal{V}\cap B_{\alpha(\sigma)}(0))/\sigma$ as well as $\chi_{g}^{\sigma}(w)=\frac{1}{\sigma^{2}}\chi_{g}(\sigma w)$ for $g\in\{f,h\}$, we have
$\frac{1}{\sigma^{2}}\frac{1}{Z_{\sigma}^{\Sigma}}\int_{\widehat{\mathcal{V}}\cap\widehat{B}(\sigma)}f(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z$
$\quad=\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\left(\frac{1}{\sigma}\nabla f(0)^{\top}w+\frac{1}{2}w^{\top}\nabla^{2}f(0)w+\chi_{f}^{\sigma}(w)\right)e^{-\frac{1}{\sigma^{2}}h(\sigma w)}\,\text{d}w$
$\quad=:H(\sigma)+I(\sigma)+J(\sigma)\,.$
Note that $H(\sigma)$ vanishes due to separability of $h$ and the particular form of $\nabla f(0)$:
$\sigma Z_{1}^{\Sigma}H(\sigma)$ $=\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\nabla f(0)^{\top}w\cdot e^{-\frac{1}{\sigma^{2}}h(\sigma w)}\,\text{d}w$
$=\int_{\widetilde{B}_{\mathcal{V}}(\sigma)}\int_{\widetilde{B}_{\mathcal{V}}(\sigma)}(q^{\top}w_{1}-q^{\top}w_{2})e^{-\frac{1}{\sigma^{2}}\overline{h}(\sigma w_{1})}e^{-\frac{1}{\sigma^{2}}\overline{h}(\sigma w_{2})}\,\text{d}w_{1}\,\text{d}w_{2}$
$=0\,.$
Next, let us consider the difference between $I(\sigma)$ and $\frac{1}{2}\operatorname{tr}(\Sigma^{-1}\nabla^{2}f(0))$ given by
$\left\lvert\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\frac{1}{2}w^{\top}\nabla^{2}f(0)w\cdot e^{-\frac{1}{2}w^{\top}\Sigma w}e^{-\chi_{h}^{\sigma}(w)}\,\text{d}w-\frac{1}{2}\operatorname{tr}(\Sigma^{-1}\nabla^{2}f(0))\right\rvert$
$\quad\leq\left\lvert\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\frac{1}{2}w^{\top}\nabla^{2}f(0)w\cdot e^{-\frac{1}{2}w^{\top}\Sigma w}(e^{-\chi_{h}^{\sigma}(w)}-1)\,\text{d}w\right\rvert$
$\qquad+\left\lvert\frac{1}{Z_{1}^{\Sigma}}\int_{\mathbb{R}^{2n}\setminus\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\frac{1}{2}w^{\top}\nabla^{2}f(0)w\cdot e^{-\frac{1}{2}w^{\top}\Sigma w}\,\text{d}w\right\rvert$
$\quad=:\lvert I_{1}(\sigma)\rvert+\lvert I_{2}(\sigma)\rvert\,$.
Then
$\lvert I_{1}(\sigma)\rvert$ $\leq\left(\sup_{w\in\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\lvert e^{-\chi_{h}^{\sigma}(w)}-1\rvert\right)\frac{1}{Z_{1}^{\Sigma}}\int_{\mathbb{R}^{2n}}\frac{1}{2}\left\lvert w^{\top}\nabla^{2}f(0)w\right\rvert\cdot e^{-\frac{1}{2}w^{\top}\Sigma w}\,\text{d}w$
[…]
Furthermore, by a variable substitution $u=\Sigma^{1/2}w$ we obtain
[…]
$\leq\frac{1}{2}L(\sigma)\lVert\Sigma^{-1/2}\nabla^{2}f(0)\Sigma^{-1/2}\rVert\,$,
where
$L(\sigma)=\frac{1}{Z_{1}}\int_{\mathbb{R}^{k}\setminus\overline{\Sigma}^{1/2}\widetilde{B}_{\mathcal{V}}(\sigma)}\lVert u\rVert^{2}e^{-\frac{1}{2}\lVert u\rVert^{2}}\,\text{d}u\,.$
Noting that for $\overline{\Sigma}=\nabla^{2}h(0)$ it holds $\Sigma^{1/2}\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)=\overline{\Sigma}^{1/2}\widetilde{B}_{\mathcal{V}}(\sigma)\times\overline{\Sigma}^{1/2}\widetilde{B}_{\mathcal{V}}(\sigma)$ and
$\overline{\Sigma}^{1/2}\widetilde{B}_{\mathcal{V}}(\sigma)=\overline{\Sigma}^{1/2}(\mathcal{V}\cap B_{\alpha(\sigma)}(0))/\sigma$ $\supseteq B_{\lambda_{\min}(\Sigma)\min\{r,\alpha(\sigma)\}/\sigma}(0)$
and hence
$L(\sigma)\leq 2\operatorname{G}_{2}\left(\lambda_{\min}(\Sigma)\frac{\min\{r,\alpha(\sigma)\}}{\sigma}\right)\,$,
with the Gaussian second moment tail
$\operatorname{G}_{2}(R):=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}\setminus B_{R}(0)}\lVert u\rVert^{2}e^{-\frac{1}{2}\lVert u\rVert^{2}}\,\text{d}u\,.$
Now consider the final expression
$\lvert J(\sigma)\rvert$ $=\left\lvert\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\chi_{f}^{\sigma}(w)e^{-h^{\sigma}(w)}\,\text{d}w\right\rvert$
$\leq\left(\sup_{w\in\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\lvert\chi_{f}^{\sigma}(w)\rvert\right)\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}e^{-h^{\sigma}(w)}\,\text{d}w$
Let us further estimate
$\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}e^{-h^{\sigma}(w)}\,\text{d}w$ $=\frac{1}{Z_{1}^{\Sigma}}\int_{\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}e^{-\frac{1}{2}w^{\top}\Sigma w}e^{-\chi_{h}^{\sigma}(w)}\,\text{d}w$
$\leq\left(\sup_{w\in\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\lvert e^{-\chi_{h}^{\sigma}(w)}-1\rvert\right)+1\,.$
Hence we see that we need to pick $\alpha$ in such a way that for $\sigma\to 0$
  • (i)
    $(Z_{\sigma}^{\Sigma})^{-1}\sigma^{-2}\exp(-\gamma_{h}(\alpha(\sigma))/\sigma^{2})\to 0$.
  • (ii)
    $\operatorname{G}_{2}\left(\lambda_{\min}(\Sigma)\frac{\min\{r,\alpha(\sigma)\}}{\sigma}\right)\to 0$
  • (iii)
    $\sup_{w\in\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\lvert\chi_{h}^{\sigma}(w)\rvert\to 0$
  • (iv)
    $\sup_{w\in\widetilde{B}_{\widehat{\mathcal{V}}}(\sigma)}\lvert\chi_{f}^{\sigma}(w)\rvert\to 0$
Again take $\alpha(\sigma)=\sigma\beta(\sigma)$ for $\beta(\sigma)=\lvert\log(\sigma)\rvert$. This yields
$\exp(-\frac{1}{\sigma^{2}}\gamma_{h}(\alpha(\sigma)))\leq\exp(-\min\{c\log(\sigma)^{2},\Delta\sigma^{-2}\})\,$,
[…]
$\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert e^{-\chi_{h}^{\sigma}(w)}-1\right\rvert\leq 2\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert\chi_{h}^{\sigma}(w)\right\rvert\leq 2C\sigma\beta(\sigma)^{3}=2C\lvert\log(\sigma)\rvert^{3}\,$,
$\sup_{w\in B_{\mathcal{V}}(\sigma)}\left\lvert\chi_{f}^{\sigma}(w)\right\rvert\leq C\sigma\beta(\sigma)^{3}=C\lvert\log(\sigma)\rvert^{3}\,$,
when $0<\sigma\leq\overline{\sigma}:=\min\{\eta^{2},(\log(2)/(2C))^{2}\}$. Combining all these estimates one obtains, as in the proof of Theorem 13, that (32) holds with
$\begin{aligned} \frac{1}{\sigma^{2}}E(\sigma)&=\frac{\sqrt{\det\Sigma}}{Z_{\sigma}\sigma^{2}}\exp(-\min\{c\log(\sigma)^{2},\Delta\sigma^{-2}\})\lVert f\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}\\ &\qquad+\lVert\Sigma^{-1/2}\nabla^{2}f(0)\Sigma^{-1/2}\rVert\left(nC\sigma\lvert\log(\sigma)\rvert^{3}+\operatorname{G}_{2}\left(\lambda_{\min}(\Sigma)\min\{r\sigma^{-1},\lvert\log(\sigma)\rvert\}\right)\right)\\ &\qquad+C\sigma\lvert\log(\sigma)\rvert^{3}(1+2C\sigma\lvert\log(\sigma)\rvert^{3})\\ &\overset{\sigma\to 0}{=}O(\sigma\lvert\log(\sigma)\rvert^{3})=\widetilde{O}(\sigma)\,.\end{aligned}$ (34)
D.2 Local version of Theorem 1
First we prove a local version of Theorem 1.
Theorem 15.
Assume that there exists an subset $\mathcal{U}\subseteq\mathcal{M}$ that is $C^{3}$-diffeomorphic to an open subset of $\mathcal{V}\subseteq\mathbb{R}^{k}$ with $B_{r}(0)\subseteq\mathcal{V}$ via $\psi:\mathcal{V}\to\mathcal{U}$ for some $0\leq k\leq d$ and that $\mu\ll\operatorname{Vol}_{\mathcal{U}}$, where $\operatorname{Vol}_{\mathcal{U}}$ is the volume measure on $\mathcal{U}$. Moreover, assume that $\mu(\cdot)=\frac{\operatorname{d}\mu}{\operatorname{d}\operatorname{Vol}_{\mathcal{U}}}\in C^{3}(\mathcal{U})$. Let $x\in\mathbb{R}^{d}$ be any point with $\psi(0)=:p=\operatorname*{arg\,min}_{p\in\mathcal{M}}\lVert x-p\rVert$ and $\delta=\inf_{y\in\mathcal{M}\setminus\mathcal{U}}\lVert x-y\rVert^{2}-\lVert x-p\rVert^{2}>0$, and such that
$\Sigma=\psi^{\prime}(0)^{\top}\psi^{\prime}(0)+\sum_{i=1}^{d}(p-x)_{i}\nabla^{2}\psi_{i}(0)>0\,.$ (35)
Then there exists some $\overline{\sigma}>0$ such that for all $\sigma\in(0,\overline{\sigma})$
  1. (i)
    it holds that
    $\lVert\mathbb{E}\nu_{x,\sigma}-p\rVert$ $\leq\frac{2\big(E_{1}(\sigma;f_{0},h,r)\lVert p-x\rVert+\sqrt{d}E_{1}(\sigma;f_{1},h,r)+(\lVert p-x\rVert+\lVert\mu\rVert_{1})\Upsilon(\sigma;\Sigma)\big)}{\mu(p)}$
  2. (ii)
    it holds for $P_{0}=\psi^{\prime}(0)\Sigma^{-1}\psi^{\prime}(0)^{\top}$ that
    $\left\lVert\frac{1}{\sigma^{2}}\operatorname{Cov}(\nu_{x,\sigma})-P_{0}\right\rVert$
    $\hskip 56.9055pt\leq\frac{4d}{\mu(p)^{2}}\frac{E_{2}(\sigma;\overline{f},\overline{h},r)}{\sigma^{2}}$
    $\hskip 56.9055pt\qquad+\left(\frac{4(\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2})}{\mu(p)^{2}}\frac{\Upsilon(\sigma;\Sigma)^{2}}{\sigma^{2}}+12\frac{E_{1}(\sigma;f_{0},h,r)+\Upsilon(\sigma;\Sigma)}{\mu(p)}\right)\lVert P_{0}\rVert$
with $h$, $f_{0}$ and $f_{1}$ given in (36), $\overline{f}_{2}$ and $\overline{h}$ in (37), $\lVert\mu\rVert_{i}=\mathbb{E}_{y\sim\mu}\lVert y-x\rVert^{i}$, $\Upsilon(\sigma;\Sigma)=\frac{\sqrt{\det\Sigma}}{Z_{\sigma}}\exp(-\frac{1}{2\sigma^{2}}\delta)$,
$E_{1}(\sigma;f_{1},h,r)=\max_{l=1,\ldots,d}E_{1}(\sigma;f_{1}^{l},h,r)\,$,
$E_{2}(\sigma;\overline{f}_{2},\overline{h},r)=\max_{l,j=1,\ldots,d}E_{2}(\sigma;\overline{f}_{2}^{jl},\overline{h},r)\,$,
with $E_{1}(\sigma;f_{1}^{l},h,r)$ given in (31) and $E_{2}(\sigma;\overline{f}_{2}^{jl},\overline{h},r)$ in (34), respectively.
Remark 16.
This theorem only requires $\mathcal{M}$ to be locally a manifold, namely at $\mathcal{U}$.
Proof.
We denote by $\lambda=\psi^{-1}\#\mu$ be corresponding pullback measure on $\mathcal{V}$ and denote its positive density again by $\lambda\in C^{3}(\mathcal{V})$. Note that $\lambda(0)=\mu(p)$ by (20). Moreover, let v=xpv=x-p and $P_{0}=\psi^{\prime}(0)\Sigma^{-1}\psi^{\prime}(0)^{\top}$ for the rest of the proof. First we need to investigate the non-asymptotic convergence of following two integrals:
$p_{\sigma}(x)$ $=\int_{\mathcal{M}}\varphi_{\sigma}(x-y)\,\text{d}\mu(y)$
$=\underbrace{\int_{\mathcal{V}}\lambda(z)\frac{1}{Z_{\sigma}}e^{-\frac{1}{2\sigma^{2}}\lVert x-\psi(z)\rVert^{2}}\,\text{d}z}_{S_{0}}+\underbrace{\int_{\mathcal{M}\setminus\mathcal{U}}\varphi_{\sigma}(x-y)\,\text{d}\mu(y)}_{R_{0}}$
and
$p_{\sigma}(x)(\mathbb{E}\nu_{x,\sigma}-x)$ $=\int_{\mathcal{M}}(y-x)\varphi_{\sigma}(x-y)\,\text{d}\mu(y)$
$=\underbrace{\int_{\mathcal{V}}\lambda(z)(\psi(z)-x)\frac{1}{Z_{\sigma}}e^{-\frac{1}{2\sigma^{2}}\lVert x-\psi(z)\rVert^{2}}\,\text{d}z}_{S_{1}}+\underbrace{\int_{\mathcal{M}\setminus\mathcal{U}}(y-x)\varphi_{\sigma}(x-y)\,\text{d}\mu(y)}_{R_{1}}$
We can write each $S_{i}$ and $R_{i}$ as
$S_{i}=\frac{1}{Z_{\sigma}}\int_{\mathcal{V}}f_{i}(z)e^{-\frac{1}{\sigma^{2}}h(z)}\,\text{d}z\,,\quad R_{i}=\int_{\mathcal{M}\setminus\mathcal{U}}g_{i}(y)\varphi_{\sigma}(x-y)\,\text{d}\mu(y)\,$,
where
$h(z)=\frac{1}{2}\lVert x-\psi(z)\rVert^{2}\,,\quad f_{0}(z)=\lambda(z)\,,\quad f_{1}(z)=\lambda(z)(\psi(z)-x)\,$, (36)
as well as $g_{0}(y)=1$ and $g_{1}(y)=y-x$. Then $f_{0},f_{1}\in C^{3}(\mathcal{V})$ and
$h(0)=\frac{1}{2}\lVert v\rVert^{2}\,,\quad\nabla h(0)=0\,,\quad\nabla^{2}h(0)=\Sigma\,$,
where the second equality is due to $p=\psi(0)$ being the closest point from $\mathcal{U}$ to $x$. Also obviously $f_{0}(0)=\lambda(0)$ and $f_{1}(0)=\lambda(0)(p-x)$. Applying Theorem 13 component-wise, we obtain for $\sigma\in(0,\overline{\sigma}_{i})$, with $\overline{\sigma}_{i}=\overline{\sigma}_{i}(\eta,C)$ given as in Theorem 13 with $\eta,C>0$ dependent on $f_{i}$ and $h$, the estimate
$\underbrace{\left\lVert e^{\frac{1}{2\sigma^{2}}\lVert v\rVert^{2}}\sqrt{\det\Sigma}S_{i}-f_{i}(0)\right\rVert}_{=:\lVert F_{i}\rVert}\leq D_{i}(\sigma)\,$,
with
$D_{0}(\sigma)=E_{1}(\sigma;f_{0},h,r)\,,\quad D_{1}(\sigma)=\sqrt{d}E_{1}(\sigma;f_{1},h,r):=\sqrt{d}\max_{l=1,\ldots,d}E_{1}(\sigma;f_{1}^{l},h,r)\,$,
and where $E_{1}$ is given in Theorem 13 and $f_{1}^{l}$ is the $l$-th component of $f_{1}$. Furthermore, we can estimate for $i=0,1$
$\lVert R_{i}\rVert$ $=\left\lVert\int_{\mathcal{M}\setminus\mathcal{U}}g_{i}(y)\varphi_{\sigma}(x-y)\,\text{d}\mu(y)\right\rVert$
$=\frac{1}{Z_{\sigma}}\left\lVert\int_{\mathcal{M}\setminus\mathcal{U}}g_{i}(y)e^{\frac{1}{2\sigma^{2}}\lVert x-y\rVert^{2}}\,\text{d}\mu(y)\right\rVert$
$\leq e^{-\frac{1}{2\sigma^{2}}\lVert v\rVert^{2}}\frac{e^{-\frac{1}{2\sigma^{2}}\delta}}{Z_{\sigma}}\int_{\mathcal{M}\setminus\mathcal{U}}\lVert g_{i}(y)\rVert\,\text{d}\mu(y)\,$,
where we have used that $\lVert x-y\rVert^{2}\geq\delta+\lVert v\rVert^{2}$ for all $y\in\mathcal{M}\setminus\mathcal{U}$. Thus we get the estimate
$\left\lVert e^{\frac{1}{2\sigma^{2}}\lVert v\rVert^{2}}R_{i}\right\rVert\leq\frac{e^{-\frac{1}{2\sigma^{2}}\delta}}{Z_{\sigma}}\lVert\mu\rVert_{i}\,$,
where the quantity $\lVert\mu\rVert_{i}=\int_{\mathcal{M}}\lVert y-x\rVert^{i}\,\text{d}\mu(y)$ denotes the $i$-th centered moment of $\mu$. Let us denote $J=e^{\frac{1}{2\sigma^{2}}\lVert v\rVert^{2}}\sqrt{\det\Sigma}$ and express
$S_{i}=f_{i}(0)/J+F_{i}/J$
Now we can estimate the distance to $p$ by
$\mathbb{E}\nu_{x,\sigma}-p=\frac{S_{1}+R_{1}}{S_{0}+R_{0}}-(p-x)=\frac{f_{1}(0)/J+F_{1}/J+R_{1}}{f_{0}(0)/J+F_{0}/J+R_{0}}-(p-x)$
$\hskip 56.9055pt=\frac{f_{1}(0)+F_{1}+JR_{1}}{f_{0}(0)+F_{0}+JR_{0}}-(p-x)=\frac{(p-x)+T_{1}}{1+T_{0}}-(p-x)=\frac{T_{1}-T_{0}(p-x)}{1+T_{0}}\,.$
where $T_{i}=(F_{i}+JR_{i})/\lambda(0)$. Let us bound
$\lVert T_{i}\rVert\leq\frac{1}{\lambda(0)}\left(D_{i}(\sigma)+\sqrt{\det\Sigma}\frac{e^{-\frac{1}{2\sigma^{2}}\delta}}{Z_{\sigma}}\right)\lVert\mu\rVert_{i}\overset{\sigma\to 0}{\longrightarrow}0\,.$
Then $\lvert T_{0}\rvert<1/2$ if $\sigma<\overline{\sigma}(f_{0},h,r):=\min\{\overline{\sigma}_{0},\widetilde{\sigma}\}$ with $\widetilde{\sigma}$ depending on $D_{0}$, $\Sigma$ and $\delta$. For this $\sigma\in(0,\overline{\sigma})$ we obtain
$\lVert\mathbb{E}\nu_{x,\sigma}-p\rVert$
$\quad\leq 2\lvert T_{0}\rvert\lVert p-x\rVert+2\lVert T_{1}\rVert$
$\quad\leq\frac{2}{\lambda(0)}\left(E_{1}(\sigma;f_{0},h,r)\lVert p-x\rVert+\sqrt{d}E_{1}(\sigma;f_{1},h,r)+(\lVert p-x\rVert+\lVert\mu\rVert_{1})\sqrt{\det\Sigma}\frac{e^{-\frac{1}{2\sigma^{2}}\delta}}{Z_{\sigma}}\right)\,.$
This proves (i). The proof of (ii) is analogous. For $(z,\widetilde{z})\in\mathcal{V}\times\mathcal{V}$ define
$\begin{aligned} \overline{f}_{2}(z,\widetilde{z})&=\lambda(z)\lambda(\widetilde{z})((\psi(z)-x)(\psi(z)-x)^{\top}-(\psi(z)-x)(\psi(\widetilde{z})-x)^{\top})\,,\\ \overline{h}(z,\widetilde{z})&=\frac{1}{2}\lVert x-\psi(z)\rVert^{2}+\frac{1}{2}\lVert x-\psi(\widetilde{z})\rVert^{2}\,.\end{aligned}$ (37)
Then $f,h\in C^{3}(\mathcal{V}\times\mathcal{V})$, $\overline{f}_{2}(0,0)=0$ and
$\overline{h}(0,0)=\lVert v\rVert^{2}\,,\quad\nabla\overline{h}(0,0)=0\,,\quad\nabla^{2}\overline{h}(0,0)=\overline{\Sigma}:=\operatorname{diag}(\Sigma,\Sigma)\,.$
Let $\overline{f}_{2}^{jl}$ be the $(j,l)$-th entry of $\overline{f}_{2}$. Then it is an elementary, but very tedious exercise to show that $\overline{f}$ and $\overline{h}$ satisfy the conditions of Theorem 14 and that moreover the second order coefficient is given by
$A_{2}(\overline{f}_{2}^{jl},\overline{h})=\lambda(0)^{2}(\psi^{\prime}(0)\Sigma^{-1}\psi^{\prime}(0)^{\top})_{lj}=\lambda(0)^{2}(P_{0})_{lj}\,$,
i.e. $\overline{A}_{2}:=(A_{2}(\overline{f}_{2}^{jl},\overline{h}))_{l,j=1}^{d}=\lambda(0)^{2}P_{0}$. Now split
$p_{\sigma}(x)^{2}\operatorname{Cov}(\nu_{x,\sigma})$
$\qquad=p_{\sigma}(x)^{2}(\operatorname{Cov}(\nu_{x,\sigma})-\mathbb{E}\nu_{x,\sigma}(\mathbb{E}\nu_{x,\sigma})^{\top})$
[…]
$\qquad=\underbrace{\int_{\mathcal{V}\times\mathcal{V}}\overline{f}_{2}(z,\widetilde{z})\frac{1}{Z_{\sigma}^{2}}e^{-\frac{1}{2\sigma^{2}}(\lVert x-\psi(z)\rVert^{2}+\lVert x-\psi(\widetilde{z})\rVert^{2})}\,\text{d}(z,\widetilde{z})}_{\overline{S}_{2}}$
$\quad+\underbrace{\int_{\mathcal{M}\times\mathcal{M}\setminus\mathcal{U}\times\mathcal{U}}((y-x)(y-x)^{\top}-(y-x)(\widetilde{y}-x)^{\top})\varphi_{\sigma}(x-y)\varphi_{\sigma}(x-\widetilde{y})\,\text{d}(\mu\otimes\mu)(y,\widetilde{y})}_{\overline{R}_{2}}\,.$
Applying Theorem 14 component-wise to $\overline{f}_{2}$ we obtain for $\sigma\in(0,\overline{\sigma}_{2})$, with $\overline{\sigma}_{2}=\overline{\sigma}_{2}(\eta,C)$ and $\eta,C>0$ dependent on $\overline{f}_{2}$ and $\overline{h}$, that
$\underbrace{\left\lVert(e^{\frac{1}{2\sigma^{2}}\lVert v\rVert^{2}}\sqrt{\det\Sigma})^{2}\overline{S}_{2}-\sigma^{2}\overline{A}_{2}\right\rVert}_{\lVert\overline{G}_{2}\rVert}\leq dE_{2}(\sigma;\overline{f}_{2},\overline{h},r):=d\max_{l,j=1,\ldots,d}E_{2}(\sigma;\overline{f}_{2}^{jl},\overline{h},r)\,$,
with $E_{2}(\sigma;\overline{f}_{2}^{jl},\overline{h},r)$ given in Theorem 34. The term $\overline{R}_{2}$ on the other hand can be estimated again by
[…]
$\leq e^{-\frac{1}{\sigma^{2}}\lVert v\rVert^{2}}\frac{e^{-\frac{1}{\sigma^{2}}\delta}}{Z_{\sigma}^{2}}(\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2})\,.$
For $\overline{S}_{2}$ it holds
$\overline{S}_{2}=\sigma^{2}\lambda(0)^{2}P_{0}(x)/J^{2}+\overline{G}_{2}/J^{2}$
and thus
$\frac{1}{\sigma^{2}}\operatorname{Cov}(\nu_{x,\sigma})-P_{0}$ $=\frac{\overline{S}_{2}/\sigma^{2}+\overline{R}_{2}/\sigma^{2}}{(S_{0}+R_{0})^{2}}-P_{0}$
$=\frac{\lambda(0)^{2}P_{0}/J^{2}+\overline{G}_{2}/(\sigma^{2}J^{2})+\overline{R}_{2}/\sigma^{2}}{(\lambda(0)/J+F_{0}/J+R_{0})^{2}}-P_{0}$
$=\frac{P_{0}(x)+\overline{T}_{2}}{(1+T_{0})^{2}}-P_{0}$
$=\frac{\overline{T}_{2}-T_{0}(2+T_{0})P_{0}}{(1+T_{0})^{2}}$
with $\overline{T}_{2}=(\overline{G}_{2}+J^{2}\overline{R}_{2})/(\lambda(0)^{2}\sigma^{2})$. Again we have the estimate
$\lVert\overline{T}_{2}\rVert\leq\frac{d}{\lambda(0)^{2}}\frac{E_{2}(\sigma)}{\sigma^{2}}+\frac{(\det\Sigma)(\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2})}{\lambda(0)^{2}}\frac{e^{-\frac{1}{\sigma^{2}}\delta}}{Z_{\sigma}^{2}\sigma^{2}}\,$,
and for $\sigma\in(0,\overline{\sigma})$ (with $\overline{\sigma}$ as before, guaranteeing $\lvert T_{0}\rvert<1/2$)
$\left\lVert\frac{1}{\sigma^{2}}\operatorname{Cov}(\nu_{x,\sigma})-P_{0}\right\rVert$
$\leq 4\lVert\overline{T}_{2}\rVert+12\lvert T_{0}\rvert\lVert P_{0}\rVert$
$\leq\frac{4d}{\lambda(0)^{2}}\frac{E_{2}(\sigma)}{\sigma^{2}}+\left(\frac{4(\det\Sigma)(\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2})}{\lambda(0)^{2}}\frac{e^{-\frac{1}{\sigma^{2}}\delta}}{Z_{\sigma}^{2}\sigma^{2}}+12\frac{D_{0}(\sigma)+\sqrt{\det\Sigma}\frac{e^{-\frac{1}{2\sigma^{2}}\delta}}{Z_{\sigma}}}{\lambda(0)}\right)\lVert P_{0}\rVert$
D.3 Bounds in terms of distribution and manifold parameters
[…]
$h:\mathcal{V}$ $\to\mathbb{R}:z\mapsto\frac{1}{2}\lVert x-\psi_{p}(z)\rVert^{2}\,$, (38)
as well as
$f_{0}:\mathcal{V}$ $\to\mathbb{R}:z\mapsto\lambda(z)\,$, (39)
$f_{1}:\mathcal{V}$ $\to\mathbb{R}^{d}:z\mapsto\lambda(z)(\psi_{p}(z)-x)\,$, (40)
$\overline{f}_{2}:\mathcal{V}\times\mathcal{V}$ $\to\mathbb{R}^{d\times d}:(z,\widetilde{z})\mapsto\lambda(z)\lambda(\widetilde{z})((\psi_{p}(z)-x)((\psi_{p}(z)-x)^{\top}-(\psi_{p}(\widetilde{z})-x)^{\top}))\,$, (41)
where $\lambda=\psi_{p}^{-1}\#\mu=\operatorname{pr}_{p}\#\mu$.
D.3.1 Conditions (24) and (26) for $h$
First we show how to express conditions (24) and (26) in terms of the manifold parameters for $h$ given in (38). To see (24), note that for $z\in\mathcal{V}$ we have by the chain and product rule as well as Lemma 7 (i) that
$\lVert\nabla^{3}h(z)\rVert$ $\leq 3\lVert\psi_{p}^{\prime\prime}(z)\rVert\lVert\psi_{p}^{\prime}(z)\rVert+\lVert x-\psi_{p}(z)\rVert\lVert\psi_{p}^{\prime\prime\prime}(z)\rVert$
$\leq 3M^{2}+(\lVert x-p\rVert+\lVert p-\psi_{p}(z)\rVert)M$
$\leq 3M^{2}+(\tau_{\mathcal{M}}+\frac{8}{7}\lVert z\rVert)M$
$\leq(3M+\frac{9}{7}\tau_{\mathcal{M}})M\,.$
Thus (24) holds with, say $C_{\tau_{\mathcal{M}},M}^{(1)}=\frac{1}{2}(M+\tau_{\mathcal{M}})M$, and $\eta=\min\{\tau_{\mathcal{M}},M\}/4$. Next, for (26) and $\eta=\min\{\tau_{\mathcal{M}},M\}/4$ we compute for $z\in B_{\eta}^{\operatorname{T}_{p}\mathcal{M}}(0)$
$h(z)-h(0)\geq\inf_{\widetilde{z}\in\mathcal{V}\setminus B_{\lVert z\rVert}^{\operatorname{T}_{p}\mathcal{M}}(0)}h(\widetilde{z})-h(0)$ $=\inf_{q\in\psi_{p}(\mathcal{V}\setminus B_{\lVert z\rVert}^{\operatorname{T}_{p}\mathcal{M}}(0))}\frac{1}{2}\lVert x-q\rVert^{2}-\frac{1}{2}\lVert x-p\rVert^{2}$
$\geq\inf_{q\in\mathcal{M}\setminus B_{\lVert z\rVert}(p)}\frac{1}{2}\lVert x-q\rVert^{2}-\frac{1}{2}\lVert x-p\rVert^{2}$
$\geq\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}\lVert z\rVert^{2}\,.$
In the first inequality we have used $\psi_{p}(\mathcal{V}\setminus B_{\lVert z\rVert}^{\operatorname{T}_{p}\mathcal{M}}(0))\subseteq\mathcal{M}\setminus B_{\lVert z\rVert}(p)$, which follows from Lemma 7 (i), whereas in the last inequality we have used Lemma 11 and $x\in\mathcal{T}(\tau)$. Thus (26) holds on $\mathcal{V}$ with $c=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and implies (29).
D.3.2 Lower bound on $\lambda_{\min}(\Sigma)$ and upper bound on $\sqrt{\det\Sigma}$
We first consider $\lambda_{\min}(\Sigma)$ for $\Sigma=\nabla^{2}h(0)$ with $h$ given in (38) and when $x\in\mathcal{T}(\tau)$. By Lemma 8 and (17) we directly obtain
$\lambda_{\min}(\Sigma)=\lVert\Sigma^{-1}\rVert^{-1}\geq 1-\lVert x-p\rVert\kappa_{\mathcal{M}}\geq 1-\frac{\tau}{\tau_{\mathcal{M}}}\,.$
Due to (17) and the definition of the shape operator $S_{p}$, an upper bound on $\sqrt{\det\Sigma}$ is given by
$\sqrt{\det\Sigma}=\sqrt{\det(I_{\operatorname{T}_{p}\mathcal{M}}+S_{p}^{p-x})}\leq\left(1+\frac{\lVert p-x\rVert}{\rho(p,p-x)}\right)^{k/2}\leq\left(1+\frac{\tau}{\tau_{\mathcal{M}}}\right)^{k/2}\leq 2^{k/2}\,.$
D.3.3 Condition (24) for $\overline{f}_{2}^{jl}$
[…]
$\overline{f}_{2}^{jl}(z,\widetilde{z})=\lambda(z)\lambda(\widetilde{z})((\psi_{p}^{j}(z)-x)(\psi_{p}^{l}(z)-x)-(\psi_{p}^{j}(z)-x)(\psi_{p}^{l}(\widetilde{z})-x))\,$,
can be obtained by direct differentiation. We don’t pursue the complete derivation here and instead say that (24) holds with, say, $C_{\tau_{\mathcal{M}},M,\mu}^{(2)}$.
D.3.4 Upper bound on $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}$, $\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}$ and $\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}$ and $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert$
We clearly have $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}=\lambda(\mathcal{V})\leq 1$. Next, due to $\lVert y-x\rVert\leq\operatorname{diam}(\mathcal{M})+\lVert p-x\rVert\leq\operatorname{diam}(\mathcal{M})+\tau$ for $y\in\mathcal{M}$, we have
$\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}\leq\lVert f_{1}\rVert_{L^{1}(\mathcal{V};\mathbb{R}^{d})}=\int_{\mathcal{U}}\lVert y-x\rVert\,\text{d}\mu(y)=\lVert\mu\rVert_{1}\leq\operatorname{diam}(\mathcal{M})+\tau\,$,
and
$\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}$ $\leq\lVert\overline{f}_{2}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V};\mathbb{R}^{d\times d})}$
[…]
$\leq\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2}$
$\leq 2(\operatorname{diam}(\mathcal{M})+\tau)^{2}\,.$
Furthermore for $v=\left(\begin{smallmatrix}v_{1}\\ v_{2}\end{smallmatrix}\right)\in\mathbb{R}^{k}\times\mathbb{R}^{k}$ we have
$\overline{f}_{2}^{\prime\prime}(0,0)[v,v]$ $=\lambda(0)^{2}\cdot(2v_{1}v_{1}^{\top}-2v_{1}v_{2}^{\top}+(p-x)(\mathrm{I\!I}_{p}(v_{1},v_{2})-\mathrm{I\!I}_{p}(v_{2},v_{2}))^{\top})$
$\qquad+2\lambda(0)\nabla\lambda(0)^{\top}(v_{1}+v_{2})\cdot(p-x)(v_{1}-v_{2})^{\top}\,.$
Hence, using (20) and the fact that $\lVert p-x\rVert\lVert\mathrm{I\!I}_{p}(v_{1},v_{2})\rVert\leq\frac{\tau}{\tau_{\mathcal{M}}}\leq 1$ for any unit vectors $v_{1},v_{2}\in\operatorname{T}_{p}\mathcal{M}$ we obtain
$\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert\leq\lVert\overline{f}_{2}^{\prime\prime}(0,0)\rVert\leq 6\mu(p)^{2}+8\tau\mu(p)\lVert\operatorname{grad}_{\mathcal{M}}\mu(p)\rVert\,.$
D.4 Proof of Theorem 1
We apply Theorem 15 to every point $x\in\mathcal{T}(\tau)$ and $p=\pi(x)$ with the graph chart $\psi=\psi_{p}$. It remains to provide universal constants for the bounds in Theorem 15 (i) and (ii). First we bound $E_{1}(\sigma;f_{0},h,r,\theta)$ with $E_{1}$ given in (31). By Section D.3 we can take $r_{\tau_{\mathcal{M}},M}=\eta=\min\{\tau_{\mathcal{M}},M\}/4$, $c_{\tau_{\mathcal{M}},M}=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and $C_{\tau_{\mathcal{M}},M,\mu}=\max\{C_{\tau_{\mathcal{M}},M}^{(1)},C_{\tau_{\mathcal{M}},M,\mu}^{(2)}\}$ and hence, due to $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}\leq 1$,
$E_{1}(\sigma;f_{0})$ $\leq\frac{2^{k/2}}{Z_{\sigma}}\exp\left(-c_{\tau_{\mathcal{M}},M}\log(\sigma)^{2}\right)$
[…]
$\qquad+C_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert(1+2C_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert^{3})\,.$
Similarly, due to $\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}\leq\lVert\mu\rVert_{1}$, we have the bound
$E_{1}(\sigma;f_{1})$ $\leq\frac{2^{k/2}}{Z_{\sigma}}\exp\left(-c_{\tau_{\mathcal{M}},M}\log(\sigma)^{2}\right)(\operatorname{diam}(\mathcal{M})+\tau)$
[…]
$\qquad+C_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert(1+2C_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert^{3})\,.$
Further, due to $\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}\leq\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2}$ and the bound on $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert$ we have
$E_{2}(\sigma;\overline{f}_{2})$ $\leq 2\frac{2^{k/2}}{Z_{\sigma}}\exp(-c_{\tau_{\mathcal{M}},M}\log(\sigma)^{2})(\operatorname{diam}(\mathcal{M})+\tau)^{2}$
$\qquad+\mu(p)\frac{6\mu(p)+8\tau\lVert\operatorname{grad}_{\mathcal{M}}\mu(p)\rVert}{1-\tau/\tau_{\mathcal{M}}}\Bigg(kC_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert^{3}$
$)\displaystyle\hskip 113.81102pt+\operatorname{G}_{2}\left((1-\frac{\tau}{\tau_{\mathcal{M}}})\min\{r_{\tau_{\mathcal{M}},M}\sigma^{-1},\lvert\log(\sigma)\rvert\right$$\Bigg)$
$\qquad+C_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert^{3}(1+2C_{\tau_{\mathcal{M}},M,\mu}\sigma\lvert\log(\sigma)\rvert^{3})$
Now by Lemma 11 we can pick $\delta=\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}\eta^{2}=2c_{\tau_{\mathcal{M}},M}r_{\tau_{\mathcal{M}},M}^{2}$ and thus
$\Upsilon(\sigma;\Sigma)\leq\frac{2^{k/2}}{Z_{\sigma}}\exp\left(-\frac{c_{\tau_{\mathcal{M}},M}r_{\tau_{\mathcal{M}},M}^{2}}{\sigma^{2}}\right)\,.$
Finally, $\lVert P_{0}\rVert\leq(1-\tau/\tau_{\mathcal{M}})^{-1}$. Note that since $\mu(\cdot)\in C^{3}(\mathcal{M})$ and $\operatorname{supp}\mu=\mathcal{M}$, there exist positive lower and upper bounds on $\mu(\cdot)$ and its derivatives on $\mathcal{M}$. This shows that there $K$ and $\overline{\sigma}$ depending only on $\tau$, $M$ and $\operatorname{diam}(\mathcal{M})$ (as $k$) and $\mu$ such that (5) hold and finishes the proof.
D.5 A useful corollary to Theorem 1
The following elementary consequence of Theorem 1 will be useful in the proofs of the results from Section 5.
Corollary 17.
Suppose that for some $s:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tau\in(0,\tau_{\mathcal{M}})$ and $\epsilon\leq\tau$ we have
$\lVert s(x)-\pi(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (42)
Then $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$. Moreover, for $x\in\mathcal{T}(\epsilon)$ we have $\lVert s(x)-x\rVert\leq 2\epsilon$ and for $x\in\mathcal{T}(\tau)$ we have $\lVert s(x)-x\rVert\leq 2\tau$.
Proof.
From (42) and the fact that $\pi(x)\in\mathcal{M}$ we clearly have $s(x)\in\mathcal{T}(\epsilon)$ for any $x\in\mathcal{T}(\tau)$, i.e. $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$. The last claims follows then from the triangle inequality $\lVert s(x)-x\rVert\leq\lVert s(x)-\pi(x)\rVert+\lVert\pi(x)-x\rVert$. ∎
Appendix E Approximate Riemannian gradient flow with landing
In this section $C=\lVert\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})})$ denote the supremum of $\nabla f$ and the Lipschitz constant of $\nabla f$ on $\mathcal{T}(\tau_{\mathcal{M}})$ for some manifold $\mathcal{M}$. We will often assume the following approximation condition on a function $s\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$:
$\lVert s(x)-\pi(x)\rVert<\epsilon\,,\quad\lVert s^{\prime}(x)-P_{0}(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (43)
We will also abbreviate the (negative) right hand side of the dynamics (8) by
$G_{s}^{\eta}(x)=s^{\prime}(x)\nabla f(x)+\eta(x-s(x))$ (44)
E.1 Stationary points of (8) and optimality criteria
We need to analyze the meaning of approximate stationary points of (8) for the optimization problem (1).
Lemma 18.
[…]
$\lVert G_{s}^{\eta}(x_{*})\rVert\leq\delta\,.$ (45)
Then for $p_{*}=\pi(x_{*})$ it holds that
$\lVert\operatorname{grad}_{\mathcal{M}}f(p_{*})\rVert\leq 2(L+C+\eta)\epsilon+2\frac{\widetilde{\tau}/\tau_{\mathcal{M}}}{1-\widetilde{\tau}/\tau_{\mathcal{M}}}C+\delta\,.$ (46)
Proof.
We have the estimates
$\lVert\eta(s(x_{*})-x_{*})-\eta(\pi(x_{*})-x_{*})\rVert\leq\eta\epsilon\,$,
and
$\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-s^{\prime}(x_{*})\nabla f(s^{\prime}(x_{*}))\rVert$
$\hskip 56.9055pt\quad\leq\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-P_{0}(\pi(x_{*}))\nabla f(s(x_{*}))\rVert$
$\hskip 56.9055pt\qquad+\lVert P_{0}(\pi(x_{*}))\nabla f(s(x_{*}))-P_{0}(x_{*})\nabla f(s(x_{*}))\rVert$
$\hskip 56.9055pt\qquad+\lVert P_{0}(x_{*})\nabla f(s(x_{*}))-s^{\prime}(x_{*})\nabla f(s(x_{*}))\rVert$
$\hskip 56.9055pt\quad\leq(\operatorname{Lip}(\nabla f)+\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty})\epsilon+\lVert P_{0}(\pi(x_{*}))-P_{0}(x_{*})\rVert\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}\,$,
where we have used that $P_{0}(\pi(x_{*}))$ is an orthogonal projection and hence has unit norm. By Lemma 23 applied to $z_{1}=s^{\prime}(x_{*})\nabla f(s(x_{*}))$ and $z_{2}=\eta(s(x_{*})-x_{*})$, Lemma 9 and the fact that $P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))=\operatorname{grad}_{\mathcal{M}}f(\pi(x_{*}))$, the inequality (46) follows. ∎
E.2 Lie derivative of manifold distance
Lemma 19.
Suppose that $\tau\in(0,\tau_{\mathcal{M}})$ and $\sigma>0$ is such that (43) is satisfied. Then it holds
$-\left\langle\nabla\operatorname{d}(x),G_{s}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$
Proof.
We have for $x\in\mathcal{T}(\tau)$
$\left\langle x-\pi(x),-G_{s}^{\eta}(x)\right\rangle$
$\hskip 56.9055pt=-2\eta\operatorname{d}(x)+\left\langle\pi(x)-x,(s^{\prime}(x)-P_{0}(x))\nabla f(s(x))\right\rangle$
$\hskip 56.9055pt\qquad+\eta\left\langle x-\pi(x),s(x)-\pi(x)\right\rangle$
$\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+\lVert\pi(x)-x\rVert\lVert s^{\prime}(x)-P_{0}(x)\rVert\lVert\nabla f(s(x))\rVert$
$\hskip 56.9055pt\qquad+\eta\lVert\pi(x)-x\rVert\lVert s(x)-\pi(x)\rVert$
$\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+(\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}\epsilon+\eta\epsilon)\sqrt{2\operatorname{d}(x)}\,.$
E.3 Exact landing with $\sigma=0$
Here we establish the following noiseless version of Theorem 3.
Theorem 20.
Consider the flow (8) for $\sigma=0$ and $\eta\geq 0$ with $x(0)\in\mathcal{T}(\tau)$ for some $\tau\in(0,\tau_{\mathcal{M}})$. Then the solution $x(t)$ exists for all times $t\geq 0$ and is contained in $\mathcal{T}(\tau)$. Moreover every accumulation point $x_{*}$ of this flow satisfies $\operatorname{grad}_{\mathcal{M}}f(x_{*})=0$ and is at most $\tau$ away from the point $p_{*}=\pi(x_{*})\in\mathcal{M}$ with $\lVert\operatorname{grad}_{\mathcal{M}}f(p_{*})\rVert\leq L\tau$. Moreover, if $\eta>0$, then $x_{*}=p_{*}\in\mathcal{M}$ and $\operatorname{grad}_{\mathcal{M}}f(p_{*})=0$, i.e. every accumulation point is critical.
Proof.
Note first that
$\frac{\operatorname{d}}{\operatorname{d}t}\operatorname{d}(x)=\left\langle x-\pi(x),\dot{x}\right\rangle=-\eta\lVert x-\pi(x)\rVert^{2}=-2\eta\operatorname{d}(x)\,$,
i.e. $\operatorname{d}(x(t))=e^{-2\eta t}\operatorname{d}(x(0))$ and hence $x(t)\in\mathcal{T}$ for all $t\geq 0$. Moreover, if $\eta>0$ then $\operatorname{d}(x(t))\to 0$ for $t\to\infty$. Let us now consider $f\circ\pi$ as a Lyapunov function for (9) with $\sigma=0$. We observe that
$\frac{\operatorname{d}}{\operatorname{d}t}f(\pi(x))$ $=-\nabla f(\pi(x))^{\top}P_{0}(x)^{\top}P_{0}(x)\nabla f(\pi(x))$
$=-\left\langle\operatorname{P}_{\operatorname{T}_{\pi(x)}\mathcal{M}}\nabla f(x),H_{x}^{-2}\operatorname{P}_{\operatorname{T}_{\pi(x)}\mathcal{M}}\nabla f(x)\right\rangle$
$=-\lVert\operatorname{grad}f(x)\rVert_{H_{x}}^{2}\,$,
with $\lVert v\rVert_{H_{x}}^{2}=\lVert H_{x}^{-1}v\rVert_{\operatorname{T}_{\pi(x)}\mathcal{M}}^{2}$. This shows that $f(x(t))$ is non-increasing for $t\to\infty$. Since $\mathcal{T}$ is bounded, $f_{*}=\lim_{t\to\infty}f(x(t))$ is finite. Moreover, we have
$\int_{0}^{\infty}\lVert\operatorname{grad}f(x(t))\rVert_{H_{x(t)}}^{2}\,\text{d}t=f(x(0))-f_{*}<\infty\,.$ (47)
Notice that $\{x(t)\mid t\geq 0\}\subseteq\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$ being compact. Since $\mathcal{T}\to\mathbb{R}:z\mapsto\lVert\operatorname{grad}f(z)\rVert_{H_{z}}^{2}$ is continuous, it is uniformly continuous on $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$. This, together with (47) implies, by Barbalat’s lemma Farkas & Wegner (2016), that $\lim_{t\to\infty}\lVert\operatorname{grad}f(x(t))\rVert_{H_{x(t)}}^{2}=0$. Clearly any accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $H_{x_{*}}^{-1}\operatorname{grad}f(x_{*})=0$, i.e. $\operatorname{grad}f(x_{*})=0$. Then $\pi(x_{*})\in\mathcal{M}$ is an accumulation point of $\{\pi(x(t))\mid t\geq 0\}\subseteq\mathcal{M}$ and
$\lVert\operatorname{grad}f(\pi(x_{*}))\rVert=\lVert\operatorname{grad}f(\pi(x_{*}))-\operatorname{grad}f(x_{*})\rVert\leq L\lVert\pi(x_{*})-x_{*}\rVert=L\tau\,.$
E.4 Proof of Theorem 3
Note first that, by Theorem 1, if (7) hold with $\epsilon\to\epsilon^{\prime}$, then 43 are satisfied with $\epsilon=\epsilon^{\prime}+K\sigma\lvert\log(\sigma)\rvert^{3}$ for $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$. In the following we will write $\epsilon>0$ for the latter quantity. First let us analyze the manifold distance of (8). We have by Lemma 19 whenever $x(t)\in\mathcal{T}(\tau)$ that
$\frac{\operatorname{d}}{\operatorname{d}t}\operatorname{d}(x)=-\left\langle\nabla\operatorname{d}(x),G_{s}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\,$,
where the right hand side is non-positive iff
$\operatorname{dist}_{\mathcal{M}}(x)\geq\frac{\epsilon}{2}\left(\frac{C}{\eta}+1\right)=:\tau_{0}\,.$
Note that $\tau_{0}<\tau$ if $\epsilon<2\tau/(1+C/\eta)$. In particular if $\operatorname{dist}_{\mathcal{M}}(x(0))\in[\tau_{0},\tau)$, then $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$ is invariant w.r.t. the flow (8) and $x(t)$ exists for all $t\geq 0$. Moreover, by a similar argument as in the standard proof of Lyapunov’s direct method Khalil & Grizzle (2002), for each $\delta>0$ there exists some $T>0$ such that for all $t\geq T$ it holds that $x(t)\in\mathcal{T}(\tau_{0}+\delta)$. If $s$ is a gradient field, i.e. $s=\nabla g$ for some function $g\in C^{1}(\mathbb{R}^{d})$, then $V(x):=f(s(x))+\eta(\lVert x\rVert^{2}/2-g(x))$ satisfies $\nabla V(x)=G_{s}^{\eta}(x)$. Similar as in the proof of Theorem 20 we can take $G_{s}$ as a Lyapunov function to obtain
$\frac{\operatorname{d}}{\operatorname{d}t}V(x)=-\lVert G_{s}^{\eta}(x)\rVert^{2}\,$
along the dynamics (8). By a similar argument $V(x)$ is non-increasing and every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $G_{s}^{\eta}(x)(x_{*})=0$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$. Note that the condition $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\tau)$ in Lemma 18 is satisfied when $\epsilon\leq\tau$. Applying Lemma 18 with $\delta=0$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and proves Theorem 3 for the case when $s$ is a gradient field. Now consider the more general case when $s$ is not necessarily a gradient field. In this case consider $V(x)=f(s(x))+\eta\operatorname{d}(x)$ to obtain
$\frac{\operatorname{d}}{\operatorname{d}t}V(x)$ $=-\left\langle s^{\prime}(x)\nabla f(s(x))+\eta(x-\pi(x)),G_{s}^{\eta}(x)\right\rangle$
$\leq-\lVert G_{s}^{\eta}(x)\rVert^{2}+\eta\lVert s(x)-\pi(x)\rVert\lVert G_{s}^{\eta}(x)\rVert$
$\leq-\lVert G_{s}^{\eta}(x)\rVert^{2}+\eta\epsilon\lVert G_{s}^{\eta}(x)\rVert\,.$
Thus, if $\lVert G_{s}^{\eta}(x)\rVert\geq\eta\epsilon$ we have $\frac{\operatorname{d}}{\operatorname{d}t}V(x)\leq 0$. By a barrier function argument Khalil & Grizzle (2002) this implies that every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $\lVert G_{s}^{\eta}(x_{*})\rVert\leq\eta\epsilon$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$. Again applying Lemma 18 with $\delta=\eta\epsilon$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and finishes the proof.
Appendix F Discretized Riemannian gradient flow and descent
In this section we provide proof of Theorem 5 as well as analysis of the approximate Riemannian gradient descent and the discretized landing flow. As before $C=\lVert\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})})$ denote the supremum of $\nabla f$ and the Lipschitz constant of $\nabla f$ on $\mathcal{T}(\tau_{\mathcal{M}})$ for some manifold $\mathcal{M}$. Moreover, the following bounds will be useful:
Lemma 21.
If $\tau\in(0,\tau_{\mathcal{M}})$ and (43) holds for some $\epsilon>0$, then
$\sup_{x\in\mathcal{T}(\tau)}\lVert s^{\prime}(x)\rVert_{\infty}\leq\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}}\,.$
Additionally, if $\epsilon\in(0,\tau]$, then
$\lVert G_{s}^{\eta}(x)\rVert\leq C\lVert s^{\prime}(x)\rVert+\eta\lVert s(x)-x\rVert\leq C(\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}})+2\eta\tau\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$
Proof.
Via triangle inequality, Lemma 8 and the definition of $C$. ∎
F.1 Discretized approximate Riemannian gradient flow
In this section we analyze the discretized version of (9), specifically the corresponding gradient descent
$x_{k+1}=x_{k}-\gamma_{k}\nabla F_{\sigma}^{\eta}(x_{k})\,$, (48)
for some sequence of step sizes $\{\gamma_{k}\}_{k=1}^{\infty}$. The selection of $\gamma_{k}$ can be inferred from any standard analysis of gradient descent for $F_{\sigma}^{\eta}$ to guarantee that all accumulation points of the resulting sequence $\{x_{k}\}_{k=0}^{\infty}$ are stationary points of $F_{\sigma}^{\eta}$. Since we can only interpret stationary points of $F_{\sigma}^{\eta}$ in terms of our original problem (1) when they are contained in a tubular neighborhood $\mathcal{T}(\tau)$ for some $\tau\in(0,\tau_{\mathcal{M}})$ (see Lemma 18), we drive conditions on the step-size to ensure $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\tau)$.
Theorem 22.
Let $\tau\in(0,\tau_{\mathcal{M}}/2)$ and $\sigma>0$ be such that (43) holds for some $\epsilon\in(0,\frac{\eta\tau}{2(C+\eta)}]$. Then if $\gamma_{k}\in[0,\gamma_{\operatorname{tubular}}]$ with
$\gamma_{\operatorname{tubular}}(\epsilon,\tau,\eta)=\tau\cdot\min\left\{\frac{1}{2(C(\epsilon+2)+2\eta\tau)},\frac{\frac{1}{4}\eta\tau-\frac{1}{2}(C+\eta)\epsilon}{4(C(\epsilon+4)+3\eta\tau)^{2}}\right\}$
and $x_{0}\in\mathcal{T}(\tau)$, the iterates $x_{k}$ of the discretized flow (48) belong to $\mathcal{T}(\tau)$.
Proof.
First we show that $x_{k}\in\mathcal{T}(\tau)$ implies $x_{k+1}\in\mathcal{T}(\tau)$. If $x_{k}\in\mathcal{T}(\tau/3)$, then, since $\gamma_{k}\lVert\nabla F_{\sigma}^{\eta}(x_{k})\rVert\leq\tau/2$ by Lemma 21, it follows $x_{k+1}\in\mathcal{T}(\tau)$. Hence assume $x_{k}\in\mathcal{T}(\tau)\setminus\mathcal{T}(\tau/3)$, i.e. $\sqrt{2\operatorname{d}(x_{k})}\geq\tau/2$. Let us write
$\operatorname{d}(x_{k+1})=\operatorname{d}(x_{k})-\gamma_{k}\left\langle\nabla\operatorname{d}(x_{k}),\nabla F_{\sigma}^{\eta}(x_{k})\right\rangle+\frac{1}{2}\gamma_{k}^{2}\lVert(I-P_{0})\mid_{\mathcal{T}(3\tau/2)}\rVert_{\infty}\lVert\nabla F_{\sigma}^{\eta}(x_{k})\rVert^{2}\,$,
where we have used $\nabla^{2}\operatorname{d}=I-P_{0}$ and $x_{k+1}\in\mathcal{T}(3\tau/2)$, since again $\gamma_{k}\lVert\nabla F_{\sigma}^{\eta}(x_{k})\rVert\leq\tau/2$. Then by Lemma 19
$-\left\langle\nabla\operatorname{d}(x),\nabla F_{\sigma}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\,.$
Now, Lemma 21 and the fact that $(1-3\tau/(2\tau_{\mathcal{M}}))^{-1}\leq 4$ imply
$\delta=\frac{1}{2}\lVert(I-P_{0})|_{\mathcal{T}(3\tau/2)}\rVert_{\infty}\lVert\nabla F_{\sigma}^{\eta}|_{\mathcal{T}(\tau)}\rVert_{\infty}^{2}\leq 4(C(\epsilon+4)+3\eta\tau)^{2}$
Therefore we have
$\operatorname{d}(x_{k+1})-\operatorname{d}(x_{k})\leq\gamma_{k}(-2\eta\operatorname{d}(x_{k})+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x_{k})}+\gamma_{k}\delta)\,$,
where right hand side is non-positive for all $\sqrt{2\operatorname{d}(x_{k})}\geq\tau/2$ if
$-\frac{1}{4}\eta\tau^{2}+\frac{1}{2}(C+\eta)\tau\epsilon+\gamma_{k}\delta\leq 0\,.$
or, equivalently,
$\gamma_{k}\leq\frac{1}{\delta}\left(\frac{1}{4}\eta\tau^{2}-\frac{1}{2}(C+\eta)\tau\epsilon\right)\,.$
If these are satisfied, then $\operatorname{d}(x_{k+1})\leq\operatorname{d}(x_{k})\leq\tau^{2}/8$ and therefore $x_{k+1}\in\mathcal{T}(\tau)$. In all, $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\tau)$ provided that $x_{0}\in\mathcal{T}(\tau)$. ∎
F.2 Proof of Theorem 5
Let us write $\epsilon^{\prime}=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$. First we note that $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\epsilon^{\prime})$ as soon as $x_{0}\in\mathcal{T}(\tau/2)$, because $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon^{\prime})$ and $x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k})\in\mathcal{T}(\tau)$, since
$\gamma_{k}\leq\frac{\tau}{2C(\epsilon^{\prime}+(1-\tau/\tau_{\mathcal{M}})^{-1})}\leq\frac{\tau}{2\lVert s^{\prime}(x_{k})\nabla f(x_{k})\rVert}\,$,
where first inequality is due to $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ and the second inequality due to Lemma 21. For brevity let us denote $y_{k}=x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k})$ and $z_{k}=x_{k}-\gamma_{k}P_{0}(x_{k})\nabla f(\pi(x_{k}))$. Also let $L_{0}=\operatorname{Lip}(\nabla(f\circ\pi)|_{\mathcal{T}(\tau)})$. By Lemma 10 and $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ it follows that
$L_{0}\leq C\lVert P_{0}^{\prime}|_{\mathcal{T}}(\tau)\rVert+(1-\tau/\tau_{\mathcal{M}})^{-1}L\leq 8C\left(2(\frac{3}{\tau_{\mathcal{M}}}+\tau M)+\frac{1}{\tau_{\mathcal{M}}}\right)+2L\,.$
Then, since $\lVert x_{k}-\pi(x_{k})\rVert\leq\epsilon^{\prime}$ for $k\geq 1$, we have
$f(x_{k+1})-f(x_{k})$ $=f(s(y_{k}))-f(x_{k})$
$=f(\pi(z_{k}))-f(\pi(x_{k}))+f(\pi(x_{k}))-f(x_{k})$
$\qquad+f(s(y_{k}))-f(\pi(y_{k}))+f(\pi(y_{k}))-f(\pi(z_{k}))$
$\leq-\gamma_{k}\left\langle P_{0}(x_{k})\nabla f(\pi(x_{k})),P_{0}(x_{k})\nabla f(\pi(x_{k}))\right\rangle+\gamma_{k}^{2}\frac{L_{0}}{2}\lVert P_{0}(x_{k})\nabla f(\pi(x_{k}))\rVert^{2}$
$\qquad+C\lVert x_{k}-\pi(x_{k})\rVert+C\epsilon+L_{0}\gamma_{k}(C\epsilon+L\lVert x_{k}-\pi(x_{k})\rVert)$
$\leq-\gamma_{k}(1-\frac{L_{0}}{2}\gamma_{k})\lVert P_{0}(x_{k})\nabla f(\pi(x_{k}))\rVert^{2}+(2C+L_{0}\gamma_{k}(C+L))\epsilon^{\prime}\,.$
Now let $\gamma_{k}\in[\gamma_{\min},\gamma_{\max}]\subseteq(0,\frac{2}{L_{0}})$. Then summing over $k=1,\ldots,N$ yields
$\gamma_{\min}(1-\frac{1}{2}\gamma_{\max}L_{0})\frac{1}{N}\sum_{k=1}^{N}\lVert P_{0}(x_{k})\nabla f(\pi(x_{k}))\rVert^{2}\leq\frac{f(x_{1})-f(x_{N+1})}{N}+(2C+L_{0}\gamma_{k}(C+L))\epsilon^{\prime}\,.$
Now note that by Lemma 9 and the fact that $(1-\epsilon^{\prime}/\tau_{\mathcal{M}})^{-1}\leq 2$ it holds
$\lVert\operatorname{grad}_{\mathcal{M}}f(\pi(x_{k}))\rVert^{2}$ $\leq 2\lVert(P_{0}(\pi(x_{k}))-P_{0}(x_{k}))\nabla f(\pi(x_{k}))\rVert^{2}+2\lVert P_{0}(x_{k})\nabla f(\pi(x_{k}))\rVert^{2}$
$\leq 8C^{2}(\epsilon^{\prime}/\tau_{\mathcal{M}})^{2}+2\lVert P_{0}(x_{k})\nabla f(\pi(x_{k}))\rVert^{2}\,$,
which implies
$\gamma_{\min}(1-\frac{1}{2}\gamma_{\max}L_{0})\frac{1}{N}\sum_{k=0}^{N}\lVert\operatorname{grad}_{\mathcal{M}}f(\pi(x_{k}))\rVert^{2}\leq\frac{4D}{N}+8C^{2}(\epsilon^{\prime}/\tau_{\mathcal{M}})^{2}+2(2C+L_{0}\gamma_{k}(C+L))\epsilon^{\prime}\,.$
This finishes the proof.
F.3 Auxiliary results
F.3.1 A lemma on norms of orthogonal vectors
Lemma 23.
Let $x,y\in\mathbb{R}^{n}$ be orthogonal and $\epsilon>0$. If there exists some $z\in\mathbb{R}^{n}$ with $\lVert x-z\rVert\leq\epsilon$ and $\lVert y-z\rVert\leq\epsilon$, then $\lVert x\rVert\leq 2\epsilon$, $\lVert y\rVert\leq 2\epsilon$ and $\lVert z\rVert\leq\sqrt{2}\epsilon$. If on the other hand for some $\delta>0$ there exist some $z_{1},z_{2}\in\mathbb{R}^{n}$ with $\lVert x-z_{1}\rVert\leq\epsilon$, $\lVert y-z_{2}\rVert\leq\epsilon$ and $\lVert z_{1}-z_{2}\rVert\leq\delta$, then $\lVert x\rVert\leq 2\epsilon+\delta$, $\lVert y\rVert\leq 2\epsilon+\delta$ and $\lVert z\rVert\leq\sqrt{2}\epsilon+\delta/\sqrt{2}$.
Proof.
The first claim follows by geometric considerations. The second follows from the first by noting that the midpoint $z=(z_{1}+z_{2})/2$ satisfies $\lVert z-z_{1}\rVert\leq\delta/2$ and $\lVert z-z_{2}\rVert\leq\delta/2$ and hence $\lVert x-z\rVert\leq\epsilon+\delta/2$ and $\lVert y-z\rVert\leq\epsilon+\delta/2$, i.e. $\lVert x\rVert\leq 2\epsilon+\delta$, $\lVert y\rVert\leq 2\epsilon+\delta$ and $\lVert z\rVert\leq\sqrt{2}\epsilon+\delta/\sqrt{2}$. ∎
F.3.2 Gaussian tail bounds
The following elementary lemmas give simplifications for some of the constants appearing in the proof of Theorem 13 and Theorem 14.
Lemma 24.
It holds for $R>0$ that
$\operatorname{G}_{0}(R):=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}\setminus B_{R}(0)}e^{-\frac{1}{2}\lVert u\rVert^{2}}\,\text{d}u\leq 2e^{-\frac{1}{2n}R^{2}}$
and for $R\geq 2n$ that
$\operatorname{G}_{2}(R):=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}\setminus B_{R}(0)}\lVert u\rVert^{2}e^{-\frac{1}{2}\lVert u\rVert^{2}}\,\text{d}u\leq\frac{2^{2-n/2}}{\Gamma(n/2)}e^{-\frac{1}{4}R^{2}}\,.$
In particular $\operatorname{G}_{0}(\lvert\log(\sigma)\rvert)=O(\sigma^{l})$ and $\operatorname{G}_{2}(\lvert\log(\sigma)\rvert)=O(\sigma^{l})$ for any $l\geq 1$ as $\sigma\to 0$.
Proof.
See Majerski (2015). ∎
Appendix G Further numerical experiments and implementation details
G.1 Optimization over $O(n)$
G.1.1 Implementation details
In all of our experiments, we discretize the flow equation 8 using the Euler scheme with a step size of $t_{\operatorname{step}}=1\cdot 10^{-4}$ and set the landing gain $\eta=3\cdot 10^{3}$.
Data generation: In our experiments, we take $Q=\operatorname{diag}(1,\ldots,n)$ and a randomly sampled symmetric $A\in\mathbb{S}^{n\times n}$ with $\mathcal{N}(0,1)$-entries.
Score architecture: We use the following score architecture:
$s_{\sigma}(X)=\frac{1}{\sigma}\widetilde{s}_{\sigma}(X)\,$,
with $X=\left(\begin{matrix}x_{1}&\cdots&x_{n}\end{matrix}\right)\in\mathbb{R}^{n\times n}$ and $Y=\left(\begin{matrix}y_{1}&\cdots&y_{n}\end{matrix}\right)=\widetilde{s}_{\sigma}(X)$, where
$y_{i}=\operatorname{MLP}_{l,w}([r;x_{i};\sigma])\text{\ for\ }i=1,\ldots,n\,$,
and $\operatorname{MLP}$ a fully connected multi-layer perceptron with ReLU activation function, $l$ layers of width $w$. The features $r=r(X)\in\mathbb{R}^{m}$ are
$r_{j}(X)=\operatorname{tr}(Q_{j}XK_{j}X^{\top})\,,\quad j=1,\ldots,m$
where $Q_{j},K_{j}\in\mathbb{R}^{n\times n}$ are learnable weight matrices (shared for all $i=1,\ldots,n$). For $n=10$ we take $l=4$, $w=512$, $m=128$ and for $n=20$ we take $l=4$, $w=2048$, $m=512$.
Diffusion and training parameters: We train minimizing equation 23 with the Adam optimizer, early stopping (i.e. $t\sim\operatorname{Unif}[\epsilon,T]$) with $T=3$ and $\epsilon=10^{-4}$ and a cosine learning rate scheduling from $\texttt{lr}=10^{-3}$ to $\texttt{lr}=5\cdot 10^{-5}$. We use $N_{\operatorname{epochs}}=10000$ and $N_{\operatorname{epochs}}=50000$ epochs for $n=10$ and $n=20$, respectively.
G.1.2 Experiments
Refer to caption Refer to caption
Figure 3: Objective value vs. flow time $t$ for the orthogonal manifold for $n=10$ (left, different $\sigma>0$) and for $n=20$ (right, $\sigma=0.05$)
G.2 Data-driven reference tracking
G.2.1 Benchmark systems and tracking goals
Benchmark systems: We consider two classical benchmark systems: The double pendulum and the unicycle car model LaValle (2006). For the double pendulum the state is $x=(\theta_{1},\omega_{1},\theta_{2},\omega_{2})$ with $\omega_{i}=\dot{\theta}_{i}$, gravity $g$, masses $m_{1},m_{2}$, lengths $l_{1},l_{2}$, dampings $d_{1},d_{2}$, control torque $u$ applied at joint 1 (first pendulum), and $\Delta\theta\coloneqq\theta_{2}-\theta_{1}$.
$\dot{\theta}_{1}$ $=\omega_{1},$ $\dot{\theta}_{2}$ $=\omega_{2},$
$\mathbf{M}(\theta)\,\ddot{\bm{\theta}}+\mathbf{C}(\theta,\dot{\theta})+\mathbf{G}(\theta)+\mathbf{D}\,\dot{\bm{\theta}}$ $=\bm{\tau},\quad\bm{\theta}=\left(\begin{matrix}\theta_{1}\\ \theta_{2}\end{matrix}\right),\quad\bm{\tau}=\left(\begin{matrix}u\\ 0\end{matrix}\right),$
with
$\mathbf{M}(\theta)$ $=\left(\begin{matrix}(m_{1}+m_{2})l_{1}^{2}&m_{2}l_{1}l_{2}\cos\Delta\\ m_{2}l_{1}l_{2}\cos\Delta\theta&m_{2}l_{2}^{2}\end{matrix}\right),$ $\mathbf{D}$ $=\left(\begin{matrix}d_{1}&0\\ 0&d_{2}\end{matrix}\right),$
$\mathbf{C}(\theta,\dot{\theta})$ $=\left(\begin{matrix}-\,m_{2}l_{1}l_{2}\sin\Delta\theta\ \omega_{2}^{\,2}\\ \ \ m_{2}l_{1}l_{2}\sin\Delta\theta\ \omega_{1}^{\,2}\end{matrix}\right),$ $\mathbf{G}(\theta)$ $=\left(\begin{matrix}(m_{1}+m_{2})gl_{1}\sin\theta_{1}\\ m_{2}gl_{2}\sin\theta_{2}\end{matrix}\right).$
We pick the output $y=(\theta_{1},\theta_{2})$ and set $m_{1}=l_{1}=g=1$, $m_{2},l_{2}=0.5$ and $d_{1},d_{2}=0.1$. For the unicycle car model the dynamics is given by $x=(\text{x},\text{y},\theta)$ with
$\dot{\text{x}}=v\cos(\theta)\,,\quad\dot{\text{y}}=v\sin(\theta)\,,\quad\dot{\theta}=\omega$
and the input $u=(v,\omega)$ and output $y=x=(\text{x},\text{y},\theta)$. Here $(\text{x},\text{y})$, $v$, $\theta$, $\omega$ is the car’s position, velocity, angle and angular velocity, respectively.
Tracking goals: For the double pendulum system the goal is to track a reference trajectory $r$ via the first joint angle $\theta_{1}$ and we pick the optimal control objective $f$ to be 15 with
$Q=\left(\begin{matrix}10&0\\ 0&10\end{matrix}\right)\,,\quad R=0.01\,$,
while for the unicycle car model the goal is to track a positional reference $r=(r_{\text{x}},r_{\text{y}})$, i.e. we pick
$Q=\left(\begin{matrix}10&0&0\\ 0&10&0\\ 0&0&0\end{matrix}\right)\,,\quad R=\left(\begin{matrix}0.01&0\\ 0&0.01\end{matrix}\right)\,.$
Here $R$ is some small penalty on the input $\bm{u}$ to keep it bounded during the optimization.
G.2.2 Implementation details
Discretization: We discretize both continuous-time dynamics $\dot{x}=\bm{f}_{\operatorname{cont}}(x,u)$ via the RK4-method and discretization step $\Delta t$ to obtain the discrete-time dynamics (13) as
$\bm{f}(x,u)=x+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4})\text{\ \ where\ \ }\begin{cases}k_{1}&=\bm{f}_{\operatorname{cont}}(x,u)\\ k_{2}&=\bm{f}_{\operatorname{cont}}(x+\frac{1}{2}\Delta tk_{1},u)\\ k_{3}&=\bm{f}_{\operatorname{cont}}(x+\frac{1}{2}\Delta tk_{2},u)\\ k_{4}&=\bm{f}_{\operatorname{cont}}(x+\Delta tk_{3},u)\end{cases}$
We use $\Delta t=0.1$ for the double pendulum and $\Delta t=0.05$ for the unicycle model.
Data generation: To generate trajectories, we use i.i.d. random inputs $u_{k}\sim\operatorname{Unif}[-5,5]$ for the double pendulum and $u_{k}=(v_{k},\omega_{k})\sim\operatorname{Unif}[0,1]\otimes\mathcal{N}(0,25)$ and a horizon of $N_{h}=100$ for both system. We use $N_{\operatorname{data}}=50000$ trajectories for the double pendulum and $N_{\operatorname{data}}=20000$ trajectories for the unicycle model.
Score architecture: The score architecture is a $1$-dimensional version of the standard UNet architecture Ronneberger et al. (2015) with a sin-cos time-embedding Song et al. (2020) and residual connections, where the different input-, state- and output dimensions are concatenated and treated as additional channels. The down- and upsampling convolutions are done w.r.t. the temporal dimension and channels.
Diffusion and training parameters: Same as in Appendix G.1.1 with this time $N_{\operatorname{epochs}}=50000$ training epochs. For the DRGD step-size we pick a fixed step-size of $\gamma=0.001$.
G.2.3 Experiments
In Figure 4 we present the objective value evolution for the experiment from Section 6.2.
Figure 4: Denoising Riemannian gradient descent: Objective value $f$ vs the iteration count $j$ for double pendulum (left) and unicycle car model (right). Note the logarithmic scale on the y-axis. The current cost (blue, dashed) is the objective value $f(\bm{u}_{j},\bm{y}_{j})$ at the current (in general infeasible $(\bm{u}_{j},\bm{y}_{j})\notin\mathcal{M}_{\operatorname{IO}}$) iterate, while the true cost (orange) is the value $f(\bm{u}_{j},\bm{y}_{j}^{\operatorname{true}})$, with $\bm{y}_{j}^{\operatorname{true}}$ obtained by simulating (13) with input $\bm{u}_{j}$.
In Figure 5 we show optimized trajectories for two other reference trajectories. Note that we have set our iteration budget at $N=4000$, while the objective is still decreasing. How to accelerate the denoising Riemannian gradient descent without losing feasibility is a core question for future work.
Figure 5: Denoising Riemannian gradient descent: Unicycle car position (right) with the optimized output trajectory $\bm{y}^{*}$ (blue, dashed), the true system trajectory $\bm{y}^{\operatorname{true}}$ (orange), the initial trajectory $\bm{y}_{0}$ (green, dotted) and the reference trajectory $\bm{r}$ (red)

### v2

@@
-Landing with the Score: Riemannian Optimization through Denoising
+arXiv:2509.23357v2[cs.LG] 28 Feb 2026
 Abstract
-Under the data manifold hypothesis, high-dimensional data are concentrated near a low-dimensional manifold.
-We study the problem of Riemannian optimization over such manifolds when they are given only implicitly through the data distribution, and the standard manifold operations required by classical algorithms are unavailable.
+Under the data manifold hypothesis, high-dimensional data concentrate near a low-dimensional manifold.
+We study Riemannian optimization when this manifold is only given implicitly through the data distribution, and standard geometric operations are unavailable.
 This formulation captures a broad class of data-driven design problems that are central to modern generative AI.
-Our key idea is to introduce a link function that connects the data distribution to the geometric operations needed for optimization.
-We show that this function enables the recovery of essential manifold operations, such as retraction and Riemannian gradient computation.
-Moreover, we establish a direct connection between our construction and the score function in diffusion models of the data distribution.
-This connection allows us to leverage well-studied parameterizations, efficient training procedures, and even pretrained score networks from the diffusion model literature to perform optimization.
-Building on this foundation, we propose two efficient inference-time algorithms—Denoising Landing Flow (DLF) and Denoising Riemannian Gradient Descent (DRGD)—and provide theoretical guarantees for both feasibility (approximate manifold adherence) and optimality (small Riemannian gradient norm).
-Finally, we demonstrate the effectiveness of our approach on finite-horizon reference tracking tasks in data-driven control, highlighting its potential for practical generative and design applications.
+Our key idea is a link function that ties the data distribution to the geometric quantities needed for optimization: its gradient and Hessian recover the projection onto the manifold and its tangent space in the small-noise regime.
+This construction is directly connected to the score function in diffusion models, allowing us to leverage well-studied parameterizations, efficient training procedures, and even pretrained score networks from the diffusion model literature to perform optimization.
+On top of this foundation, we develop two efficient inference-time algorithms for optimization over data manifolds: Denoising Landing Flow (DLF) and Denoising Riemannian Gradient Descent (DRGD).
+We provide theoretical guarantees for approximate feasibility (manifold adherence) and optimality (small Riemannian gradient norm).
+We demonstrate the effectiveness of our approach on finite-horizon reference tracking tasks in data-driven control, illustrating their potential for practical generative and design applications.
 1 Introduction
@@
 (2018), the control of rigid bodies Duong et al.
-(2024), as well as sensor network localization Patwari & Hero (2004) and many others.
-Compared to general constrained optimization, Riemannian optimization promises the advantage of exploiting the natural geometry of the problem, producing feasible iterates and increased numerical robustness Boumal (2023).
-In contrast to the above classical setup, in this work we focus on the setting where the manifold $\mathcal{M}$ is given implicitly through a measure $\mu=\mu_{\mathrm{data}}$ supported on $\mathcal{M}$.
-This perspective is especially relevant in view of the manifold hypothesis (Loaiza-Ganem et al., 2024), which posits that many real-world data sets lie (approximately) on a manifold with dimension much smaller than that of the ambient space Fefferman et al.
+(2024), as well as sensor network localization Patwari and Hero (2004) and many others.
+Compared to general constrained optimization, Riemannian optimization promises the advantage of exploiting the natural geometry of the problem, producing feasible iterates and increased numerical robustness Boumal (2023), thus allowing for early termination and making it more suitable for real-time implementation.
+In contrast to the above classical setup, in this work we focus on the setting where the manifold $\mathcal{M}$ is given implicitly through a finite set of samples from an underlying population data distribution $\mu_{\mathrm{data}}$ supported on $\mathcal{M}$.
+This perspective is especially relevant in view of the data manifold hypothesis (Loaiza-Ganem et al., 2024), which posits that many real-world data sets lie (approximately) on a manifold with dimension much smaller than that of the ambient space Fefferman et al.
 (2016).
-Importantly, such data manifolds are not just low-dimensional geometric structures—they also capture rich semantic meaning.
+Importantly, such data manifolds are not just low-dimensional geometric structures – they also capture rich semantic meaning.
 For instance, the image manifold corresponds to photo-realistic images Pope et al.
-, the system behavior manifold to dynamically feasible input-output trajectories Willems & Polderman (1997), while the manifold of airfoils represents aerodynamically viable shapes Zheng et al.
+, the system behavior manifold to dynamically feasible input-output trajectories Willems and Polderman (1997), while the manifold of airfoils represents aerodynamically viable shapes Zheng et al.
 (2025).
-The optimization problem (1) in this setting thus encompasses a broad class of modern tasks, including airfoil design (Chen et al., 2025), reinforcement learning (Lee & Choi, 2025), and Bayesian inverse problems (Chung et al., 2022b).
-Driven by the growing demand in generative AI and data-driven design, this calls for a paradigm shift: moving from optimization over explicitly known manifolds, where classical Riemannian optimization applies, to optimization over data manifolds that are accessible only implicitly through samples.
-In this data-driven regime, methods from classical Riemannian optimization cannot be applied directly, since they rely on explicit manifold operations such as tangent-space projection, retraction, or exponential maps (Boumal et al., 2019; Boumal, 2023).
-While there has been extensive research on manifold learning Meilă & Zhang (2024); Lin & Zha (2008); Cayton et al.
-(2005); Belkin & Niyogi (2005) in the past, none of these works addressed (1) from an optimization point of view and focused instead on learning the manifold geometry.
-Moreover, the non-parametric nature of these methods limits their applicability to very low-dimensional manifolds and cannot exploit the successful inductive bias and the exceptional representation power of modern neural networks.
-In this work, we propose a data-driven approach to recover the fundamental operations needed for optimization on manifolds.
-Starting from the data distribution $\mu_{\mathrm{data}}$, we smooth it with a Gaussian kernel to obtain
+The optimization problem (1) in this implicit setting thus encompasses a broad class of modern tasks, including airfoil- and ship hull design (Chen et al., 2025; Bagazinski and Ahmed, 2023), additive manufacturing Peng et al.
+(2023), reinforcement learning (Lee and Choi, 2025), and Bayesian inverse problems (Chung et al., 2022b).
+In this data-driven regime, methods from classical smooth Riemannian optimization cannot be applied directly, since they rely on explicit manifold operations such as tangent-space projection, retraction, or exponential maps (Boumal et al., 2019; Boumal, 2023).
+Graph-based Riemannian optimization has recently been proposed in (Wang et al., 2025), but the setup is quite different, as to be elaborated in Section 1.2.
+Furthermore, while there has been extensive research on manifold learning Meilă and Zhang (2024); Lin and Zha (2008); Cayton and others (2005); Belkin and Niyogi (2005) in the past, none of these works addressed (1) from an optimization point of view and focused instead on learning the manifold geometry, rather than representation models that are suitable to be included as a constraint in an optimization problem.
+Instead, in this work, we propose a novel data-driven approach to recover the fundamental operations needed for optimization on manifolds: Starting first from the population data distribution $\mu_{\mathrm{data}}$, we smooth it with a Gaussian kernel to obtain
 $p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu_{\mathrm{data}},$ (2)
 and define the associated link function
-$\ell_{\sigma}(x)=\frac{1}{2}\lVert x\rVert^{2}-\sigma^{2}\log p_{\sigma}(x).$ (3)
+$\ell_{\sigma}(x)=\frac{1}{2}\lVert x\rVert^{2}+\sigma^{2}\log p_{\sigma}(x)\,,\quad x\in\mathbb{R}^{d}\,.$ (3)
 We show that, as the smoothing parameter $\sigma$ decreases, the gradient $\nabla\ell_{\sigma}$ recovers the projection back to the manifold, while the Hessian $\nabla^{2}\ell_{\sigma}$ recovers the projection onto its tangent space.
-These results reveal that core ingredients of Riemannian optimization—such as retraction and gradient computation—can be implemented directly from the derivative information of $\ell_{\sigma}$, even when the manifold itself is only given implicitly through data.
-Given the above novel theoretical findings, a key practical challenge is how to access (even approximately) the gradient $\nabla\ell_{\sigma}$ and Hessian $\nabla^{2}\ell_{\sigma}$ when only samples from $\mu_{\mathrm{data}}$ are available.
-A crucial observation is that $p_{\sigma}$ coincides with the marginal distribution of the Variance-Exploding SDE (VE-SDE), with analogous arguments for VP-SDE and DDPM formulations (Ho et al., 2020; Song et al., 2020).
-In this setting, the gradient of $\log p_{\sigma}$ – commonly referred to as the score function – plays a central role.
-In diffusion models, this score is parameterized by a neural network and learned directly from samples of $\mu_{\mathrm{data}}$, typically via denoising score matching (Vincent, 2011; Song et al., 2020).
-It is therefore natural to adopt the same methodology here: a neural network can be trained to represent $\nabla\ell_{\sigma}$, while the Hessian $\nabla^{2}\ell_{\sigma}$ can be recovered by computing its Jacobian.
+These results reveal that core ingredients of Riemannian optimization – such as retraction and gradient computation – can be implemented directly from the derivative information of $\ell_{\sigma}$, which itself can be constructed from $\mu_{\operatorname{data}}$.
+Given the above novel theoretical findings, a key practical challenge is how to access (even approximately) the gradient $\nabla\ell_{\sigma}$ and Hessian $\nabla^{2}\ell_{\sigma}$, when, in real-world applications, only samples from $\mu_{\mathrm{data}}$ are available.
+A crucial observation is that $p_{\sigma}$ coincides with the marginal distribution of the Variance-Exploding SDE (VE-SDE) [^1] Analogous extensions to VP-SDE and DDPM can also be similarly derived..
+Note that the gradient of $\log p_{\sigma}$, i.e.
+the score function, is central to diffusion models.
+It is parameterized by a neural network and learned directly from samples of $\mu_{\mathrm{data}}$, typically via denoising score matching (Vincent, 2011).
+We can hence exploit this well-developed toolkit for a practical implementation of our idea: Let $\texttt{s}(x,\sigma)$ be a pretrained score network.
+We use $\texttt{v}(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ to approximately represent $\nabla\ell_{\sigma}$, while the Hessian $\nabla^{2}\ell_{\sigma}$ can be recovered by computing its Jacobian.
 This approach offers two key advantages: (i) strong inductive biases can be incorporated through the neural network parameterization, and (ii) efficient, well-established training techniques from the diffusion model literature can be directly leveraged.
-Taken together, the theoretical link between $\ell_{\sigma}$ and manifold geometry, and the practical machinery for learning the score, form the foundation of the paradigm shift: from classical Riemannian optimization with explicit manifold knowledge to a data-driven framework where geometry is recovered from samples—thus enabling principled manifold optimization in generative and design-driven applications.
-Building on these insights, we propose two algorithms for optimization over data manifolds – denoising landing flow (DLF) and denoising Riemannian gradient descent (DRGD) – to the best of our knowledge, the first such framework in the literature.
-The prerequisite is access to an approximate function $s_{\theta}^{\sigma}\simeq\nabla\ell_{\sigma}$, which can be learned from data using standard diffusion model parameterization and training given samples from $\mu_{\mathrm{data}}$.
-  • Our first method directly employs $s_{\theta}^{\sigma}$ and its Jacobian (with $\sigma$ fixed as a small constant) to substitute the corresponding manifold operations in Riemannian gradient descent: each iteration computes a Riemannian gradient step and then applies retraction to (approximately) return to the manifold.
-  • Since learning the score is inevitably imperfect when $\sigma$ is small, we also design a complementary landing-type algorithm that avoids staying close to the feasible set at every step.
-Instead, it augments the target function $f$ with the regularization term $\ell_{\sigma}$ and performs gradient descent on this regularized objective.
-For both algorithms, we establish non-asymptotic convergence guarantees to approximate stationary points, measured by small Riemannian gradient norm, as $\sigma\to 0$.
+Taken together, the theoretical link between $\nabla\ell_{\sigma}$ and manifold geometry, and the practical machinery for learning the score, form the foundation of the paradigm shift: from classical smooth Riemannian optimization with explicit manifold knowledge to a data-driven framework where geometry is recovered from samples—thus enabling principled manifold optimization in generative and design-driven applications.
+Building on these insights, we propose the first score-based framework for optimization over data manifolds.
+We derive two algorithms: denoising landing flow (DLF) and denoising Riemannian gradient descent (DRGD).
+Both rely on an approximate score network $\texttt{s}(x,\sigma)\approx\nabla\log p_{\sigma}$ learned from samples of $\mu_{\mathrm{data}}$ using standard diffusion-model training.
+DRGD uses the learned gradient and its Jacobian (for a small, fixed $\sigma$) to mimic Riemannian gradient descent with retraction, while DLF performs gradient flow on a penalized objective, relaxing feasibility at intermediate iterates.
+We establish non-asymptotic convergence to approximate stationary points as $\sigma\to 0$.
 Importantly, our methods require only inference of the neural network and gradients with respect to its inputs—not with respect to the network parameters.
@@
 1.1 Our contributions
-We summarize our contributions and give the outline of the paper as follows.
-  • Data-driven recovery of manifold operations.
-By introducing the link function $\ell_{\sigma}$, we develop strategies to recover core manifold operations directly from data.
-In Section 3, by leveraging the score function from diffusion models to train and parameterize $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$, we bridge the gap between the requirements of classical Riemannian optimization (which assumes explicit manifold knowledge) and the emerging need to optimize over data manifolds that are only implicitly available.
-  • First algorithms for optimization over data manifolds.
-We propose two algorithms – denoising landing flow (DLF) via (8) in Section 4 and denoising Riemannian gradient descent (DRGD) via (12) in Section 5to the best of our knowledge, the first in the literature – that exploit operations enabled by a pretrained score function.
+We next summarize our main contributions and outline the structure of the paper.
+  • Link function and data-driven manifold operations.
+Building on the smoothed data distribution $p_{\sigma}$ and the associated link function $\ell_{\sigma}$ introduced above, we show in Section 3 that its gradient and Hessian recover, in the small-$\sigma$ regime, the projection onto the manifold and its tangent space.
+Further, with the help of score learning in diffusion model training, we bridge the gap between the requirements of classical Riemannian optimization (which assumes explicit manifold knowledge) and the emerging need to optimize over data manifolds that are only implicitly available.
+  • First score-based algorithms for optimization over data manifolds.
+We propose two algorithms – denoising landing flow (DLF) in Section 4 and denoising Riemannian gradient descent (DRGD) in Section 5 – to the best of our knowledge, the first in the literature that exploit operations enabled by a pretrained score function.
 Our methods require only inexpensive inference-time queries and back-propagation with respect to the input of the neural network, making them computationally efficient and readily applicable when pretrained scores are available.
-  • Rigorous non-asymptotic guarantees.
-We provide the first non-asymptotic convergence analysis for optimization over data manifolds.
-Our main results (Theorem 3 and Theorem 5) ensure both approximate feasibility (output is close to the manifold) and approximate optimality (small Riemannian gradient norm).
-A key technical contribution is a novel non-asymptotic analysis of how $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$ approximate the true manifold operations when $\sigma$ is small but non-zero (Theorem 1).
-Finally in Section 6 we validate our findings by numerically simulating the proposed flow on well-known manifolds from Riemannian optimization literature and data-driven optimal control for reference trajectory tracking.
-Our finding are that we can generate feasible points on the manifold with objective values far lower than the ones available in the training set – see Figure 1 – and shows the effectiveness of exploiting the strong inductive biases of modern deep learning for the classical field of constrained optimization.
+  • Non-asymptotic guaranties.
+We establish non-asymptotic convergence guarantees for both algorithms, showing approximate feasibility (outputs close to the data manifold) and approximate optimality (small Riemannian gradient norm) as $\sigma\to 0$ (Theorems 3 and 5).
+A key technical ingredient is a uniform control of how $\nabla\ell_{\sigma}$ and $\nabla^{2}\ell_{\sigma}$ approximate the ideal manifold operations at finite $\sigma$ (Theorem 1).
+Finally, in Section 6 we validate our approach on classical Riemannian optimization benchmarks and a data-driven optimal control problem for reference trajectory tracking.
+We demonstrate that our methods can generate feasible points on the manifold with objective values significantly lower than those observed in the training data (Figure 1), illustrating how strong inductive biases of modern deep networks can be harnessed for constrained optimization.
 Figure 1: Optimized trajectory (orange) on the system trajectory manifold for the unicycle car model that is desired to track the reference trajectory (black, thin).
@@
 1.2 Related Work
-Riemannian Optimization.
+In this section we briefly give an overview over the related work. More connections of our approach to previous work can be found in Appendix A.
+Smooth Riemannian Optimization.
 Riemannian optimization (RO) was originally developed under the assumption that the constraint manifold is explicitly known, either through closed-form descriptions such as matrix manifolds or via nonlinear equality constraints (Sato, 2021; Boumal, 2023).
@@
 Such algorithms guarantee that the iterates remain on the manifold at every iteration and are thus referred to as feasible methods.
-In contrast, infeasible methods, where the iterates are not constrained to stay on $\mathcal{M}$, have also been studied—for example, augmented Lagrangian approaches (Xie & Wright, 2021).
+In contrast, infeasible methods, where the iterates are not constrained to stay on $\mathcal{M}$, have also been studied – for example, augmented Lagrangian approaches (Xie and Wright, 2021).
 However, these typically involve solving complicated optimization subproblems in each inner loop, making them computationally expensive in practice.
 More recently, a new class of landing-type algorithms has emerged: instead of enforcing feasibility at every step, they regularize the objective function with the distance to the manifold and then perform gradient flow or descent on this regularized objective.
-Such methods have demonstrated strong empirical performance, offering a promising alternative to classical approaches (Ablin & Peyré, 2022; Schechtman et al., 2023).
+Such methods have demonstrated strong empirical performance, offering a promising alternative to classical approaches (Ablin and Peyré, 2022; Schechtman et al., 2023).
+Optimization over graph
+Recently, Wang et al.
+(2025) propose a graph-based strategy for manifold denoising, i.e.
+given observations sampled from the data manifold, for a new query point, find the closest point on the data manifold.
+The mixed-order optimization algorithm in this work consists of two types of updates: (1) By extending the concept of tangent space in classical smooth manifolds to the graph setting, first-order steps can be taken over the tangent bundle graph; (2) By enumerating the points connected by the “zero-order edge”, zero-th order steps can be taken to avoid local minima and achieve a strong result of convergence to global minima.
+We highlight that their setup is quite different from ours: (i) Their method is purely non-parametric and hence cannot exploit the recent development of deep learning; (ii) The proposed mix-order algorithm is more like a discrete search method.
+As an example, their strong global convergence result is built on the existence of zero-order edges, which “teleport” the iterate out of the local minima automatically.
+This is not the typical behavior of continuous optimization method, either deterministic or stochastic; (iii) Their method is purely designed for the denoising task while our focus is for minimizing general smooth objectives over data manifold.
+We highlight that due to the requirement of their analysis, for different objectives, the zero-order edges need to be reconstructed to avoid local minima.
 Manifold Learning.
@@
 The task of uncovering this structure is commonly referred to as manifold estimation or manifold learning, with a large body of work devoted to this goal.
-Classical approaches include Isomap (Tenenbaum et al., 2000), Laplacian Eigenmaps (Belkin & Niyogi, 2003), and Locally Linear Embedding (LLE) (Roweis & Saul, 2000), Diffusion Map (Coifman & Lafon, 2006), among many subsequent developments, e.g.
-(Zhou et al., 2020).
-We defer the discussion on diffusion-model-related manifold learning literature to the next paragraph.
+  • Non-parametric approaches include Isomap (Tenenbaum et al., 2000), Laplacian Eigenmaps (Belkin and Niyogi, 2003), and Locally Linear Embedding (LLE) (Roweis and Saul, 2000), Diffusion Map (Coifman and Lafon, 2006), among many subsequent developments, e.g.
+(Zhou et al., 2020; Yao et al., 2023).
+  • Parametric manifold learning techniques such as generative adversarial networks (GAN) and various types of autoencoders (AE) are used to estimate data manifolds $\mathcal{M}$ by learning a map $\psi:\mathcal{Z}\to\mathcal{M}$ that parametrizes $\mathcal{M}$ through some lower-dimensional “latent space” $\mathcal{Z}$ and which can be interpreted as a learned coordinate chart on the former Goodfellow et al.
+(2014); Berahmand et al.
+(2024).
+In the case of GAN this map corresponds to the generator, while for AE it corresponds to the decoder.
+For both GAN and AE, it has been noted that $\psi$ is not always a local chart in the differential-geometric sense Lee et al.
+(2022) and an additional line of work considered enforcing the latter property for the decoder through regularization or architectural constraints Sorrenson et al.
+(2023); Kumar et al.
+(2020).
+Manifold learning flows (M-flows) Brehmer and Cranmer (2020) address this problem by parametrizing $\psi$ in such a way that $\psi$ is injective and $\psi\circ\psi^{-1}$ can be efficiently computed and have been used for manifold learning and density estimation.
+Having obtained a chart $\psi$, one can locally trivialize the problem equation 1 to $\min_{\mathcal{Z}}f\circ\psi$, variants of which have also been known under the name of latent space optimization Tripp et al.
+(2020); Notin et al.
+(2021).
+We provide a detailed comparison of this paragraph to our work in Appendix A.
+Comparison with Pull-back Gradient Flow.
+When the map $\psi$ is learned, it induces a pull-back metric on the latent space $\mathcal{Z}$.
+Smooth Riemannian optimization algorithms on $\mathcal{M}$ can then be equivalently formulated in $\mathcal{Z}$; for instance, Riemannian gradient flow on $\mathcal{M}$ corresponds to a pull-back gradient flow on $\mathcal{Z}$.
+While we are not aware of prior work that explicitly exploits this idea to address optimization over data manifolds, we provide a comparison with this approach in Section A.2 and highlight several nontrivial challenges that should not be overlooked.
 Diffusion models.
 Diffusion models have achieved remarkable success in generative modeling, where the goal is to generate new samples consistent with an underlying data distribution $\mu_{\mathrm{data}}$.
-A central ingredient of these methods is the learning of the score function—the gradient of the log-density of the diffused distribution $p_{\sigma}$ Tang & Zhao (2025); Song et al.
+A central ingredient of these methods is the learning of the score function – the gradient of the log-density of the diffused distribution $p_{\sigma}$ Tang and Zhao (2025); Song et al.
 (2020).
@@
 affine subspaces), the Jacobian of the score scaled by the diffusion temperature asymptotically approximates the projection of the manifold onto the normal space and used it to study the geometric phases of diffusion models.
-Recent work on the statistical complexity of diffusion models under the manifold hypothesis provides indirect evidence that these models capture geometric information about the underlying data manifold (Oko et al., 2023; Tang & Yang, 2024).
+Recent work on the statistical complexity of diffusion models under the manifold hypothesis provides indirect evidence that these models capture geometric information about the underlying data manifold (Oko et al., 2023; Tang and Yang, 2024).
 In particular, the sample complexity required to learn the data distribution $\mu_{\mathrm{data}}$ depends only on the intrinsic dimension of the manifold, rather than on the ambient dimension.
@@
 This ensures (approximate) feasibility at the final step and guarantees that the optimization process remains semantically meaningful, which is not guaranteed by the sampling formulation, as discussed below.
-Across the literatures Classifier(-Free) Guidance in Diffusion Models (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) and the Plug-and-Play Framework in Bayesian Inverse Problems (Venkatakrishnan et al., 2013; Laumont et al., 2022; Pesme et al., 2025; Graikos et al., 2022; Chung et al., 2022a), a unifying perspective is the task of sampling from a posterior distribution of the form
+Across the literatures Classifier(-Free) Guidance in Diffusion Models (Dhariwal and Nichol, 2021; Ho and Salimans, 2022) and the Plug-and-Play Framework in Bayesian Inverse Problems (Venkatakrishnan et al., 2013; Laumont et al., 2022; Pesme et al., 2025; Graikos et al., 2022; Chung et al., 2022a), a unifying perspective is the task of sampling from a posterior distribution of the form
 $p_{\mathrm{post}}\propto p_{\mathrm{pre}}\exp\left(-\tfrac{r}{\alpha}\right),$
-where $r$ denotes a cost function to be minimized—corresponding to our objective ff—and $\alpha>0$ is a temperature-like parameter (Domingo-Enrich et al., 2024).
+where $r$ denotes a cost function to be minimized, corresponding to our objective $f$, and $\alpha>0$ is a temperature-like parameter (Domingo-Enrich et al., 2024).
 In Classifier(-Free) Guidance, $r$ encodes a classifier signal (e.g., specified by a prompt) and in Bayesian inverse problems, $r$ is the negative log-likelihood of observations.
@@
 2 Preliminaries
-Here we introduce some preliminary notation and concepts to state our results. We refer the reader to the Appendices A and C for more details.
+Here we introduce some preliminary notation and concepts to state our results. We refer the reader to the Appendices B and D for more details.
 2.1 Manifolds and distance functions
@@
 For any point $p\in\mathcal{M}$ we denote by $\operatorname{T}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ and $\operatorname{N}_{p}\mathcal{M}\subseteq\mathbb{R}^{d}$ the tangent and normal spaces of $\mathcal{M}$ at $p$, respectively, and their orthogonal projections by $\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}$ and $\operatorname{P}_{\operatorname{N}_{p}\mathcal{M}}$.
-The squared distance function is defined by $\operatorname{d}(x)=\inf_{p\in\mathcal{M}}\frac{1}{2}\lVert x-p\rVert^{2}$.
-For a $C^{2}$-submanifold, then there exists a radius $\tau_{\mathcal{M}}>0$ such that every point in $x\in\mathcal{T}(\tau)=\bigcup_{p\in\mathcal{M}}B_{\tau}(p)$ has a unique projection $\pi(x)\in\mathcal{M}$ such that $\frac{1}{2}\lVert x-\pi(x)\rVert^{2}=\operatorname{d}(x)$ and $x-\pi(x)\in\operatorname{N}_{\pi(x)}\mathcal{M}$.
-Sets of the form $\mathcal{T}(\tau)$ will be called tubular neighborhoods of $\mathcal{M}$ with closure denoted by $\overline{\mathcal{T}}(\tau)$.
-One can prove that $\operatorname{d}$ and $\pi$ are differentiable on $\mathcal{T}(\tau_{\mathcal{M}})$ with $\operatorname{d}^{\prime}(x)=x-\pi(x)$ and $\pi^{\prime}(x)$ given explicitly in Appendix A.
-The maximal principal curvature of a manifold will be denoted by $\kappa_{\mathcal{M}}$.
-It always holds that $\tau_{\mathcal{M}}\leq 1/\kappa_{\mathcal{M}}$.
+For a $C^{2}$-submanifold, then there exists a radius $\tau_{\mathcal{M}}>0$ such that every point in $x\in\mathcal{T}(\tau_{\mathcal{M}})$, where
+$\mathcal{T}(\tau):=\{x\in\mathbb{R}^{d}\mid\operatorname{dist}(x,\mathcal{M})<\tau\}\,,\quad\operatorname{dist}(x,\mathcal{M})=\inf_{p\in\mathcal{M}}\lVert x-p\rVert\,$,
+has a unique projection $\pi(x)\in\mathcal{M}$.
+Sets of the form $\mathcal{T}(\tau)$ with $\tau\in(0,\tau_{\mathcal{M}})$ are called tubular neighborhoods of $\mathcal{M}$.
+The squared distance function is defined by $\operatorname{d}(x)=\frac{1}{2}\operatorname{dist}(x,\mathcal{M})^{2}$ and $\operatorname{d}$ and $\pi$ are both differentiable on $\mathcal{T}(\tau_{\mathcal{M}})$ with $\frac{1}{2}\lVert x-\pi(x)\rVert^{2}=\operatorname{d}(x)$ and $x-\pi(x)\in\operatorname{N}_{\pi(x)}\mathcal{M}$ and $\operatorname{d}^{\prime}(x)=x-\pi(x)$ and $\pi^{\prime}(x)$ given explicitly in Appendix B.
 Finally, for a function $f:\mathcal{M}\to\mathbb{R}$ we denote the Riemannian gradient by $\operatorname{grad}_{\mathcal{M}}f(p)$, which in the case of $f:\mathbb{R}^{d}\to\mathbb{R}$ is given by $\operatorname{grad}_{\mathcal{M}}f(p)=\operatorname{P}_{\operatorname{T}_{p}\mathcal{M}}\nabla f(p)$.
-We write
 2.2 The Stein score function and score-based diffusion models
 For a Borel probability measure we denote its Gaussian blurring by $p_{\sigma}=\mathcal{N}(0,\sigma^{2}I)\ast\mu$ for $\sigma>0$ with $p_{0}=\mu$.
-The score function $\nabla\log p_{\sigma}$ of $p_{\sigma}$ and its Jacobian allow for the following interpretation (see Jaffer & Gupta (1972), also Appendix B):
+The score function $\nabla\log p_{\sigma}$ of $p_{\sigma}$ and its Jacobian allow for the following interpretation (see Jaffer and Gupta (1972), also Appendix C):
 $x+\sigma^{2}\nabla\log p_{\sigma}(x)=\nabla\ell_{\sigma}(x)=\mathbb{E}\nu_{x,\sigma}\,,\text{\ \ }I+\sigma^{2}\nabla^{2}\log p_{\sigma}(x)=\nabla^{2}\ell_{\sigma}(x)=\frac{1}{\sigma^{2}}\operatorname{Cov}(\nu_{x,\sigma})\,$, (4)
@@
 Specifically in the so-called variance exploding (VE) diffusion scheme one seeks to learn $\nabla\log p_{\sigma}$ for different noise scales $\sigma$ via a neural network $\texttt{s}(\cdot,\sigma)$ by minimizing the conditional score matching loss $L_{\operatorname{CSM}}(\texttt{s}(\cdot,\sigma))$, which attains its unique minimum in $\texttt{s}(\cdot,\sigma)=\nabla\log p_{\sigma}$.
-Then $\texttt{s}(\cdot,\sigma)$ is used for sampling from $\mu$ by following a particular reverse-time SDE or ODE flow in the noise scale $\sigma$ (see Appendix C for more details).
+Then $\texttt{s}(\cdot,\sigma)$ is used for sampling from $\mu$ by following a particular reverse-time SDE or ODE flow in the noise scale $\sigma$ (see Appendix D for more details).
 3 Score as retraction and tangent space projection
@@
 for all $\sigma\in(0,\overline{\sigma})$ and $x\in\mathcal{T}(\tau)$.
-The proof is deferred to Appendix D and is based on a careful non-asymptotic estimate of the Laplace integral method.
-As a consequence and together with fact that $\pi^{\prime}(x)$ coincides with $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ for $x\in\mathcal{M}$ (see Appendix A) we obtain the following result.
+The proof is deferred to Appendix E and is based on a careful non-asymptotic estimate of the Laplace integral method.
+As a consequence and together with fact that $\pi^{\prime}(x)$ coincides with $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ for $x\in\mathcal{M}$ (see Appendix B) we obtain the following result.
 Corollary 2.
@@
 We stress that the expressions in (6) are defined for all $x\in\mathbb{R}^{d}$, whereas $\operatorname{d}$, $\pi$ and $\pi^{\prime}$ in are only sensible in a tubular neighborhood $\mathcal{T}$ of $\mathcal{M}$.
-Thus, in theory, a well-trained diffusion model score $\texttt{s}(\cdot,\sigma)$ and its Jacobian $\texttt{s}^{\prime}(\cdot,\sigma)$ allow us to approximate the closest-point projection $\pi(x)$ and the tangent space projection $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ as $\sigma\to 0$ arbitrarily well via its Tweedie score $x+\sigma^{2}\texttt{s}(x,\sigma)$ and its Jacobian, respectively.
+Thus, in theory, a well-trained diffusion model score $\texttt{s}(\cdot,\sigma)$ and its Jacobian $\texttt{s}^{\prime}(\cdot,\sigma)$ allow us to approximate the closest-point projection $\pi(x)$ and the tangent space projection $\operatorname{P}_{\operatorname{T}_{x}\mathcal{M}}$ as $\sigma\to 0$ arbitrarily well via the operator $\texttt{v}(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and its Jacobian, respectively.
 4 Denoising Riemannian gradient flow with landing
 In this section we show how to use the score function from Section 3 for Riemannian optimization of (1) for some smooth $f\in C^{1}(\mathbb{R}^{d})$.
-Assuming that we have access to a sufficiently accurate estimate of the Tweedie score in form of a vector function $s\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$ such that $s(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and
-$\lVert s(x)-\pi_{\sigma}(x)\rVert\leq\epsilon\text{\ and\ }\lVert s^{\prime}(x)-P_{\sigma}(x)\rVert\leq\epsilon\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$ (7)
+Assuming that we have access to a sufficiently accurate estimate of the score in form of a vector function $\texttt{v}\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$ such that $\texttt{v}(x)=x+\sigma^{2}\texttt{s}(x,\sigma)$ and
+$\lVert\texttt{v}(x)-\pi_{\sigma}(x)\rVert\leq\epsilon\text{\ and\ }\lVert\texttt{v}^{\prime}(x)-P_{\sigma}(x)\rVert\leq\epsilon\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$ (7)
 for some $\tau\in(0,\tau_{\mathcal{M}})$, we propose for $\eta\geq 0$ the denoising landing flow (DLF)
-$\dot{x}$ $=-s^{\prime}(x)\nabla f(s(x))+\eta(s(x)-x)\,.$ (8)
-In the exact case $s(x)=\pi_{\sigma}(x)$ and $s^{\prime}(x)=P_{\sigma}(x)$ (i.e. $\epsilon=0$ in (7)), flow (8) is the gradient flow
+$\dot{x}$ $=-\texttt{v}^{\prime}(x)\nabla f(\texttt{v}(x))+\eta(\texttt{v}(x)-x)\,.$ (8)
+In the exact case $\texttt{v}(x)=\pi_{\sigma}(x)$ and $\texttt{v}^{\prime}(x)=P_{\sigma}(x)$ (i.e. $\epsilon=0$ in (7)), flow (8) is the gradient flow
 $\dot{x}$ $=-\nabla F_{\sigma}^{\eta}(x)=-P_{\sigma}(x)\nabla f(\pi_{\sigma}(x))+\eta(\pi_{\sigma}(x)-x)\text{\ \ with\ \ }F_{\sigma}^{\eta}(x)=f(\pi_{\sigma}(x))+\eta\operatorname{d}_{\sigma}(x)$ (9)
 and the dynamics in (9) consists of two parts: An approximate projection $P_{\sigma}(x)\nabla f(\pi_{\sigma}(x))$ of the gradient $\nabla f(\pi_{\sigma}(x))$ and an approximate landing term $\eta(\pi_{\sigma}(x)-x)$ corresponding to the penalty function $\eta\operatorname{d}_{\sigma}(x)$.
-In the further case of $\sigma=0$ and $x(0)\in\mathcal{M}$ the flow (9) reduces to the ordinary Riemannian gradient flow, which has been extensively studied Helmke & Moore (2012); Ambrosio et al.
+In the further case of $\sigma=0$ and $x(0)\in\mathcal{M}$ the flow (9) reduces to the ordinary Riemannian gradient flow, which has been extensively studied Helmke and Moore (2012); Ambrosio et al.
 (2005).
@@
 […]
-with $H_{x}^{-1}$ a linear operator on $\operatorname{T}_{\pi(x)}\mathcal{M}$ given in Appendix A.
-In particular the two terms in (10) belong to $\operatorname{T}_{\pi(x)}\mathcal{M}$ and $\operatorname{N}_{\pi(x)}\mathcal{M}$, respectively, and are orthogonal to each other, implying that the distance between $x$ and $\mathcal{M}$ is non-increasing, which allows for perfect landing on the manifold via similar arguments as in Ablin & Peyré (2022); Schechtman et al.
-(2023), see Theorem 20 in Appendix E.3.
+with $H_{x}^{-1}$ a linear operator on $\operatorname{T}_{\pi(x)}\mathcal{M}$ given in Appendix B.
+In particular the two terms in (10) belong to $\operatorname{T}_{\pi(x)}\mathcal{M}$ and $\operatorname{N}_{\pi(x)}\mathcal{M}$, respectively, and are orthogonal to each other, implying that the distance between $x$ and $\mathcal{M}$ is non-increasing, which allows for perfect landing on the manifold via similar arguments as in Ablin and Peyré (2022); Schechtman et al.
+(2023), see Theorem 22 in Appendix F.3.
 For $\sigma>0$ or $\epsilon>0$, the two summands in (9) and (8) in general not perpendicular to each other and the landing is not exact.
@@
 Set $C=\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f)$.
-Suppose that for some $\epsilon>0$ and $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ with $\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\min\{\tau,\frac{2\tau}{1+C/\eta}\}$ the function $s$ satisfies (7).
+Suppose that for some $\epsilon>0$ and $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ with $\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\min\{\tau,\frac{2\tau}{1+C/\eta}\}$ the function v satisfies (7).
 Then for any $x(0)\in\mathcal{T}(\tau)$ the solution $x(t)$ to (8) exists for all $t\geq 0$ and is contained in $\mathcal{T}(\tau)$.
@@
 Remark 4.
-We can evaluate the right hand side of the flow (8) in a single forward-backward pass of the network $s$.
-Namely, given an input $x$, we compute and store $p=s(x)$ by a forward pass of $s$, while keeping the computational graph of $s(x)$.
-Then we evaluate $v=\nabla f(p)$ and build the computational graph of $y=\left\langle s(x),v\right\rangle$, while detaching $v$.
-Finally we backpropagate on $x$ in $y$ to obtain $s^{\prime}(x)v=s^{\prime}(x)\nabla f(s(x))$.
+We can evaluate the right hand side of the flow (8) in a single forward-backward pass of the network v.
+Namely, given an input $x$, we compute and store $p=\texttt{v}(x)$ by a forward pass of v, while keeping the computational graph of $\texttt{v}(x)$.
+Then we evaluate $g=\nabla f(p)$ and build the computational graph of $y=\left\langle\texttt{v}(x),g\right\rangle$, while detaching $g$.
+Finally we backpropagate on $x$ in $y$ to obtain $\texttt{v}^{\prime}(x)g=\texttt{v}^{\prime}(x)\nabla f(\texttt{v}(x))$.
 5 Denoising Riemannian gradient descent
@@
 (2008); Boumal (2023)
-$x_{k+1}=s(x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k}))\,$, (12)
+$x_{k+1}=\texttt{v}(x_{k}-\gamma_{k}\texttt{v}^{\prime}(x_{k})\nabla f(x_{k}))\,$, (12)
 which we term the denoising Riemannian gradient descent (DRGD).
-Here $s$ acts as an approximate retraction and $s$ as an approximate projection onto the tangent space.
+Here v acts as an approximate retraction and $\texttt{v}^{\prime}$ as an approximate projection onto the tangent space.
 We obtain the following convergence result for this algorithm.
@@
 $L_{0}=8C\left(2(\frac{3}{\tau_{\mathcal{M}}}+\tau M)+\frac{1}{\tau_{\mathcal{M}}}\right)+2L\,.$
-Suppose that for some $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ and $\epsilon>0$ with $\epsilon^{\prime}:=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$ the function $s$ satisfies (7) and that the step-size $\gamma_{k}$ is constrained by $\gamma_{k}\in[\gamma_{\min},\gamma_{\max}]$ with
+Suppose that for some $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$ and $\epsilon>0$ with $\epsilon^{\prime}:=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$ the function v satisfies (7) and that the step-size $\gamma_{k}$ is constrained by $\gamma_{k}\in[\gamma_{\min},\gamma_{\max}]$ with
 $0<\gamma_{\min}<\gamma_{\max}<\min\left\{\frac{2}{L_{0}},\frac{\tau}{C(4+\tau)}\right\}\,.$
@@
 Remark 6.
-Note that both Theorem 3 and Theorem 5 require the rather strong $L^{\infty}$-approximation assumption (7) on the Tweedie score $s$ and its Jacobian.
+If we set $\epsilon=0$ and $\sigma=0$, we recover, up to constants, the classical result on iterates of the Riemannian gradient descent with known manifold $\mathcal{M}$ and non-convex objective $f$ with Lipschitz gradient (Boumal, 2023, Corollary 4.9) (see Appendix A.3).
+Remark 7.
+Note that both Theorem 3 and Theorem 5 require the rather strong $L^{\infty}$-approximation assumption (7) on v and its Jacobian.
 The analysis under a weaker $L^{2}$-bound is out of score for this paper and left for future work.
@@
 (2008); Boumal (2023).
-Specifically we consider the orthogonal group manifold $\mathcal{M}=\operatorname{O}(n)\subseteq\mathbb{R}^{n\times n}$ with $\mu=\operatorname{Vol}_{\mathcal{M}}$ being the uniform volume measure and Brockett’s cost function Helmke & Moore (2012) defined by
+Specifically we consider the orthogonal group manifold $\mathcal{M}=\operatorname{O}(n)\subseteq\mathbb{R}^{n\times n}$ with $\mu=\operatorname{Vol}_{\mathcal{M}}$ being the uniform volume measure and Brockett’s cost function Helmke and Moore (2012) defined by
 $\min_{X\in\operatorname{O}(n)}f(X):=\operatorname{tr}(AXQX^{\top})\,$,
 where $A,Q\in\mathbb{S}^{n\times n}$ are given.
-We consider the cases $n\in\{10,20\}$ and assume that we are given a set $\mathcal{D}_{\operatorname{train}}\subseteq\mathcal{M}$ of $N_{\operatorname{data}}=20000$ data points from $\mu$ and train the score function $s$ with denoising score matching (see Appendix C for diffusion models and Appendix G.1.1 for implementation details).
-In Figure 3 (left) (in Appendix G.1.2 for space reasons) we compare the evolution of the objective value of our approximation of (8) to the exact landing flow (10) for different noise levels $\sigma>0$ and dimension $n$.
+We consider the cases $n\in\{10,20\}$ and assume that we are given a set $\mathcal{D}_{\operatorname{train}}\subseteq\mathcal{M}$ of $N_{\operatorname{data}}=20000$ data points from $\mu$ and train the score function s with denoising score matching (see Appendix D for diffusion models and Appendix H.1.1 for implementation details).
+In Figure 2 (left) we compare the evolution of the objective value of our approximation of (8) to the exact landing flow (10) for different noise levels $\sigma>0$ and dimension $n$.
 We observe that we can obtain objective values with cost lower than the best possible point in the training set and that the accuracy improves as $\sigma\to 0$.
+Refer to caption Refer to caption
+Figure 2: Objective value vs.
+flow time $t$ for the orthogonal manifold for $n=10$ (left, different $\sigma>0$) and for $n=20$ (right, $\sigma=0.05$).
+Here “exact landing” refers to equation 8 with exact operations $\texttt{v}=\pi$ and $\texttt{v}^{\prime}=\pi^{\prime}$.
 6.2 Reference tracking via data-driven control
@@
 […]
-Under certain smoothness assumptions on the dynamics $\bm{f}$ and $\bm{g}$ the set $\mathcal{M}_{\operatorname{IO}}$ is a (non-compact) embedded smooth submanifold of $(\mathbb{R}^{n_{u}})^{N_{h}}\times(\mathbb{R}^{n_{y}})^{N_{h}+1}$.
+Under smoothness assumptions on the dynamics $\bm{f}$ and $\bm{g}$, the set $\mathcal{M}_{\operatorname{IO}}$, being a graph of a smooth map, is a (non-compact) embedded smooth submanifold of $(\mathbb{R}^{n_{u}})^{N_{h}}\times(\mathbb{R}^{n_{y}})^{N_{h}+1}$.
 The problem (14) is ubiquitous in receding horizon control applications such as model predictive control (MPC) and used for e.g.
@@
 given by (white) noise Ljung (1999).
-In particular the so-called system behavior manifold $\mathcal{M}_{\operatorname{IO}}$ Willems & Polderman (1997) is given by samples from a distribution $\mu$ on its in- and outputs and fits precisely into our framework of data-driven Riemannian optimization (1).
-We test our proposed denoising Riemannian gradient descent on two classical systems from the control domain: The discretized double pendulum system and the unicycle car model LaValle (2006) (see Appendix G.2.1 detailed information on the systems and the particular choice of $\bm{r}$, $R$ and $Q$ in (15)), each on a horizon of $N_{h}=100$.
-To apply our proposed methods, we train a diffusion model (see Appendix G.2.2 for implementation details) on the measured trajectories $\mathcal{D}_{\operatorname{train}}$ and solve (14) via the denoising Riemannian gradient descent to obtain a solution $(\bm{u}^{*},\bm{y}^{*})$.
+In particular the so-called system behavior manifold $\mathcal{M}_{\operatorname{IO}}$ Willems and Polderman (1997) is given by samples from a distribution $\mu$ on its in- and outputs and fits precisely into our framework of data-driven Riemannian optimization (1).
+A similar setup with observable state $\bm{y}=\bm{x}$ has been considered in the domain of reinforcement learning Janner et al.
+(2022).
+We test our proposed denoising Riemannian gradient descent on two classical systems from the control domain: The discretized double pendulum system and the unicycle car model LaValle (2006) (see Appendix H.2.1 detailed information on the systems and the particular choice of $\bm{r}$, $R$ and $Q$ in (15)), each on a horizon of $N_{h}=100$.
+To apply our proposed methods, we train a diffusion model (see Appendix H.2.2 for implementation details) on the measured trajectories $\mathcal{D}_{\operatorname{train}}$ and solve (14) via the denoising Riemannian gradient descent to obtain a solution $(\bm{u}^{*},\bm{y}^{*})$.
 As initial values we take the trajectories from the training set that minimize the objective cost $f$, i.e.
@@
 To account for this deviation we back-test our generated input trajectory $\bm{u}^{*}$ by implementing it on the true underlying system (13) to obtain the real output $\bm{y}^{\operatorname{true}}$.
-Results and discussion: In Figure 4 (in Appendix G.2.3) we depict the evolution of the objective value w.r.t.
+Results and discussion: In Figure 4 (in Appendix H.2.3) we depict the evolution of the objective value w.r.t.
 iteration count and in Figure 3 we depict the final optimizing trajectories $\bm{y}^{*}$ and $\bm{y}^{\operatorname{true}}$.
@@
 Future work will consist of deriving error bounds for this flow when the denoising score is trained with a non-zero $L^{2}$-error as well as the study of more sophisticated classical Riemannian optimization algorithms such as Newton and trust region methods when using the approximate manifold operations with the trained score to accelerate convergence.
-Appendix A More on manifolds and distance functions
-In addition to the notation introducted in Section 2.1, we note that $\tau_{\mathcal{M}}$ is also the largest $\tau\geq 0$ such that the map $\{(p,v)\in\operatorname{N}\mathcal{M}\mid\lVert v\rVert<\tau\}\to\mathbb{R}^{d}:(p,v)\mapsto p+v$ is a diffeomorphism.
+Acknowledgments and Disclosure of Funding
+We thank the reviewers for their constructive suggestions.
+This work was supported as a part of NCCR Automation, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 51NF40_225155).
+Riccardo De Santi is supported by the ETH AI Center through an ETH AI Center doctoral fellowship.
+Appendix A Connections and comparisons to prior work
+A.1 Connection to denoising autoencoders
+In the following we show how our results connect to learning denoising auto-encoders. The following theorem is shown in Alain and Bengio (2014):
+Theorem 8 (Theorem 1 in Alain and Bengio (2014)).
+Let $\mu$ be a probability measure on $\mathbb{R}^{d}$ and consider for $\sigma>0$ the reconstruction problem
+$\min_{r}L_{\operatorname{DAE}}(r)\quad\text{for}\quad L_{\operatorname{DAE}}(r)=\mathbb{E}_{x\sim\mu\,,\;\epsilon\sim\mathcal{N}(0,I)}\|x-r(x+\sigma\epsilon)\|^{2}\,.$
+If $\mu$ is absolutely continuous with (Lebesgue) density $p$, the optimal reconstruction function $r_{\sigma}^{*}$ satisfies
+$r_{\sigma}^{*}(x)=\frac{\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}p(x-\sigma\epsilon)(x-\sigma\epsilon)}{\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}p(x-\sigma\epsilon)}\overset{\sigma\to 0}{=}x+\sigma^{2}\nabla\log p(x)+o(\sigma^{2})$
+at any $x\in\mathbb{R}^{d}$ with $p(x)>0$.
+To see how this result relates to Theorem 1, note that function $r_{\sigma}^{*}$ here is precisely our “approximate projection” $\pi_{\sigma}$, since the first above expression corresponds to the posterior of observing $x$ under the convolution of the prior $p$ with $\mathcal{N}(0,\sigma^{2}I)$, i.e.
+in our notation
+$r_{\sigma}^{*}(x)=\mathbb{E}\nu_{x,\sigma}=\pi_{\sigma}(x).$
+Therefore, if $\mu$ has the Lebesgue-density $p$ (i.e. the data manifold would have full dimension $k=d$), then Alain and Bengio (2014) show that
+$x+\sigma^{2}\nabla\log p_{\sigma}(x)\overset{\sigma\to 0}{=}x+\sigma^{2}\nabla\log p(x)+o(\sigma^{2}).$
+In particular, if $p(x)>0$, then the last equation states that
+$\lim_{\sigma\to 0}x+\sigma^{2}\nabla\log p_{\sigma}(x)=x.$
+Theorem 1 on the other hand deals with the case $k<d$, i.e.
+when the manifold is truly lower-dimensional.
+In this case, $\mu$ has no density $p$ w.r.t.
+the Lebesgue measure on $\mathbb{R}^{d}$, and it holds that
+$\lim_{\sigma\to 0}x+\sigma^{2}\nabla\log p_{\sigma}(x)=\pi(x),$
+where $\pi$ denotes the projection operation onto the data manifold $\mathcal{M}$.
+While Alain and Bengio (2014) does consider manifold support in their discussion (see (Alain and Bengio, 2014, Section 3.4, Figure 5)), note that the population data measure in this case is not absolute continuous w.r.t.
+the Lebesgue measure on $\mathbb{R}^{d}$, and hence $\log p(x)$ has no meaning.
+Consequently, results of Alain and Bengio (2014), e.g.
+the expansion above regarding $r_{\sigma}^{*}$, are not applicable.
+Theorem 1 can therefore be seen a (uniform) generalization of Alain and Bengio (2014) to an order of $\widetilde{O}(\sigma)$, establishing the asymptotic convergence of $r_{\sigma}^{*}$ and its Jacobian for $\sigma\to 0$ as
+$r_{\sigma}^{*}(x)=\pi(x)+\widetilde{O}(\sigma)\,,\quad(r_{\sigma}^{*})^{\prime}(x)=\pi^{\prime}(x)+\widetilde{O}(\sigma)\,.$
+A.2 Trivializations and natural gradients using parametric manifold learning
+In this section we compare our approach to latent space optimization.
+For this purpose, we first describe the latter problem in terms of trivializations of (1) using local parametrizations: Given a finite atlas $\mathcal{A}=(\psi_{i},\mathcal{U}_{i},\mathcal{V}_{i})_{i=1}^{m}$ of $\mathcal{M}$ with open $\mathcal{V}_{i}\subseteq\mathbb{R}^{k}$ and $\mathcal{U}_{i}\subseteq\mathcal{M}$ and charts $\psi_{i}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ it clearly holds
+$\min_{x\in\mathcal{M}}f(x)=\min_{i=1,\ldots,m}\min_{z\in\mathcal{V}_{i}}f(\psi_{i}(z))\,.$
+In the following we will focus on one inner subproblem by fixing a single chart $(\psi,\mathcal{U},\mathcal{V})$ and consider
+$\min_{x\in\mathcal{U}}f(x)=\min_{z\in\mathcal{V}}f(\psi(z))\,.$ (16)
+While for the above equality we only need $\psi$ to be surjective (i.e.
+$\psi(\mathcal{V})=\mathcal{U}$), for equivalence of first-order stationary points of both problems we need $\psi$ to be a submersion: To see this, note that stationarity of $x_{*}\in\mathcal{U}$ for the left hand side in equation 16 is equivalent to $\operatorname{grad}_{\mathcal{M}}f(x_{*})=\operatorname{P}_{\operatorname{T}_{x_{*}}\mathcal{M}}\nabla f(x_{*})=0$ and stationarity of any $z_{*}\in\mathcal{V}\cap\psi^{-1}(\{x_{*}\})$ for the right hand side is equivalent to $\psi^{\prime}(z_{*})^{\top}\nabla f(\psi(z_{*}))=\psi^{\prime}(z_{*})^{\top}\nabla f(x_{*})=0$.
+In particular, for these two notions to be equivalent a-priori for any $f$ and $x_{*}$, the columns of $\psi^{\prime}(z_{*})$ must span $\operatorname{T}_{x_{*}}\mathcal{M}$, which implies that $\psi$ is a submersion.
+In particular, if $\psi$ is a chart, then the stationary points of both sides in equation 16 are in a one-to-one correspondence and only then both problems are equivalent from an first-order optimization point of view.
+Two possible ways to solve the right hand side in (16) are the latent space gradient flow (LSGF) and the pull-back Riemannian gradient flow (PBRGF) given by
+$\dot{z}$ $=-\nabla(f\circ\psi)(z)=-\psi^{\prime}(z)^{\top}\nabla f(\psi(z))\,$, (17)
+$\dot{z}$ $=-G_{\psi}(z)\psi^{\prime}(z)^{\top}\nabla f(\psi(z))\,$, (18)
+respectively, where $G_{\psi}(z)=\psi^{\prime}(z)^{\top}\psi^{\prime}(z)$ is the metric tensor of $\psi$ at $z\in V$.
+Here PBRGF [^2] The Euler discretization of the continuous dynamics (18) is also known as natural gradient descent (Amari, 1998, Theorem 1) in the classical Riemannian optimization literature is obtained from LSGF via a change in inner product on the tangent space $\operatorname{T}_{\psi(z)}\mathcal{M}$ and is equivalent to the Riemannian gradient flow in the sense that for $x=\psi(z)$ it holds that
+$\dot{x}=-\operatorname{grad}f(x)\,.$
+While (17) and (18) are both possible approaches to solving (1) and have been applied in the past Pegios et al. ; Tripp et al. (2020), there are a few challenges present these methods:
+  • Parameterizing the full atlas of a manifold is hard: The atlas $\mathcal{A}$ of a manifold typically consists of multiple charts $(\psi_{i},U_{i},V_{i})$.
+The number of charts as well as their span across the manifold is typically unknown and hard to estimate a-priori.
+  • Intrinsic dimension of the manifold unknown: The dimension of the latent space $\mathcal{Z}$ must correspond to the manifold dimension, the estimation of which is not a trivial task.
+  • Generative model training: The implementation of (18) or (17) require $\psi$ to be invertible and the inverse coordinate map $\psi^{-1}$ to be well approximated by a neural network.
+It is not straightforward to align this requirement with the standard training objectives commonly used in generative models, such as the likelihood-based loss in VAEs, the denoising score-matching loss in diffusion models, or the adversarial loss in GANs.
+Further approaches such as M-flows Brehmer and Cranmer (2020) require $\psi$ to be essentially a normalizing flow, which is fundamentally incompatible with common architectures such as U-Nets or transformers.
+  • Expensive update rules for PBRGF: The update rule of PBRGF requires (i) computing the matrix $G_{\psi}(z)$, which in turn involves the explicit evaluation of the Jacobian $\psi^{\prime}(z)$, and (ii) inverting $G_{\psi}(z)$.
+Both operations become computationally expensive when $\psi$ is implemented by a neural network even for a moderate intrinsic dimension $k$.
+  • Limited latent space validity region: Since $\mathcal{Z}$ is representing the local coordinates of a single chart, the validity of the model $\psi$ is only given in a bounded domain of $\mathcal{Z}$.
+Optimizing outside of this domain requires changing charts, since otherwise the obtained latent points will have no meaning in the data-space.
+Estimating the validity domain of $\psi$ is difficult a-priori Notin et al.
+(2021).
+In the above light, the advantages of the algorithms DLF and DRGD described in Sections 4 and 5 are as follows:
+  • Easy to parameterize with a neural network: The central object in our approach is the projection operator onto the data manifold $\mathcal{M}$, which is provably $C^{2}$ in the tubular neighborhood of any $C^{3}$ manifold.
+Assuming $\mathcal{M}$ is compact, standard universal approximation results ensure that the target map $\pi$ can be well-approximated by a single neural network
+  • Dimension-agnostic algorithms: Both DLF and DRGD require no knowledge about the manifold dimension $k$.
+  • Alignment with denoising score matching: DLF and DRGD use an approximation of $\pi$ via the score function of $p_{\sigma}$ (up to scaling and a residual term of $x$), so the denoising score-matching loss used to train diffusion models is directly aligned with our target quantity.
+Our approach imposes no architectural constraints: in our experiments, for example, we successfully use both MLPs and U-Nets.
+  • Simple and efficient update rule The update rules (8) and (12) are highly compatible with modern hardware and software stacks and requires only a single forward–backward pass of the network per iteration (see Remark 4)
+  • Convergence guarantees in the embedding space: Theorem 3 and 5 provide explicit convergence analysis with guarantees on both optimality and feasibility, stated directly for the original data manifold in the data space, where convergence is semantically meaningful.
+A.3 Recovery of iterate complexity of Riemannian gradient descent
+In this section we show that in the case of $\sigma=\epsilon=0$, Theorem 5 recovers the corresponding result for the Riemannian gradient descent.
+To see this, let us recall the standard complexity guarantee for Riemannian gradient descent under a general smooth non-convex objective $f$: In (Boumal, 2023, Corollary 4.9) it is stated that (taking $R_{x}(v)=\pi(x+v)$ as the retraction defined on some ball $B_{r}(x)\cap T_{x}\mathcal{M}$, where for us $r>0$ can be picked uniformly over $x\in\mathcal{M}$) if $f\circ\pi$ satisfies the pullback $L$-Lipschitz condition
+$f(\pi(x+v))-f(x)\leq v^{\top}\nabla f(x)+\frac{L}{2}\|v\|^{2}\text{\ for all\ }v\in B_{r}(x)\cap T_{x}\mathcal{M}\,.$
+then for step-size $\gamma_{k}=\frac{1}{L}$ the RGD generates iterates $(x_{k},v_{k}=-\gamma_{k}\pi^{\prime}(x_{k})\nabla f(x_{k}))$ that satisfy
+$\min_{k=0,\ldots,N-1}\|\operatorname{grad}f(x_{k})\|\leq\frac{\sqrt{2L(f(x_{0})-f_{*})}}{\sqrt{N}}\,.$
+provided that $v_{k}\in B_{r}(x)\cap T_{x}\mathcal{M}$ for all $k=0,\ldots,N-1$.
+Now we set the following in Theorem 5 in our paper: $s=\pi$, $\sigma=0$.
+The pullback $L_{0}$-Lipschitz condition is satisfied under our assumption $L=\operatorname{Lip}(\nabla f)$ for a different constant $L_{0}=\operatorname{Lip}(\nabla(f\circ\pi))$, which can be estimated in terms of $L$, $C$, $\tau$ and $M$ (see the proof of Theorem 5).
+The rate obtained in Theorem 5 reads then
+$\min_{k=0,\ldots,N}\|\operatorname{grad}f(x_{k})\|^{2}\leq\frac{4D}{N\gamma_{\min}(1-L_{0}\gamma_{\max}/2)}\,$,
+where $\gamma_{\min}(1-L_{0}\gamma_{\max}/2)$ can be lower-bounded by a constant constructed from $\tau$ and $L_{0}$.
+Moreover, $2D$ is simply an estimation of $f(x_{0})-f_{*}$ via the triangle inequality (see the proof of Theorem 5).
+Thus in this case both theorems guarantee an $O(1/\sqrt{N})$ best-norm-of-gradient convergence rate with similar constants.
+Appendix B More on manifolds and distance functions
+In addition to the notation introduced in Section 2.1, we note that $\tau_{\mathcal{M}}$ is also the largest $\tau\geq 0$ such that the map $\{(p,v)\in\operatorname{N}\mathcal{M}\mid\lVert v\rVert<\tau\}\to\mathbb{R}^{d}:(p,v)\mapsto p+v$ is a diffeomorphism.
 By a tubular neighborhood of radius $\tau\in(0,\tau_{\mathcal{M}}]$ we mean a set of the form $\mathcal{T}(\tau)=\{p+v\mid p\in\mathcal{M}\,,v\in\operatorname{N}_{p}\mathcal{M}\,,\lVert v\rVert<\tau\}$.
@@
 The Weingarten map $S_{p}^{u}$ at a point $p\in\mathcal{M}$ in the direction $u\in\operatorname{N}_{p}\mathcal{M}$ is defined as the unique self-adjoint linear operator $S_{p}^{u}:\operatorname{T}_{p}\mathcal{M}\to\operatorname{T}_{p}\mathcal{M}$ such that $\left\langle w,S_{p}^{u}(v)\right\rangle_{\operatorname{T}_{p}\mathcal{M}}=\mathrm{I\!I}_{p}^{u}(v,w)$ for all $v,w\in\operatorname{T}_{p}\mathcal{M}$.
-A useful operator that has been studied in Abatzoglou (1978); Breiding & Vannieuwenhoven (2021); Alvarez et al.
+A useful operator that has been studied in Abatzoglou (1978); Breiding and Vannieuwenhoven (2021); Alvarez et al.
 (2004) is
 $H_{x}=I_{T_{\pi(x)}\mathcal{M}}+S_{\pi(x)}^{\pi(x)-x}:\operatorname{T}_{\pi(x)}\mathcal{M}\to\operatorname{T}_{\pi(x)}\mathcal{M}\,.$ (19)
-In Breiding & Vannieuwenhoven (2021) it has been shown (see Lemma 8) that $H_{x}$ is invertible on $\mathcal{T}$. Now we have the following useful identities that hold for $x\in\mathcal{T}$
+In Breiding and Vannieuwenhoven (2021) it has been shown (see Lemma 10) that $H_{x}$ is invertible on $\mathcal{T}$. Now we have the following useful identities that hold for $x\in\mathcal{T}$
 $\pi^{\prime}(x)$ $=\operatorname{I}_{\operatorname{T}_{\pi(x)}\mathcal{M}}H_{x}^{-1}\operatorname{P}_{\operatorname{T}_{\pi(x)}\mathcal{M}}$
@@
 and hence $\lVert\psi_{p}^{\prime}(0)\rVert\leq 1$ and $\lVert\psi_{p}^{\prime\prime}(0)\rVert=\lVert\mathrm{I\!I}_{p}\rVert\leq 1/\tau_{\mathcal{M}}$.
-A.1 Properties of (16)
+B.1 Properties of (19)
 We study the invertibility and boundedness of the operator (19).
@@
 The operator (19) is invertible on $\mathcal{T}(\tau_{\mathcal{M}})$.
-Moreover, (16) satisfies [^1] In (Breiding & Vannieuwenhoven, 2021, Theorem 4.3) the quantity $\lVert H_{x}^{-1}\rVert$ has been shown to equal the condition number of a certain critical point problem associated with $\mathcal{M}$.
+Moreover, (19) satisfies [^3] In (Breiding and Vannieuwenhoven, 2021, Theorem 4.3) the quantity $\lVert H_{x}^{-1}\rVert$ has been shown to equal the condition number of a certain critical point problem associated with $\mathcal{M}$.
 $\lVert P_{0}(x)\rVert=\lVert H_{x}^{-1}\rVert=\left(1-\frac{\lVert\pi(x)-x\rVert}{\rho(\pi(x),x-\pi(x))}\right)^{-1}\leq\frac{1}{1-\lVert x-\pi(x)\rVert\kappa_{\mathcal{M}}}\leq\frac{1}{1-\lVert x-\pi(x)\rVert/\tau_{\mathcal{M}}}\,$,
@@
 Proof.
-In (Breiding & Vannieuwenhoven, 2021, Lemma A.2) the following condition has been established: Let $\mathcal{S}=\{(a,p)\in\mathbb{R}^{n}\times\mathcal{M}\mid a-p\in\operatorname{N}_{p}\mathcal{M}\}$.
+In (Breiding and Vannieuwenhoven, 2021, Lemma A.2) the following condition has been established: Let $\mathcal{S}=\{(a,p)\in\mathbb{R}^{n}\times\mathcal{M}\mid a-p\in\operatorname{N}_{p}\mathcal{M}\}$.
 Then $\mathcal{S}$ is diffeomorphic to the normal bundle $\operatorname{N}\mathcal{M}$ via the diffeomorphism $\Phi:\operatorname{N}\mathcal{M}\to\mathcal{S}:(v,p)\mapsto(p+\operatorname{I}_{\operatorname{N}_{p}\mathcal{M}}(v),p)$.
@@
 $\geq\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}\eta^{2}\,.$
-A.2 Densities on manifolds
-Given a manifold $\mathcal{M}$ as in Section A and any chart $\psi:\mathcal{V}\to\mathcal{U}\subseteq\mathcal{M}$, the volume measure $\operatorname{Vol}_{\mathcal{M}}$ is uniquely defined by
+B.2 Densities on manifolds
+Given a manifold $\mathcal{M}$ as in Section B and any chart $\psi:\mathcal{V}\to\mathcal{U}\subseteq\mathcal{M}$, the volume measure $\operatorname{Vol}_{\mathcal{M}}$ is uniquely defined by
 $\operatorname{Vol}_{\mathcal{M}}(E)=\int_{\psi^{-1}(E)}G_{\psi}(z)\,\text{d}z\text{\ \ for\ \ }E\subseteq\mathcal{U}\text{\ Borel measurable}.$
@@
 $\lambda(0)=\mu(p)\,,\quad\nabla\lambda(0)=\operatorname{grad}_{\mathcal{M}}\mu(p)\in\operatorname{T}_{p}\mathcal{M}\cong\mathbb{R}^{k}\,.$ (23)
-Appendix B More on the Stein score function
+Appendix C More on the Stein score function
 In this section we derive the representations (4).
@@
 which yields the second representation in (4). ∎
-Appendix C Variance-exploding diffusion models
-In the variance-exploding scheme of score-based diffusion models Song et al. (2020); Tang & Zhao (2025), one considers the following stochastic differential equation on the finite interval $[0,T]$:
+Appendix D Variance-exploding diffusion models
+In the variance-exploding scheme of score-based diffusion models Song et al. (2020); Tang and Zhao (2025), one considers the following stochastic differential equation on the finite interval $[0,T]$:
 $\operatorname{d}X_{t}=\sqrt{2t}\operatorname{d}W_{t}\,,\quad t\in[0,T]\,,\quad X_{0}\sim\mu\,.$ (24)
@@
 The loss $L_{\operatorname{CSM}}$ can be evaluated because $\nabla\log p_{\sigma}(x\mid x_{0})=-\frac{1}{\sigma^{2}}(x-x_{0})$ is known explicitly.
-Appendix D Proof of Theorem 1
-D.1 Nonasymptotic Laplace method
+Appendix E Proof of Theorem 1
+E.1 Nonasymptotic Laplace method
 In this section we derive some non-asymptotic error estimates for the Laplace method Hashorva et al. (2015); Hwang (1980), which concerns itself with the asymptotic of integrals of the form
@@
 for some functions $f,h:\mathcal{V}\to\mathbb{R}$ from an open set $\mathcal{V}\subseteq\mathbb{R}^{k}$, where $h$ is non-negative and attains a unique minimum in, say, $z=0\in\mathcal{V}$.
-While there have been results providing such an error estimation to the first order expansion Inglot & Majerski (2014); Majerski (2015); Lapinski (2019), we need an error estimate up to the second order expansion, as provided in Theorem 14.
+While there have been results providing such an error estimation to the first order expansion Inglot and Majerski (2014); Majerski (2015); Lapinski (2019), we need an error estimate up to the second order expansion, as provided in Theorem 16.
 For the sake of completeness we also include the statement and proof for the first order expansion in Theorem 15.
@@
 $\begin{aligned} \frac{1}{\sigma^{2}}E(\sigma)&=\frac{\sqrt{\det\Sigma}}{Z_{\sigma}\sigma^{2}}\exp(-\min\{c\log(\sigma)^{2},\Delta\sigma^{-2}\})\lVert f\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}\\ &\qquad+\lVert\Sigma^{-1/2}\nabla^{2}f(0)\Sigma^{-1/2}\rVert\left(nC\sigma\lvert\log(\sigma)\rvert^{3}+\operatorname{G}_{2}\left(\lambda_{\min}(\Sigma)\min\{r\sigma^{-1},\lvert\log(\sigma)\rvert\}\right)\right)\\ &\qquad+C\sigma\lvert\log(\sigma)\rvert^{3}(1+2C\sigma\lvert\log(\sigma)\rvert^{3})\\ &\overset{\sigma\to 0}{=}O(\sigma\lvert\log(\sigma)\rvert^{3})=\widetilde{O}(\sigma)\,.\end{aligned}$ (37)
-D.2 Local version of Theorem 1
+E.2 Local version of Theorem 1
 First we prove a local version of Theorem 1.
@@
 $\leq\frac{4d}{\lambda(0)^{2}}\frac{E_{2}(\sigma)}{\sigma^{2}}+\left(\frac{4(\det\Sigma)(\lVert\mu\rVert_{2}+\lVert\mu\rVert_{1}^{2})}{\lambda(0)^{2}}\frac{e^{-\frac{1}{\sigma^{2}}\delta}}{Z_{\sigma}^{2}\sigma^{2}}+12\frac{D_{0}(\sigma)+\sqrt{\det\Sigma}\frac{e^{-\frac{1}{2\sigma^{2}}\delta}}{Z_{\sigma}}}{\lambda(0)}\right)\lVert P_{0}\rVert$
-D.3 Bounds in terms of distribution and manifold parameters
+E.3 Bounds in terms of distribution and manifold parameters
 […]
@@
 where $\lambda=\psi_{p}^{-1}\#\mu=\operatorname{pr}_{p}\#\mu$.
-D.3.1 Conditions (24) and (26) for $h$
+E.3.1 Conditions (27) and (29) for $h$
 First we show how to express conditions (27) and (29) in terms of the manifold parameters for $h$ given in (41).
@@
 Thus (29) holds on $\mathcal{V}$ with $c=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and implies (32).
-D.3.2 Lower bound on $\lambda_{\min}(\Sigma)$ and upper bound on $\sqrt{\det\Sigma}$
+E.3.2 Lower bound on $\lambda_{\min}(\Sigma)$ and upper bound on $\sqrt{\det\Sigma}$
 We first consider $\lambda_{\min}(\Sigma)$ for $\Sigma=\nabla^{2}h(0)$ with $h$ given in (41) and when $x\in\mathcal{T}(\tau)$. By Lemma 10 and (20) we directly obtain
@@
 $\sqrt{\det\Sigma}=\sqrt{\det(I_{\operatorname{T}_{p}\mathcal{M}}+S_{p}^{p-x})}\leq\left(1+\frac{\lVert p-x\rVert}{\rho(p,p-x)}\right)^{k/2}\leq\left(1+\frac{\tau}{\tau_{\mathcal{M}}}\right)^{k/2}\leq 2^{k/2}\,.$
-D.3.3 Condition (24) for $\overline{f}_{2}^{jl}$
+E.3.3 Condition (27) for $\overline{f}_{2}^{jl}$
 […]
@@
 can be obtained by direct differentiation. We don’t pursue the complete derivation here and instead say that (27) holds with, say, $C_{\tau_{\mathcal{M}},M,\mu}^{(2)}$.
-D.3.4 Upper bound on $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}$, $\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}$ and $\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}$ and $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert$
+E.3.4 Upper bound on $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}$, $\lVert f_{1}^{j}\rVert_{L^{1}(\mathcal{V})}$ and $\lVert\overline{f}_{2}^{jl}\rVert_{L^{1}(\mathcal{V}\times\mathcal{V})}$ and $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert$
 We clearly have $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}=\lambda(\mathcal{V})\leq 1$.
@@
 $\lVert\nabla^{2}\overline{f}_{2}^{jl}(0)\rVert\leq\lVert\overline{f}_{2}^{\prime\prime}(0,0)\rVert\leq 6\mu(p)^{2}+8\tau\mu(p)\lVert\operatorname{grad}_{\mathcal{M}}\mu(p)\rVert\,.$
-D.4 Proof of Theorem 1
+E.4 Proof of Theorem 1
 We apply Theorem 17 to every point $x\in\mathcal{T}(\tau)$ and $p=\pi(x)$ with the graph chart $\psi=\psi_{p}$.
@@
 First we bound $E_{1}(\sigma;f_{0},h,r,\theta)$ with $E_{1}$ given in (34).
-By Section D.3 we can take $r_{\tau_{\mathcal{M}},M}=\eta=\min\{\tau_{\mathcal{M}},M\}/4$, $c_{\tau_{\mathcal{M}},M}=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and $C_{\tau_{\mathcal{M}},M,\mu}=\max\{C_{\tau_{\mathcal{M}},M}^{(1)},C_{\tau_{\mathcal{M}},M,\mu}^{(2)}\}$ and hence, due to $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}\leq 1$,
+By Section E.3 we can take $r_{\tau_{\mathcal{M}},M}=\eta=\min\{\tau_{\mathcal{M}},M\}/4$, $c_{\tau_{\mathcal{M}},M}=\frac{1}{2}\frac{\tau_{\mathcal{M}}-\tau}{\tau_{\mathcal{M}}+\tau}$ and $C_{\tau_{\mathcal{M}},M,\mu}=\max\{C_{\tau_{\mathcal{M}},M}^{(1)},C_{\tau_{\mathcal{M}},M,\mu}^{(2)}\}$ and hence, due to $\lVert f_{0}\rVert_{L^{1}(\mathcal{V})}\leq 1$,
 $E_{1}(\sigma;f_{0})$ $\leq\frac{2^{k/2}}{Z_{\sigma}}\exp\left(-c_{\tau_{\mathcal{M}},M}\log(\sigma)^{2}\right)$
@@
 This shows that there $K$ and $\overline{\sigma}$ depending only on $\tau$, $M$ and $\operatorname{diam}(\mathcal{M})$ (as $k$) and $\mu$ such that (5) hold and finishes the proof.
-D.5 A useful corollary to Theorem 1
+E.5 A useful corollary to Theorem 1
 The following elementary consequence of Theorem 1 will be useful in the proofs of the results from Section 5.
 Corollary 19.
-Suppose that for some $s:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tau\in(0,\tau_{\mathcal{M}})$ and $\epsilon\leq\tau$ we have
-$\lVert s(x)-\pi(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (42)
-Then $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$.
-Moreover, for $x\in\mathcal{T}(\epsilon)$ we have $\lVert s(x)-x\rVert\leq 2\epsilon$ and for $x\in\mathcal{T}(\tau)$ we have $\lVert s(x)-x\rVert\leq 2\tau$.
+Suppose that for some $\texttt{v}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tau\in(0,\tau_{\mathcal{M}})$ and $\epsilon\leq\tau$ we have
+$\lVert\texttt{v}(x)-\pi(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (45)
+Then $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$.
+Moreover, for $x\in\mathcal{T}(\epsilon)$ we have $\lVert\texttt{v}(x)-x\rVert\leq 2\epsilon$ and for $x\in\mathcal{T}(\tau)$ we have $\lVert\texttt{v}(x)-x\rVert\leq 2\tau$.
 Proof.
-From (42) and the fact that $\pi(x)\in\mathcal{M}$ we clearly have $s(x)\in\mathcal{T}(\epsilon)$ for any $x\in\mathcal{T}(\tau)$, i.e.
-$s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$.
-The last claims follows then from the triangle inequality $\lVert s(x)-x\rVert\leq\lVert s(x)-\pi(x)\rVert+\lVert\pi(x)-x\rVert$.
+From (45) and the fact that $\pi(x)\in\mathcal{M}$ we clearly have $\texttt{v}(x)\in\mathcal{T}(\epsilon)$ for any $x\in\mathcal{T}(\tau)$, i.e.
+$\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon)$.
+The last claims follows then from the triangle inequality $\lVert\texttt{v}(x)-x\rVert\leq\lVert\texttt{v}(x)-\pi(x)\rVert+\lVert\pi(x)-x\rVert$.
 ∎
-Appendix E Approximate Riemannian gradient flow with landing
+Appendix F Approximate Riemannian gradient flow with landing
 In this section $C=\lVert\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})}\rVert_{\infty}$ and $L=\operatorname{Lip}(\nabla f|_{\mathcal{T}(\tau_{\mathcal{M}})})$ denote the supremum of $\nabla f$ and the Lipschitz constant of $\nabla f$ on $\mathcal{T}(\tau_{\mathcal{M}})$ for some manifold $\mathcal{M}$.
-We will often assume the following approximation condition on a function $s\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$:
-$\lVert s(x)-\pi(x)\rVert<\epsilon\,,\quad\lVert s^{\prime}(x)-P_{0}(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (43)
+We will often assume the following approximation condition on a function $\texttt{v}\in C^{1}(\mathbb{R}^{d};\mathbb{R}^{d})$:
+$\lVert\texttt{v}(x)-\pi(x)\rVert<\epsilon\,,\quad\lVert\texttt{v}^{\prime}(x)-P_{0}(x)\rVert<\epsilon\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$ (46)
 We will also abbreviate the (negative) right hand side of the dynamics (8) by
-$G_{s}^{\eta}(x)=s^{\prime}(x)\nabla f(x)+\eta(x-s(x))$ (44)
-E.1 Stationary points of (8) and optimality criteria
+$G_{\texttt{v}}^{\eta}(x)=\texttt{v}^{\prime}(x)\nabla f(x)+\eta(x-\texttt{v}(x))$ (47)
+F.1 Stationary points of (8) and optimality criteria
 We need to analyze the meaning of approximate stationary points of (8) for the optimization problem (1).
@@
 […]
-$\lVert G_{s}^{\eta}(x_{*})\rVert\leq\delta\,.$ (45)
+$\lVert G_{\texttt{v}}^{\eta}(x_{*})\rVert\leq\delta\,.$ (48)
 Then for $p_{*}=\pi(x_{*})$ it holds that
@@
 We have the estimates
-$\lVert\eta(s(x_{*})-x_{*})-\eta(\pi(x_{*})-x_{*})\rVert\leq\eta\epsilon\,$,
+$\lVert\eta(\texttt{v}(x_{*})-x_{*})-\eta(\pi(x_{*})-x_{*})\rVert\leq\eta\epsilon\,$,
 and
-$\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-s^{\prime}(x_{*})\nabla f(s^{\prime}(x_{*}))\rVert$
-$\hskip 56.9055pt\quad\leq\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-P_{0}(\pi(x_{*}))\nabla f(s(x_{*}))\rVert$
-$\hskip 56.9055pt\qquad+\lVert P_{0}(\pi(x_{*}))\nabla f(s(x_{*}))-P_{0}(x_{*})\nabla f(s(x_{*}))\rVert$
-$\hskip 56.9055pt\qquad+\lVert P_{0}(x_{*})\nabla f(s(x_{*}))-s^{\prime}(x_{*})\nabla f(s(x_{*}))\rVert$
+$\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-\texttt{v}^{\prime}(x_{*})\nabla f(\texttt{v}(x_{*}))\rVert$
+$\hskip 56.9055pt\quad\leq\lVert P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))-P_{0}(\pi(x_{*}))\nabla f(\texttt{v}(x_{*}))\rVert$
+$\hskip 56.9055pt\qquad+\lVert P_{0}(\pi(x_{*}))\nabla f(\texttt{v}(x_{*}))-P_{0}(x_{*})\nabla f(\texttt{v}(x_{*}))\rVert$
+$\hskip 56.9055pt\qquad+\lVert P_{0}(x_{*})\nabla f(\texttt{v}(x_{*}))-\texttt{v}^{\prime}(x_{*})\nabla f(\texttt{v}(x_{*}))\rVert$
 $\hskip 56.9055pt\quad\leq(\operatorname{Lip}(\nabla f)+\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty})\epsilon+\lVert P_{0}(\pi(x_{*}))-P_{0}(x_{*})\rVert\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}\,$,
 where we have used that $P_{0}(\pi(x_{*}))$ is an orthogonal projection and hence has unit norm.
-By Lemma 23 applied to $z_{1}=s^{\prime}(x_{*})\nabla f(s(x_{*}))$ and $z_{2}=\eta(s(x_{*})-x_{*})$, Lemma 9 and the fact that $P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))=\operatorname{grad}_{\mathcal{M}}f(\pi(x_{*}))$, the inequality (46) follows.
+By Lemma 25 applied to $z_{1}=\texttt{v}^{\prime}(x_{*})\nabla f(\texttt{v}(x_{*}))$ and $z_{2}=\eta(\texttt{v}(x_{*})-x_{*})$, Lemma 11 and the fact that $P_{0}(\pi(x_{*}))\nabla f(\pi(x_{*}))=\operatorname{grad}_{\mathcal{M}}f(\pi(x_{*}))$, the inequality (49) follows.
 ∎
-E.2 Lie derivative of manifold distance
+F.2 Lie derivative of manifold distance
 Lemma 21.
 Suppose that $\tau\in(0,\tau_{\mathcal{M}})$ and $\sigma>0$ is such that (46) is satisfied. Then it holds
-$-\left\langle\nabla\operatorname{d}(x),G_{s}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$
+$-\left\langle\nabla\operatorname{d}(x),G_{\texttt{v}}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\text{\ \ for all\ \ }x\in\mathcal{T}(\tau)\,.$
 Proof.
 We have for $x\in\mathcal{T}(\tau)$
-$\left\langle x-\pi(x),-G_{s}^{\eta}(x)\right\rangle$
-$\hskip 56.9055pt=-2\eta\operatorname{d}(x)+\left\langle\pi(x)-x,(s^{\prime}(x)-P_{0}(x))\nabla f(s(x))\right\rangle$
-$\hskip 56.9055pt\qquad+\eta\left\langle x-\pi(x),s(x)-\pi(x)\right\rangle$
-$\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+\lVert\pi(x)-x\rVert\lVert s^{\prime}(x)-P_{0}(x)\rVert\lVert\nabla f(s(x))\rVert$
-$\hskip 56.9055pt\qquad+\eta\lVert\pi(x)-x\rVert\lVert s(x)-\pi(x)\rVert$
+$\left\langle x-\pi(x),-G_{\texttt{v}}^{\eta}(x)\right\rangle$
+$\hskip 56.9055pt=-2\eta\operatorname{d}(x)+\left\langle\pi(x)-x,(\texttt{v}^{\prime}(x)-P_{0}(x))\nabla f(\texttt{v}(x))\right\rangle$
+$\hskip 56.9055pt\qquad+\eta\left\langle x-\pi(x),\texttt{v}(x)-\pi(x)\right\rangle$
+$\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+\lVert\pi(x)-x\rVert\lVert\texttt{v}^{\prime}(x)-P_{0}(x)\rVert\lVert\nabla f(\texttt{v}(x))\rVert$
+$\hskip 56.9055pt\qquad+\eta\lVert\pi(x)-x\rVert\lVert\texttt{v}(x)-\pi(x)\rVert$
 $\hskip 56.9055pt\leq-2\eta\operatorname{d}(x)+(\lVert\nabla f|_{\mathcal{T}(\tau)}\rVert_{\infty}\epsilon+\eta\epsilon)\sqrt{2\operatorname{d}(x)}\,.$
-E.3 Exact landing with $\sigma=0$
+F.3 Exact landing with $\sigma=0$
 Here we establish the following noiseless version of Theorem 3.
@@
 Since $\mathcal{T}\to\mathbb{R}:z\mapsto\lVert\operatorname{grad}f(z)\rVert_{H_{z}}^{2}$ is continuous, it is uniformly continuous on $\overline{\mathcal{T}}(\operatorname{dist}_{\mathcal{M}}(x(0)))$.
-This, together with (47) implies, by Barbalat’s lemma Farkas & Wegner (2016), that $\lim_{t\to\infty}\lVert\operatorname{grad}f(x(t))\rVert_{H_{x(t)}}^{2}=0$.
+This, together with (50) implies, by Barbalat’s lemma Farkas and Wegner (2016), that $\lim_{t\to\infty}\lVert\operatorname{grad}f(x(t))\rVert_{H_{x(t)}}^{2}=0$.
 Clearly any accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $H_{x_{*}}^{-1}\operatorname{grad}f(x_{*})=0$, i.e.
@@
 $\lVert\operatorname{grad}f(\pi(x_{*}))\rVert=\lVert\operatorname{grad}f(\pi(x_{*}))-\operatorname{grad}f(x_{*})\rVert\leq L\lVert\pi(x_{*})-x_{*}\rVert=L\tau\,.$
-E.4 Proof of Theorem 3
-Note first that, by Theorem 1, if (7) hold with $\epsilon\to\epsilon^{\prime}$, then 43 are satisfied with $\epsilon=\epsilon^{\prime}+K\sigma\lvert\log(\sigma)\rvert^{3}$ for $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$.
+F.4 Proof of Theorem 3
+Note first that, by Theorem 1, if (7) hold with $\epsilon\to\epsilon^{\prime}$, then 46 are satisfied with $\epsilon=\epsilon^{\prime}+K\sigma\lvert\log(\sigma)\rvert^{3}$ for $\sigma\in(0,\overline{\sigma}(\tau,\mathcal{M},\mu))$.
 In the following we will write $\epsilon>0$ for the latter quantity.
@@
 We have by Lemma 21 whenever $x(t)\in\mathcal{T}(\tau)$ that
-$\frac{\operatorname{d}}{\operatorname{d}t}\operatorname{d}(x)=-\left\langle\nabla\operatorname{d}(x),G_{s}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\,$,
+$\frac{\operatorname{d}}{\operatorname{d}t}\operatorname{d}(x)=-\left\langle\nabla\operatorname{d}(x),G_{\texttt{v}}^{\eta}(x)\right\rangle\leq-2\eta\operatorname{d}(x)+\epsilon(C+\eta)\sqrt{2\operatorname{d}(x)}\,$,
 where the right hand side is non-positive iff
@@
 the flow (8) and $x(t)$ exists for all $t\geq 0$.
-Moreover, by a similar argument as in the standard proof of Lyapunov’s direct method Khalil & Grizzle (2002), for each $\delta>0$ there exists some $T>0$ such that for all $t\geq T$ it holds that $x(t)\in\mathcal{T}(\tau_{0}+\delta)$.
-If $s$ is a gradient field, i.e.
-$s=\nabla g$ for some function $g\in C^{1}(\mathbb{R}^{d})$, then $V(x):=f(s(x))+\eta(\lVert x\rVert^{2}/2-g(x))$ satisfies $\nabla V(x)=G_{s}^{\eta}(x)$.
-Similar as in the proof of Theorem 20 we can take $G_{s}$ as a Lyapunov function to obtain
-$\frac{\operatorname{d}}{\operatorname{d}t}V(x)=-\lVert G_{s}^{\eta}(x)\rVert^{2}\,$
+Moreover, by a similar argument as in the standard proof of Lyapunov’s direct method Khalil and Grizzle (2002), for each $\delta>0$ there exists some $T>0$ such that for all $t\geq T$ it holds that $x(t)\in\mathcal{T}(\tau_{0}+\delta)$.
+If v is a gradient field, i.e.
+$\texttt{v}=\nabla g$ for some function $g\in C^{1}(\mathbb{R}^{d})$, then $V(x):=f(\texttt{v}(x))+\eta(\lVert x\rVert^{2}/2-g(x))$ satisfies $\nabla V(x)=G_{\texttt{v}}^{\eta}(x)$.
+Similar as in the proof of Theorem 22 we can take $G_{\texttt{v}}$ as a Lyapunov function to obtain
+$\frac{\operatorname{d}}{\operatorname{d}t}V(x)=-\lVert G_{\texttt{v}}^{\eta}(x)\rVert^{2}\,$
 along the dynamics (8).
-By a similar argument $V(x)$ is non-increasing and every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $G_{s}^{\eta}(x)(x_{*})=0$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$.
-Note that the condition $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\tau)$ in Lemma 18 is satisfied when $\epsilon\leq\tau$.
+By a similar argument $V(x)$ is non-increasing and every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $G_{\texttt{v}}^{\eta}(x)(x_{*})=0$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$.
+Note that the condition $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\tau)$ in Lemma 20 is satisfied when $\epsilon\leq\tau$.
 Applying Lemma 20 with $\delta=0$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and proves Theorem 3 for the case when $s$ is a gradient field.
-Now consider the more general case when $s$ is not necessarily a gradient field.
-In this case consider $V(x)=f(s(x))+\eta\operatorname{d}(x)$ to obtain
-$\frac{\operatorname{d}}{\operatorname{d}t}V(x)$ $=-\left\langle s^{\prime}(x)\nabla f(s(x))+\eta(x-\pi(x)),G_{s}^{\eta}(x)\right\rangle$
-$\leq-\lVert G_{s}^{\eta}(x)\rVert^{2}+\eta\lVert s(x)-\pi(x)\rVert\lVert G_{s}^{\eta}(x)\rVert$
-$\leq-\lVert G_{s}^{\eta}(x)\rVert^{2}+\eta\epsilon\lVert G_{s}^{\eta}(x)\rVert\,.$
-Thus, if $\lVert G_{s}^{\eta}(x)\rVert\geq\eta\epsilon$ we have $\frac{\operatorname{d}}{\operatorname{d}t}V(x)\leq 0$.
-By a barrier function argument Khalil & Grizzle (2002) this implies that every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $\lVert G_{s}^{\eta}(x_{*})\rVert\leq\eta\epsilon$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$.
+Now consider the more general case when v is not necessarily a gradient field.
+In this case consider $V(x)=f(\texttt{v}(x))+\eta\operatorname{d}(x)$ to obtain
+$\frac{\operatorname{d}}{\operatorname{d}t}V(x)$ $=-\left\langle\texttt{v}^{\prime}(x)\nabla f(\texttt{v}(x))+\eta(x-\pi(x)),G_{\texttt{v}}^{\eta}(x)\right\rangle$
+$\leq-\lVert G_{\texttt{v}}^{\eta}(x)\rVert^{2}+\eta\lVert\texttt{v}(x)-\pi(x)\rVert\lVert G_{\texttt{v}}^{\eta}(x)\rVert$
+$\leq-\lVert G_{\texttt{v}}^{\eta}(x)\rVert^{2}+\eta\epsilon\lVert G_{\texttt{v}}^{\eta}(x)\rVert\,.$
+Thus, if $\lVert G_{\texttt{v}}^{\eta}(x)\rVert\geq\eta\epsilon$ we have $\frac{\operatorname{d}}{\operatorname{d}t}V(x)\leq 0$.
+By a barrier function argument Khalil and Grizzle (2002) this implies that every accumulation point $x_{*}$ of $\{x(t)\mid t\geq 0\}$ satisfies $\lVert G_{\texttt{v}}^{\eta}(x_{*})\rVert\leq\eta\epsilon$ and belongs to $\overline{\mathcal{T}}(\tau_{0})$.
 Again applying Lemma 20 with $\delta=\eta\epsilon$ and $\widetilde{\tau}=\tau_{0}$ yields (11) and finishes the proof.
-Appendix F Discretized Riemannian gradient flow and descent
+Appendix G Discretized Riemannian gradient flow and descent
 In this section we provide proof of Theorem 5 as well as analysis of the approximate Riemannian gradient descent and the discretized landing flow.
@@
 If $\tau\in(0,\tau_{\mathcal{M}})$ and (46) holds for some $\epsilon>0$, then
-$\sup_{x\in\mathcal{T}(\tau)}\lVert s^{\prime}(x)\rVert_{\infty}\leq\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}}\,.$
+$\sup_{x\in\mathcal{T}(\tau)}\lVert\texttt{v}^{\prime}(x)\rVert_{\infty}\leq\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}}\,.$
 Additionally, if $\epsilon\in(0,\tau]$, then
-$\lVert G_{s}^{\eta}(x)\rVert\leq C\lVert s^{\prime}(x)\rVert+\eta\lVert s(x)-x\rVert\leq C(\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}})+2\eta\tau\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$
+$\lVert G_{\texttt{v}}^{\eta}(x)\rVert\leq C\lVert\texttt{v}^{\prime}(x)\rVert+\eta\lVert\texttt{v}(x)-x\rVert\leq C(\epsilon+\frac{1}{1-\tau/\tau_{\mathcal{M}}})+2\eta\tau\text{\ \ for\ \ }x\in\mathcal{T}(\tau)$
 Proof.
 Via triangle inequality, Lemma 10 and the definition of $C$. ∎
-F.1 Discretized approximate Riemannian gradient flow
+G.1 Discretized approximate Riemannian gradient flow
 In this section we analyze the discretized version of (9), specifically the corresponding gradient descent
@@
 ∎
-F.2 Proof of Theorem 5
+G.2 Proof of Theorem 5
 Let us write $\epsilon^{\prime}=\epsilon+K(\tau,\mathcal{M},\mu)\sigma\lvert\log(\sigma)\rvert^{3}\leq\tau/2$.
-First we note that $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\epsilon^{\prime})$ as soon as $x_{0}\in\mathcal{T}(\tau/2)$, because $s(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon^{\prime})$ and $x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k})\in\mathcal{T}(\tau)$, since
-$\gamma_{k}\leq\frac{\tau}{2C(\epsilon^{\prime}+(1-\tau/\tau_{\mathcal{M}})^{-1})}\leq\frac{\tau}{2\lVert s^{\prime}(x_{k})\nabla f(x_{k})\rVert}\,$,
+First we note that $\{x_{k}\}_{k=0}^{\infty}\subseteq\mathcal{T}(\epsilon^{\prime})$ as soon as $x_{0}\in\mathcal{T}(\tau/2)$, because $\texttt{v}(\mathcal{T}(\tau))\subseteq\mathcal{T}(\epsilon^{\prime})$ and $x_{k}-\gamma_{k}\texttt{v}^{\prime}(x_{k})\nabla f(x_{k})\in\mathcal{T}(\tau)$, since
+$\gamma_{k}\leq\frac{\tau}{2C(\epsilon^{\prime}+(1-\tau/\tau_{\mathcal{M}})^{-1})}\leq\frac{\tau}{2\lVert\texttt{v}^{\prime}(x_{k})\nabla f(x_{k})\rVert}\,$,
 where first inequality is due to $(1-\tau/\tau_{\mathcal{M}})^{-1}\leq 2$ and the second inequality due to Lemma 23.
-For brevity let us denote $y_{k}=x_{k}-\gamma_{k}s^{\prime}(x_{k})\nabla f(x_{k})$ and $z_{k}=x_{k}-\gamma_{k}P_{0}(x_{k})\nabla f(\pi(x_{k}))$.
+For brevity let us denote $y_{k}=x_{k}-\gamma_{k}\texttt{v}^{\prime}(x_{k})\nabla f(x_{k})$ and $z_{k}=x_{k}-\gamma_{k}P_{0}(x_{k})\nabla f(\pi(x_{k}))$.
 Also let $L_{0}=\operatorname{Lip}(\nabla(f\circ\pi)|_{\mathcal{T}(\tau)})$.
@@
 Then, since $\lVert x_{k}-\pi(x_{k})\rVert\leq\epsilon^{\prime}$ for $k\geq 1$, we have
-$f(x_{k+1})-f(x_{k})$ $=f(s(y_{k}))-f(x_{k})$
+$f(x_{k+1})-f(x_{k})$ $=f(\texttt{v}(y_{k}))-f(x_{k})$
 $=f(\pi(z_{k}))-f(\pi(x_{k}))+f(\pi(x_{k}))-f(x_{k})$
-$\qquad+f(s(y_{k}))-f(\pi(y_{k}))+f(\pi(y_{k}))-f(\pi(z_{k}))$
+$\qquad+f(\texttt{v}(y_{k}))-f(\pi(y_{k}))+f(\pi(y_{k}))-f(\pi(z_{k}))$
 $\leq-\gamma_{k}\left\langle P_{0}(x_{k})\nabla f(\pi(x_{k})),P_{0}(x_{k})\nabla f(\pi(x_{k}))\right\rangle+\gamma_{k}^{2}\frac{L_{0}}{2}\lVert P_{0}(x_{k})\nabla f(\pi(x_{k}))\rVert^{2}$
@@
 This finishes the proof.
-F.3 Auxiliary results
-F.3.1 A lemma on norms of orthogonal vectors
+G.3 Auxiliary results
+G.3.1 A lemma on norms of orthogonal vectors
 Lemma 25.
@@
 ∎
-F.3.2 Gaussian tail bounds
+G.3.2 Gaussian tail bounds
 The following elementary lemmas give simplifications for some of the constants appearing in the proof of Theorem 15 and Theorem 16.
@@
 See Majerski (2015). ∎
-Appendix G Further numerical experiments and implementation details
-G.1 Optimization over $O(n)$
-G.1.1 Implementation details
+Appendix H Further numerical experiments and implementation details
+H.1 Optimization over $O(n)$
+H.1.1 Implementation details
 In all of our experiments, we discretize the flow equation 8 using the Euler scheme with a step size of $t_{\operatorname{step}}=1\cdot 10^{-4}$ and set the landing gain $\eta=3\cdot 10^{3}$.
@@
 We use $N_{\operatorname{epochs}}=10000$ and $N_{\operatorname{epochs}}=50000$ epochs for $n=10$ and $n=20$, respectively.
-G.1.2 Experiments
-Refer to caption Refer to caption
-Figure 3: Objective value vs. flow time $t$ for the orthogonal manifold for $n=10$ (left, different $\sigma>0$) and for $n=20$ (right, $\sigma=0.05$)
-G.2 Data-driven reference tracking
-G.2.1 Benchmark systems and tracking goals
+H.2 Data-driven reference tracking
+H.2.1 Benchmark systems and tracking goals
 Benchmark systems: We consider two classical benchmark systems: The double pendulum and the unicycle car model LaValle (2006).
@@
 Here $R$ is some small penalty on the input $\bm{u}$ to keep it bounded during the optimization.
-G.2.2 Implementation details
+H.2.2 Implementation details
 Discretization: We discretize both continuous-time dynamics $\dot{x}=\bm{f}_{\operatorname{cont}}(x,u)$ via the RK4-method and discretization step $\Delta t$ to obtain the discrete-time dynamics (13) as
@@
 the temporal dimension and channels.
-Diffusion and training parameters: Same as in Appendix G.1.1 with this time $N_{\operatorname{epochs}}=50000$ training epochs. For the DRGD step-size we pick a fixed step-size of $\gamma=0.001$.
-G.2.3 Experiments
+Diffusion and training parameters: Same as in Appendix H.1.1 with this time $N_{\operatorname{epochs}}=50000$ training epochs. For the DRGD step-size we pick a fixed step-size of $\gamma=0.001$.
+H.2.3 Experiments
 In Figure 4 we present the objective value evolution for the experiment from Section 6.2.